### v1

# A Reduction-Driven Local Search for the Generalized Independent Set Problem

Abstract
The Generalized Independent Set (GIS) problem extends the classical maximum independent set problem by incorporating profits for vertices and penalties for edges. This generalized problem has been identified in diverse applications in fields such as forest harvest planning, competitive facility location, social network analysis, and even machine learning. However, solving the GIS problem in large-scale, real-world networks remains computationally challenging. In this paper, we explore data reduction techniques to address this challenge. We first propose 14 reduction rules that can reduce the input graph with rigorous optimality guarantees. We then present a reduction-driven local search (RLS) algorithm that integrates these reduction rules into the pre-processing, the initial solution generation, and the local search components in a computationally efficient way. The RLS is empirically evaluated on 278 graphs arising from different application scenarios. The results indicates that the RLS is highly competitive – For most graphs, it achieves significantly superior solutions compared to other known solvers, and it effectively provides solutions for graphs exceeding 260 million edges, a task at which every other known method fails. Analysis also reveals that the data reduction plays a key role in achieving such a competitive performance.
Keywords: Generalized independent set; Data reduction; Local search; Large graph
1 Introduction
Graphs are essential tools for modeling complex systems including social networks, geometric networks, and forest management systems. The analysis of graph data has led to the development of numerous combinatorial problems that help address real-world challenges. In this paper, we study the generalized independent set (GIS) problem. Let $G=(V,E=E_{p}\cup E_{r})$ be an undirected graph where $V$ is the vertex set and $E$ is the edge set that is partitioned into a permanent edge set $E_{p}$ and a removable edge set $E_{r}$. Each vertex $v\in V$ is assigned a profit $w(v)\in\mathbb{R}$. Each removable edge $e\in E_{r}$ is associated with a penalty $p(e)\in\mathbb{R}$. The generalized independent set of the graph $G$ is a vertex set $I\subseteq V$ such that no permanent edge exists between any two vertices of $I$ in $G$. Then, the GIS problem asks for a generalized independent set $I$ of maximum net benefit, which is the total vertex profits of $I$ minus the total penalties of removable edges whose end vertices are in $I$. An example is shown in Figure 1.
Figure 1: A toy example of the GIS problem. The set of red vertices is a generalized independent set of net benefit of $2-1+6+5=12$.
When all the edges are permanent edges, the GIS problem equals the well-known maximum independent set problem (MIS) which asks for a maximum set of non-adjacent vertices from a given graph. As the GIS allows a more flexible definition of vertex profit and edge penalties, it captures a broader range of real-world scenarios than the classical independent set problem. We leave detailed examples of its applicability in diverse domains in Section 2.1. However, since MIS is NP-hard, W[1]-hard and difficult to approximate (Cygan et al. 2015), the equivalence indicates that the GIS is also computationally hard.
Existing works have proposed various approaches to solve the GIS problem practically. Exact solution methods include mixed integer linear programming (Colombi et al. 2017) and unconstrained binary quadratic programming (Hosseinian and Butenko 2019). They can solve instances with up to 400 vertices and 71,820 edges in three hours so far (Zheng et al. 2024). On the other hand, heuristic search methods have demonstrated their effectiveness in providing high-quality solutions within reasonable computational time (e.g., Colombi et al. (2017), Nogueira et al. (2021)). For example, the state-of-the-art ALS algorithm proposed by Zheng et al. (2024) achieves near-optimal solutions on all standard benchmark instances and scales to problems with up to 18,000 vertices. Nevertheless, we observe that these heuristic algorithms struggle to maintain solution quality in even larger instances.
The data reduction, which reduces input size while preserving problem optimality, has recently emerged as a promising technique for improving algorithm scalability on large instances. The technique has been successfully applied to problems such as independent set, vertex cover, and maximum clique (Großmann et al. 2023, Abu-Khzam et al. 2022, Verma et al. 2015, Walteros and Buchanan 2020, Lamm et al. 2019). For the GIS problem, research on data reduction is still in its early stages. There are three main challenges in developing data reduction techniques for the GIS problem: (1) The complex interplay among vertex profits, edge penalties, and the underlying graph structure, (2) Translating reduction rules into efficient, practical algorithms, often requiring incremental implementations (Akiba and Iwata 2016, Chang et al. 2017), (3) Integrating reductions into local search algorithms while balancing search time and efficiency.
In this paper, we address these challenges by proposing a novel reduction-driven local search algorithm for solving the GIS problem. Our contributions are summarized as follows.
  1. 1.
    To the best of our knowledge, this is the first work to apply data reduction techniques to the GIS problem. We introduce 14 reduction rules that are classified into four groups: edge transformation, neighborhood-based reduction, low-degree reduction, and vertex-pair reduction. We provide rigorous proofs that these rules preserve the problem optimality. To efficiently implement these reduction rules, we introduce an ordering strategy and construct a dependency graph to manage their applications.
  2. 2.
    We propose an efficient reduction-driven local search algorithm, named RLS, to solve the GIS problem. We incorporate our proposed reduction rules into three main components of RLS: the pre-processing, which incrementally reduces the size of original input instance, the initial solution generation, which interplays the vertex selection and data reduction, and the neighborhood search, which uses a novel reduction-based move operator to guide the search trajectory.
  3. 3.
    We evaluate the proposed RLS algorithm on a total of 278 instances. On small instances, RLS consistently achieves the known optimal solutions with significantly shorter running time than exact algorithms. On large instances, RLS exhibits superior scalability over existing heuristic algorithms in solution quality. Notably, RLS successfully handles instances with over 260 million edges, where existing heuristic methods fail. Our experimental results also demonstrate that data reduction techniques significantly improve the computational efficiency of RLS.
The rest of the paper is organized as follows. Section 2 reviews applications and related works, followed by Section 3 which introduces the notations and definitions. In Section 4, we present our reduction rules for GIS. Then, in Section 5, we describe the RLS algorithm. Section 6 presents computational results and comparisons with existing methods. In the last section, we conclude the paper and outline future directions.
2 Background
2.1 Applications Related to GIS
In the following, we highlight several important applications of the GIS problem:
Forest harvest planning. The planning of forest harvest involves selecting forest areas (cells) for timber, aiming to maximize benefits like timber value while minimizing ecological damage from excessive cutting. As a GIS problem, forest cells are vertices weighted by harvesting benefits, and edges represent penalties for harvesting adjacent cells (Hochbaum and Pathria (1997)).
Competitive facility location. In determining optimal locations for service facilities, each potential site has a utility value. However, the utility diminishes when multiple facilities compete for customers in the same area. The facility location problem, modeled with vertices as facilities and edges as geographic relationships, can be framed as a GIS problem to maximize utilities (Hochbaum (2004)). Similar spatial conflicts appear in airplane seating (Pavlik et al. 2021) and label placement (Mauri et al. 2010).
Social network analysis. The GIS problem has broad applications in social network analysis. A maximum independent set corresponds to the complement of a minimum vertex cover and is equivalent to a maximum clique in the graph’s complement (Sanchis and Jagota (1996)). These properties make GIS useful for tasks such as detecting cliques or communities (Gschwind et al. (2021)) and optimizing coverage and reach within social networks (Puthal et al. (2015)).
Other applications. GIS has been applied in machine learning to aggregate low-quality results into better results (Brendel and Todorovic 2010), to track objects in multi-object tracking tasks (Brendel et al. 2011), and to serve as a pooling layer in graph neural networks (Nouranizadeh et al. 2021).
2.2 Existing Algorithms
The GIS problem was first introduced in Hochbaum and Pathria (1997). Existing methods for solving the GIS problem can be mainly divided into exact and heuristic algorithms.
Exact Algorithms. In pursuit of optimal solutions, Hochbaum (2004) introduced an integer programming model that considers both nodes and edges as variables, with each edge associated with a specific constraint. However, this model relies on binary variables for removable edges, resulting in large formulations; for example, a graph with 400 vertices requires 80,000 variables, making direct linear programming infeasible. To address this, Colombi et al. (2017) reformulated the GIS problem as an unconstrained binary quadratic program (UBQP), and developed a branch-and-cut algorithm, along with additional valid inequalities. Note that they also generated a total of 216 random instances which often serve as benchmark datasets in the later work. Hosseinian and Butenko (2019) employed another linear program to bound the optimal solution, which is later incorporated into a branch-and-bound framework. Recently, Zheng et al. (2024) used a Lagrangian relaxation in their branch-and-bound framework to estimate tight upper bounds. They showed that their algorithm is able to solve the GIS instances with up to 400 vertices within $3$ hours.
Heuristic Algorithms. Kochenberger et al. (2007) also developed a UBQP formulation for the GIS problem solved via tabu search. Colombi et al. (2017) enhanced performance by integrating advanced tabu search with UBQP. Nogueira et al. (2021) directly solved the original problem by using an iterated local search heuristic. They explored move operations of adding one or two vertices to the incumbent solution. More recently, Zheng et al. (2024) proposed a more refined local search heuristic, namely ALS, which explores three move operations – adding one vertex to the solution, dropping one vertex from the solution, and replacing a vertex in the solution with another one not in the solution. The ALS algorithm reported state-of-the-art results on the set of 216 benchmark instances proposed by Colombi et al. (2017). Their algorithm is able to find high-quality solutions on large instances with up to $17903$ vertices within $5$ minutes. However, its performance on even larger graphs, such as the uk2002 graph with 18 million vertices, remains unknown.
Data Reduction Techniques. The data reduction techniques are originally introduced in the framework of parameterized complexity analysis by the name of kernelization (Cygan et al. 2015). Recently, data reduction techniques are intensively used for practically tackling MIS (Abu-Khzam et al. 2022). For example, Akiba and Iwata (2016) shows that branch-and-reduce outperforms pure branch-and-bound, while Dahlum et al. (2016) and Chang et al. (2017) combine reductions with local search and heuristics to achieve near-optimal performance. These techniques enable handling graphs with millions of vertices and edges (Lamm et al. 2019, Xiao et al. 2021). Besides, the effectiveness of reductions has also been acknowledged in other fundamental problems like the clique problem (Walteros and Buchanan 2020), and the matching problem (Koana et al. 2021). However, as for the GIS problem, no existing work focuses on reduction rules. And extending existing reductions to vertex- and edge-weighted graphs for the GIS problem remains challenging.
3 Preliminary
By the definition, the input instance of the GIS problem is an (undirected weighted) graph $G=(V,E=E_{p}\cup E_{r})$ where each vertex $v\in V$ is associated with a profit $w(v)\in\mathbb{R}$ and each removable edge $e\in E_{r}$ is associated with a penalty $p(e)$. In order to reformulate the problem in a uniform way, we extend the definition by associating each edge $e\in E_{p}$ also with a penalty $p(e)=M$ where $M\in R$ is a sufficiently large value, e.g., the absolute sum of all vertex profits and removable edge penalties, $\sum_{v\in V}|w(v)|+\sum_{e\in E_{r}}|p(e)|$). For a set $S\subseteq V$, we formally define $nb(S)$ as the net benefit of $S$, i.e.
$nb(S)=\sum_{v\in S}w(v)-\sum_{\{u,v\}\subseteq S,(u,v)\in E}p(u,v).$
Then, the GIS problem is equal to finding a set $S\subseteq V$ that maximizes the net benefit of $S$ from a given graph $G$. The optimal objective value of the GIS problem in $G$ is denoted by $\alpha(G)$, i.e.
$\alpha(G)=\max_{S\subseteq V}nb(S)$
We also denote $MGIS(G)$ as the set of all vertex sets $S$ with optimal objective value. It is clear that every $S\in MGIS(S)$ is a generalized independent set which is defined in the introductory section. We say a generalized independent set $S\subseteq V$ is maximal if there is no other generalized independent set $S^{\prime}$ such that $S\subset S^{\prime}$ and $nb(S^{\prime})\geq nb(S)$.
For convenience, we import more notations in Table 1. Note that we use $\tilde{w}(v)$ to represent the profit of $v$ plus the non-negative penalties of edges that are incident to $v$. The $w^{+}(S)$ is an upper bound on the net benefit of any subset of $S$. For example, in Figure 1, $\tilde{w}(u)=2+1=3$. When $S$ denotes the set of red vertices in Figure 1, $w^{+}(S)=3+6+5=14$.
In the rest of the paper, we use notation $\coloneqq$ to represent assignment and $=$ to indicate equivalence.
Table 1: Table of Notations
$G=(V,E)$ a graph $G$ with vertex set $V$ and edge set $E$
$E_{p},E_{r}$ the set of pernament edges and the set of removable edges respectively
$w(v),p(u,v)$ the profit of including a vertex $v$, and the penalty of including both vertices $u$ and $v$
$N(v),N_{p}(v),N_{r}(v)$ the sets of all neighbors, permanent neighbors and removable neighbors of $v$, respectively
$|N(v)|,|N_{p}(v)|,|N_{r}(v)|$ the degree, permanent degree, and removable degree of $v$, respectively
$N[v],N_{p}[v],N_{r}[v]$
the sets of all neighbors, permanent neighbors and removable neighbors of $v$ including $v$
itself, respectively
$\Delta$ the maximum degree of vertices in $G$
$\tilde{w}(v)$
$\tilde{w}(v)=w(v)+\sum_{x\in N_{r}(v)}\max(0,-p(v,x))$
$S,G\setminus S$ a vertex subset of $G$, and the graph obtained by removing $S$ and all incident edges
$w(S),w^{+}(S)$ $w(S)=\sum_{v\in S}w(v)$, $w^{+}(S)=\sum_{v\in S}\max(0,\tilde{w}(v))$
$nb(S)$ the net benefit of vertex set $S$, $nb(S)=\sum_{v\in S}w(v)-\sum_{\{u,v\}\subseteq S}p(u,v)$
$\alpha(G),MGIS(G)$ the largest net benefit value, and the set of all vertex sets whose net benefit is $\alpha(G)$
Integer Programming. The GIS problem can be easily formulated as (Colombi et al. 2017):
$\alpha(G)=\max\quad$ $\sum_{i\in V}w(i)x_{i}-\sum_{\{i,j\}\in E_{r}}p(i,j)y_{ij}$ (1)
s.t. $x_{i}+x_{j}\leq 1$ $(i,j)\in E_{p}$ (2)
$x_{i}+x_{j}-y_{ij}\leq 1$ $(i,j)\in E_{r}$ (3)
$x_{i}\in\{0,1\}$ $i\in V$ (4)
$y_{ij}\in\{0,1\}$ $(i,j)\in E_{r}$ (5)
where binary variables $x_{i}$ indicate whether node $i$ is selected, and binary variables $y_{ij}$ indicate whether the edge is included.
4 Reduction Rules
In this section, we develop reduction rules that identify reducible graph structures without compromising optimality. We rely on four operations to reduce the current instance to another one. (1) Edge transformation, which identifies removable edges that can be deleted or transferred to permanent edges; (2) Vertex inclusion, which identifies vertices that are in some set $I\in MGIS(G)$; (3) Vertex exclusion, which identifies vertices that are not in some set $I\in MGIS(G)$. (4) Folding, which identifies vertices that can be merged into a single vertex. Based on the four operations, we will introduce a total of 14 reduction rules.
4.1 Edge Transformation
We first introduce two edge transformation rules.
R 1 (Edge Removal Reduction).
If there exists an edge $(u,v)\in E$ such that $p(u,v)=0$, then $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}$ is obtained by removing the edge $(u,v)$ from $G$.
Technical proofs of the correctness of this reduction rule, as well as the following reduction rules, are deferred in Appendix A. R1 implies that removing an edge $(u,v)$ that $p(u,v)=0$ does not change optimality.
R 2 (Edge-penalty Reduction).
If there is an edge $(u,v)$ such that $p(u,v)<M$ and $p(u,v)>min(\tilde{w}(u),\tilde{w}(v))$, then $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}$ is obtained by setting $p(u,v):=M$ in $G$.
Compared to permanent edges, removable edges are more challenging to reduce because of the difficulties of identifying whether both endpoints of a removable edge are in some $S\in MGIS(G)$. R2 allows us to transform removable edges into permanent edges without sacrificing optimality, underpinning many further reduction rules.
4.2 Neighborhood-based Reductions
We further consider the local neighborhood of a vertex.
R 3 (Neighborhood Weight Reduction).
If there is a vertex $u\in V$ that $w(u)\geq w^{+}(N(u))$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus N_{p}[u]$, and $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$.
The neighborhood weight reduction generalizes the neighborhood reduction rule in (Lamm et al. 2019) for solving the MIS problem. We further extend R3 by considering more neighboring structures and obtain R4-R6.
R 4 (Neighborhood Penalty Reduction).
If there exists a vertex $u\in V$ such that $w(u)\geq w^{+}(N_{p}(u))+\sum_{x\in N_{r}(u)}\max(0,p(u,x))$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus N_{p}[u]$ and $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$.
R 5 (Negative Profit Reduction).
If there exists a vertex $u\in V$ such that $w(u)<0$ and $\sum_{v\in N(u)}\min(0,p(u,v))>w(u)$, then $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$.
R 6 (Clique Reduction).
If there exists a vertex $u\in V$ such that $N_{p}(u)$ is a clique and $w(u)\geq\sum_{v\in N_{r}(u)}\max(0,p(u,v))+\max(0,\max_{v\in N_{p}(u)}\tilde{w}(v))$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus N_{p}[u]$ and $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$.
4.3 Low-degree Reductions
Low-degree vertices such as degree-one and degree-two vertices are common in large real-world graphs. Next, we introduce five reduction rules to reduce these vertices.
4.3.1 Degree-one Vertices Reduction
R 7 (Degree-one Vertices Reduction).
Let $u$ be a vertex with a single neighbor $v$.
Case 1: $w(u)\geq p(u,v)$. (1) If $w(u)\geq 0$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G\setminus\{u\}$ and $w(v)\coloneqq w(v)-p(u,v)$. (2) If $w(u)<0$, then $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G\setminus\{u\}$, and $w(v)\coloneqq w(v)-p(u,v)+w(u)$.
Case 2: $w(u)<p(u,v)$. (1) If $w(u)\geq 0$ and $w(u)\geq\tilde{w}(v)$, then $\alpha(G)=\alpha(G\setminus\{u,v\})+w(u)$. (2) If $w(u)\geq 0$ and $w(u)<\tilde{w}(v)$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G\setminus\{u\}$ and $w(v)\coloneqq w(v)-w(u)$. (3) If $w(u)<0$, then $\alpha(G)=\alpha(G\setminus\{u\})$.
Note that R7 covers all degree-one vertices. In other words, any vertex of degree one can be removed using this rule.
4.3.2 Degree-two Vertices Reductions
R 8 (Permanently Connected Neighbors Reduction).
Let $u$ be a vertex with exactly two neighbors $x,y$ and $p(x,y)=M$. The following table shows the reduction rules for reducing $u$ under different conditions.
conditions reduction
Case 1: $w(u)\geq\max(p(u,x),p(u,y))$ $w(u)\geq 0$
R8.1.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G\setminus\{u\}$,
$w(x)\coloneqq w(x)-p(u,x)$, $w(y)\coloneqq w(y)-p(u,y)$.
$w(u)<0$
R8.1.2: $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G\setminus\{u\}$,
$w(x)\coloneqq w(x)+w(u)-p(u,x)$, $w(y)\coloneqq w(y)+w(u)-p(u,y)$.
Case 2: $p(u,x)>w(u)\geq p(u,y)$ w.l.o.g, assume $p(u,x)>p(u,y)$
$w(u)\geq 0$ and
$w(u)\geq\tilde{w}(x)$
R8.2.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u,x\}$
and $w(y)\coloneqq w(y)-p(u,y)$.
$w(u)\geq 0$ and
$w(u)<\tilde{w}(x)$
R8.2.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$
$w(x)\coloneqq w(x)-w(u)$, $w(y)\coloneqq w(y)-p(u,y)$.
$w(u)<0$
R8.2.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$
$w(y)\coloneqq w(y)+w(u)-p(u,y)$.
Case 3: $\min(p(u,x),p(u,y))>w(u)$ w.l.o.g, assume $\tilde{w}(x)\geq\tilde{w}(y)$
$w(u)\geq 0$ and
$w(u)\geq\tilde{w}(x)$
R8.3.1: $\alpha(G)=\alpha(G\setminus\{u,x,y\})+w(u)$
$w(u)\geq 0$ and
$\tilde{w}(y)\leq w(u)<\tilde{w}(x)$
R8.3.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u,y\}$
$w(x)\coloneqq w(x)-w(u)$.
$w(u)\geq 0$
and $w(u)<\tilde{w}(y)$
R8.3.3: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$
$w(x)\coloneqq w(x)-w(u)$, $w(y)\coloneqq w(y)-w(u)$.
$w(u)<0$ R8.3.4: $\alpha(G)=\alpha(G\setminus\{u\})$.
Figure 2: Examples of R8. Suppose that vertex $u$ is adjacent to $x$ and $y$ and, $(x,y)\in E_{p}$. $G_{R}$ represents the residual graph $G\setminus\{u,x,y\}$. $\alpha$ and $\alpha^{\prime}$ indicate the optimal objective value in the original graph and the graph after reduction, respectively.
Examples of R8 are presented in Figure 2. For example, in the first plot, $u$ satisfies case 1 and $w(u)\geq 0$. Then by R8.1.1, we can remove $u$ from the original graph $G$ (left) to get the reduced graph $G^{\prime}$(right) and $\alpha(G)=\alpha(G^{\prime})+10$.
To describe the $R9$ reduction, we further define $p(u,v)=0$ for $(u,v)\notin E$. Note that this does not alter the objective value in the graph.
R 9 (Non-permanently Connected Neighbors Reduction).
Let $u$ be a vertex with two neighbors $x,y$ and $p(x,y)<M$. W.l.o.g, assume $p(u,x)\geq p(u,y)$. The following table summarizes the reduction rules under different conditions.
conditions reduction
Case 1: $w(u)\geq p(u,x)$
$w(u)\geq 0$ and
$w(u)\geq p(u,x)+p(u,y)$
R9.1.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
$w(x)\coloneqq w(x)-p(u,x)$, $w(y)\coloneqq w(y)-p(u,y)$
$w(u)\geq 0$ and
$w(u)<p(u,x)+p(u,y)$
R9.1.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(x)\coloneqq w(x)-p(u,x)$,
$w(y)\coloneqq w(y)-p(u,y)$, $p(x,y)\coloneqq p(x,y)-p(u,x)-p(u,y)+w(u)$
$w(u)<0$
R9.1.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(x)\coloneqq w(x)+w(u)-p(u,x)$,
$w(y)\coloneqq w(y)+w(u)-p(u,y)$, $p(x,y)\coloneqq p(x,y)+w(u)$.
Case 2: $p(u,x)>w(u)\geq p(u,y)$
$w(u)\geq 0$ and
$w(u)\geq p(u,x)+p(u,y)$
R9.2.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(y)\coloneqq w(y)-p(u,y)$,
$w(x)\coloneqq w(x)-w(u)$, $p(x,y)\coloneqq p(x,y)+p(u,x)-w(u)$.
$w(u)\geq 0$ and
$w(u)<p(u,x)+p(u,y)$
R9.2.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
$w(y)\coloneqq w(y)-p(u,y)$, $w(x)\coloneqq w(x)-w(u)$, $p(x,y)\coloneqq p(x,y)-p(u,y)$.
$w(u)<0$ and
$w(u)\geq p(u,x)+p(u,y)$
R9.2.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
$w(y)\coloneqq w(y)+w(u)-p(u,y)$, $p(x,y)\coloneqq p(x,y)+p(u,x)$.
$w(u)<0$ and
$w(u)<p(u,x)+p(u,y)$
R9.2.4: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
$w(y)\coloneqq w(y)+w(u)-p(u,y)$, $p(x,y)\coloneqq p(x,y)+w(u)-p(u,y)$.
Case 3: $p(u,y)>w(u)$ $w(u)\geq 0$
R9.3.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
$w(x)\coloneqq w(x)-w(u)$, $w(y)\coloneqq w(y)-w(u)$, $p(x,y)\coloneqq p(x,y)-w(u)$.
$w(u)<0$ and
$w(u)<p(u,x)+p(u,y)$
R9.3.2: $\alpha(G)=\alpha(G\setminus\{u\})$.
$w(u)<0$ and
$w(u)\geq p(u,x)+p(u,y)$
R9.3.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
$p(x,y)\coloneqq p(x,y)-w(u)+p(u,x)+p(u,y)$.
Figure 3: Examples of R9. Suppose that vertex $u$ is adjacent to vertices $x$ and $y$, and $(x,y)\notin E_{p}$.
Examples of reduction rule $R9$ are presented in Figure 3. As a matter of fact, $R8$ and $R9$ together can reduce all degree-two vertices. To be specific, R8 covers the cases where the two neighbors are connected by a permanent edge, while R9 covers the remaining cases where the two neighbors are connected by a removable edge or not connected.
4.3.3 Low Permanent-degree Reductions
We further introduce two reduction rules to reduce high-degree vertices that have only one or two permanent neighbors.
R 10 (Permanent Degree-one Vertices Reduction).
If there exists a vertex $u$ that has exactly one permanent neighbor $x$ and $w(u)\geq\sum_{v\in N_{r}(u)}max(0,p(u,v))$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(x)\coloneqq w(x)-w(u)$, $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$, and $p(x,v)\coloneqq p(x,v)-p(u,v)$ for each $v\in N_{r}(u)\setminus N_{p}(x)$.
R 11 (Permanent Degree-two Vertices Reduction).
Let $u$ be a vertex with exactly two permanent neighbors $x,y$, that $p(x,y)=M$, and $w(u)\geq\sum_{v\in N_{r}(u)}max(0,p(u,v))$. W.l.o.g, assume $\tilde{w}(x)\geq\tilde{w}(y)$.
Case 1: If $w(u)\geq\tilde{w}(x)$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u,x,y\}$, and $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$.
Case 2: If $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u,y\}$, $w(x)\coloneqq w(x)-w(u)$, $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$, and $p(x,v)\coloneqq p(x,v)-p(u,v)$ for each $v\in N_{r}(u)\setminus N_{p}(x)$.
Case 3: If $w(u)<\tilde{w}(y)$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(x)\coloneqq w(x)-w(u)$, $w(y)\coloneqq w(y)-w(u)$, $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$, $p(x,v)\coloneqq p(x,v)-p(u,v)$ for each $v\in N_{r}(u)\setminus N_{p}(x)$, and $p(y,v)\coloneqq p(y,v)-p(u,v)$ for each $v\in N_{r}(u)\setminus N_{p}(y)$.
4.4 Vertex-pair Reductions
We introduce three additional rules that can reduce some specific vertex pairs.
R 12 (Permanent Edge Reduction).
If there is a permanent edge $(u,v)$ such that $w(u)\geq\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}[v])+\min(w^{+}(N_{r}(u)\setminus N_{p}(v)),\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p(u,x)))$, then $\alpha(G)=\alpha(G\setminus\{v\})$.
R 13 (Common Neighbors Reduction).
If there is a permanent edge $(u,v)$ such that $w(v)\geq w^{+}(N(v))-\max(0,w(u))$, then $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G\setminus(N_{p}(u)\cap N_{p}(v))$.
R 14 (Twin Vertices Reduction).
If there exist two vertices $u,v$ such that $N_{p}(u)=N_{p}(v)$, $w(u)\geq\sum_{x\in N_{r}(u)}\max(0,p(u,x))$ and $w(v)\geq\sum_{x\in N_{r}(v)}\max(0,p(v,x))$, then $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}$ is obtained by folding $u$ and $v$ into a vertex $v^{\prime}$, $w(v^{\prime})\coloneqq w(u)+w(v)-p(u,v)$, $p(u,x)\coloneqq M$ for each $x\in N_{p}(u)$, and $p(v^{\prime},x)\coloneqq p(u,x)+p(v,x)$ for each $x\in N_{r}(v)\cup N_{r}(u)$.
5 Reduction-driven Local Search
Now, we elaborate on the local search algorithm driven by these 14 reduction rules for solving the GIS problem. This reduction-driven local search (RLS) algorithm is presented in Algorithm 1.
Input: a graph $G=(V,E)$ with vertex profits and edge penalties, cutoff time
Output: the best solution $S^{*}$ found
$S_{init}\leftarrow\varnothing$; /* initial an empty solution */
$G,S_{init}\leftarrow\textit{Complete\_Reduction}(G,S_{init})$; /* Pre-process to obtain a kernel graph via $R1$-$R14$/
1 while cutoff time is not satisfied do
       /* construct an initial solution using $R1$-$R14$/
2       $S\leftarrow S_{init}\cup\textit{Random\_Peeling}(G)$;
       /* improve $S$ via local search and partial reductions */
3       $S\leftarrow\textit{Neighborhood\_Search\_with\_Partial\_Reduction}(G,S)$;
       /* update the best visited solution */
4       if $nb(S)>nb(S^{*})$ then  $S^{*}\leftarrow S$ ;
5
6return $S^{*}$;
Algorithm 1 The reduction-driven local search for GIS
The RLS begins by pre-processing the original graph $G$. It uses the Complete_Reduction procedure which applies all the reduction rules to obtain a kernel graph. Due to the correctness of the reduction rules, solving the GIS problem in the original graph is equivalent to solving it in this kernel graph. The $S_{init}$ is a partial vertex set that must be in the optimal solution.
The remaining steps of RLS follow the traditional iterated local search algorithm (Glover and Laguna (1998), Lourenço et al. (2019)): At each iteration, it supplements $S_{init}$ with a random generalized independent set obtained by Random_Peeling, then uses Neighborhood_Search_with_Partial_Reduction to improve the current solution (line 1)-1) . The current solution is denoted by $S$ and the best-known solution visited by the iterative search is kept in $S^{*}$ (line 1). The iteration continues until the cutoff time limit is met. Compared with traditional local search, RLS seamlessly integrates reduction techniques into the pre-processing, the initial solution building, and the neighborhood searching procedures.
Input: a graph $G=(V,E)$ with vertex profits and edge penalties, initial solution $S\subseteq V$
Output: updated graph $G$ and partial solution $S$
/* initial candidate sets for 14 reduction rules */
1 $\text{Cand}_{i}\leftarrow E_{r}$ for $i=1,2$; $\text{Cand}_{i}\leftarrow V$ for $i=3,4,6$; $\text{Cand}_{5}\leftarrow\{v\in V\mid w(v)<0\}$;
2 $\text{Cand}_{7}\leftarrow\{v\in V\mid|N(v)|=1\}$; $\text{Cand}_{8}\leftarrow\{v\in V\mid|N(v)|=2\}$; $\text{Cand}_{9}\leftarrow\text{Cand}_{8}$;
3 $\text{Cand}_{10}\leftarrow\{v\in V\mid|N_{p}(v)|=1\}$; $\text{Cand}_{11}\leftarrow\{v\in V\mid|N_{p}(v)|=2\}$;
4 $\text{Cand}_{i}\leftarrow E_{p}$ for $i=12,13$; $\text{Cand}_{14}\leftarrow V\times V\setminus E_{p}$;
5 while $\exists i:\text{Cand}_{i}\neq\varnothing$ do
       /* exhaustively apply reduction rules */
6       for each $i=1,2,\ldots,14$ do
7             for each $elem\in\text{Cand}_{i}$ do
8                   if the condition of $Ri$ is satisfied by checking $elem$ then
9                         change the structure, profits and penalties of $G$ and set $S$ according to $Ri$;
10                         update the Candj if $Rj$ reachable from $Ri$ in Figure 4;
11
12                  Remove $elem$ from Candi;
return $G,S$;
Algorithm 2 The Complete_Reduction algorithm
5.1 Complete Reduction as Pre-processing
The Complete_Reduction algorithm exhaustively applies the 14 reduction rules in Section 4. A naive implementation is to check every rule independently, process the graph when applicable, and stop when no rules apply. The times of checking all these rules is $O(14(|V|+|E|))$ as it requires scanning the entire graph for each rule. Given the large size of real-world graphs, this is inefficient in practice. We propose a more efficient implementation which incrementally detects applicable rules. Specifically, we observe that rules are interdependent, with one application triggering others. For example, if only the condition of $R1$ holds in graph $G$ (i.e., there exists an edge $(u,v)$ with $p(u,v)=0$), the edge $(u,v)$ is removed, making R3-15 potentially applicable but not the other rules. Hence, we make an in-depth analysis of this dependency relationship for all these rules, and we show the result in Figure 4.
In Algorithm 2, we implement this reduction algorithm. It takes a graph $G$ as its input and outputs a kernel graph and a set $S\subseteq V$. In detail, it initializes a special candidate set Candi which contains either vertices, edges, or vertex pairs, for a rule $Ri$ where $i=1,...,14$. Lines 1-4 show how Candi is initialized. When an element $elem\in\text{Cand}_{i}$ satisfies the condition of $Ri$, the current graph $G$ is changed and the current set $S$ is updated according to $Ri$. Due to the change of graph $G$, if another reduction rule $Rj$ can be applied (i.e. $Rj$ is reachable from $Ri$ in the dependency graph in Figure 4), then the vertices, edges, or vertex pairs which are “neighbor elements” of $elem$ are moved into Candj. Consequently, the algorithm checks whether each reduction rule is satisfied by tracking the candidate set in an incremental manner.
Figure 4: The dependency graph of the reduction rules. A directed edge on the left side indicates how the reduction rule changes the graph $G$ to $G^{\prime}$. A directed edge on the right side indicates what reduction rules can be satisfied due to this change.
5.2 Random Peeling for Initial Solutions
Random-peeling is a simple random procedure for generating an initial solution. It starts with an empty vertex set $S$. Then when $V\setminus N_{p}(S)$ is not empty (meaning that $S$ can possibly be a larger generalized independent set), a vertex $u$ is picked randomly from $V\setminus N_{p}(S)$ and $u$ is moved into $S$. Due to the removal of $u$ from $G$, the Complete_Reduction is used again to further shrink the remaining graph $G$ (line 3). When $V\setminus N_{p}(S)$ becomes empty, we further extend $S$ by randomly picking a vertex $v$ such that $nb(S\cup\{v\})>nb(S)$. The $S$ is returned when it becomes a maximal generalized independent set.
Input: a graph $G=(V,E)$ with vertex profits and edge penalties
Output: a solution $S$ to the GIS problem
1 initialize an empty set $S$;
2 while $|V\setminus N_{p}[S]|>0$ do
3       $u\leftarrow$ randomly select a vertex from $V\setminus N_{p}[S]$;
4
5      $G,S\leftarrow\textit{Complete\_Reduction}(G\setminus\{u\},S)$;
6extend $S$ to be a maximal generalized independent set;
return $S$;
Algorithm 3 The Random_Peeling algorithm
5.3 Neighborhood Search
To improve the incumbent solution, we use the iterated tabu search with three move operators.
Move operators. In RLS, there are three move operators that update the current solution: ADD, SWAP, and the novel REDUCE operator based on two reduction rules. Given the current solution $S\subseteq V$, the ADD$(v)$ operator displaces a vertex $v$ (if any) from $V\setminus N_{p}[S]$ to the $S$ and the SWAP$(u,v)$ operator adds a vertex $v\in N_{p}(S)$ if $|N_{p}(v)\cap S|=1$ and removes the only vertex $u\in N_{p}(v)\cap S$ from the current solution $S$. Indeed, ADD and SWAP are often used in existing local search methods such as Zhou and Hao (2017), Jin and Hao (2015). The REDUCE$(V\setminus N_{p}[S])$ operator applies $R3$ and $R4$ to include candidate vertices that are definitely in a set in $MGIS(G)$.
Input: a graph $G=(V,E)$ with vertex profits and edge penalties, initial solution $S$, search depth $L\in\mathbb{N}$, perturbation parameter $\epsilon\in(0,1)$
Output: a solution $S$ to the GIS problem
1 $S^{*}\leftarrow S$, $tabu\_list\leftarrow\varnothing$;
2 while True do
3       $M_{1}\leftarrow V\setminus N_{p}[S]$, $M_{2}\leftarrow\{v\in N_{p}(S)\mid u\in S,(u,v)\in E_{p},|N_{p}(v)\cap S|=1\}\setminus tabu\_list$
4       $v\leftarrow$ a vertex with the largest move gain from $M_{1}\cup M_{2}$;
       if $v\in M_{1}$ then  $S\leftarrow S\cup\{v\}$, $M_{1}\leftarrow M_{1}\setminus\{v\}$ /* add $v$ into $S$/
5       else if $v\in M_{2}$ then
             $u\leftarrow N_{p}(v)\cap S$, $S\leftarrow S\cup\{v\}\setminus\{u\}$; /* Swap $v$ and $u$/
6             add $u$ into $tabu\_list$ with tabu tenure $T_{u}=10+random(0,|M_{2}|)$
       $REDUCE(M_{1})$ and reassign $S$; /* apply $R3$ and $R4$/
7       if $nb(S)>nb(S^{*})$ then  $S^{*}\leftarrow S$, $l\leftarrow 0$ else  $l\leftarrow l+1$
8      if $l>|S|$ or $M_{1}=\varnothing\&M_{2}=\varnothing$ then
             sequentially drop $\epsilon|S|$ vertices with the least move gain from $S$, ties breaking randomly; /* Perturbation */
9
10       if $l>L$ then break;
11
return $S^{*}$;
Algorithm 4 The Neighborhood_Search_with_Partial_Reduction algorithm
Fast Calculation of Move Gain. For each vertex $u\in V$, we maintain $B_{u}=w(u)-\sum_{v\in S,(u,v)\in E}p(u,v)$. Then, the move gain of $ADD(v)$, which is the gain of the net benefit after applying a $ADD(v)$, is $B_{v}$, and the move gain of $SWAP(u,v)$ is $B_{v}-B_{u}$. After an $ADD(v)$ move, for each $u\in N_{r}(v)$, we update $B_{u}$ as $B_{u}\coloneqq B_{u}-p(u,v)$ After a $SWAP(u,v)$ move, for each $x\in N_{r}(u)$, $B_{x}\coloneqq B_{x}+p(x,u)$ and for each $y\in N_{r}(v)$, $B_{y}\coloneqq B_{y}-p(y,v)$. The time complexity of performing $ADD$ or $SWAP$ is $O(\Delta)$, while the REDUCE operator operates exactly. The overall complexity of applying all three operators is $O(|E|)$. Notably, the REDUCE operator is not specific to the GIS problem and can enhance the efficiency of local search in other optimization problems.
Neighborhood Search with Tabu and Perturbation. The neighborhood search in RLS iteratively performs the three basic operators above to improve the current solution $S$, as shown in Algorithm 4.
At each iteration in lines 4-4, it first builds $M_{1}$, the set of vertices that can be added into $S$ by $ADD$, and $M_{2}$, the set of vertices that can be swapped into $S$ by $SWAP$. By evaluating $M_{1}$ and $M_{2}$, the $ADD$ or $SWAP$ with the largest move gain is applied, as presented in lines 4-8. After performing the ADD or SWAP operator, we additionally use the REDUCE operator to update $M_{1}$. This decreases the search space of the $ADD$ operator because $V\setminus N_{p}(S)$ is much larger than $S$ in practice.
To avoid short-term search cycles, a tabu list is used to exclude recently visited vertices for successive iterations (line 4). The tabu tenure for a vertex is set to $10+random(0,|M_{2}|)$, as commonly used in local search algorithms (Wu et al. 2023, Zhou and Hao 2017, Jin and Hao 2015). To further avoid local optima, a perturbation step is applied during neighborhood search (lines 4-4). Specifically, when we detect that no operators can be applied or the best solution remains unchanged after a certain number of iterations, we remove a portion of low-contribution vertices. The whole neighborhood search terminates when the consecutive number of iterations without improvement reaches the search depth parameter $L$ (line 4).
6 Computational Experiments
This section is dedicated to a comprehensive evaluation of the proposed reduction rules and heuristic approaches. In order to evaluate the algorithm performance, we compare both solution quality and runtime between RLS and other benchmark algorithms; to evaluate the effect of reduction rules, we compare the algorithms with reductions and without reductions; we also conduct an ablation study to evaluate the effect of each reduction rule.
6.1 Datasets
The proposed reduction rules and the RLS algorithm are evaluated on three sets and a total of $278=216+47+15$ instances, covering various application scenarios.
SET-1. The first set of instances consists of $216$ standard instances with up to $400$ vertices and up to $71,820$ edges. These instances were first introduced by Colombi et al. (2017) and tested in Nogueira et al. (2021), Hosseinian and Butenko (2019), Zheng et al. (2024). This set of instances consists of 12 graphs with 125 to 400 vertices and 6963 to 71820 edges (available at https://or-dii.unibs.it/instances/Instances_GISP.zip).
SET-2. The second set of instances consists of $47$ medium-sized instances with up to $17,903$ vertices and up to $1,997,732$ edges, introduced by Zheng et al. (2024) as benchmark instances for the GIS problem (available at https://github.com/m2-Zheng/GISP/tree/main).
SET-3. To further evaluate the algorithms on large instances and diverse real-world scenarios, we introduce 15 instances with up to 18 million vertices and 261 million edges. The first four smaller instances represent ecosystem and road/facility networks, while the remaining 11 larger instances are drawn from hyperlink, citation, and social networks. The instances were generated using the same method for generating SET-2 (available at https://github.com/PlutoAiyi/RLS/tree/main).
6.2 Experimental Settings
All experiments are conducted on a machine with an Intel(R) Xeon(R) Platinum 8360Y CPU @ 2.40GHz and 2T RAM.
Benchmark Algorithms. To evaluate the effectiveness of our proposed reduction rules and RLS algorithm, we compare them with the best-performing algorithms for the GIS problem.
  • CB&B: The exact branch-and-bound algorithm using the UBQP formulation proposed by Hosseinian and Butenko (2019).
  • LA-B&B: The exact branch-and-bound algorithm using an Lagrange relaxation techniques by Zheng et al. (2024).
  • ILS-VND: The iterative local search algorithm proposed by Nogueira et al. (2021).
  • ALS: The adaptive multi-neighborhood local search algorithm proposed by Zheng et al. (2024)
  • RLS: Our proposed reduction-driven local search algorithm. The algorithm is implemented in C++.
The source codes of ILS-VND, LA-B&B and ALS were made available by the authors. All source codes are available at https://github.com/PlutoAiyi/RLS.
Parameter Settings. The RLS algorithm requires two parameters: the search depth $L$ and the perturbation parameter $\epsilon\in(0,1)$. We set $L=10|V|$ where $|V|$ is the number of vertices of the input graph and $\epsilon=0.2$. This setting is obtained using the general IRACE automatic parameter configuration tool (López-Ibáñez et al. 2016). For large instances like soc-pokec, LiveJ and uk2002, we decrease the search depth as $L=0.5|V|$ to encourage explorations within the cutoff time.
6.3 Performance Comparison with Existing Algorithms
We first compared five algorithms on the SET-1 instances. The exact algorithms CB&B and LA-B&B ran with a cutoff time of 3 hours (10800 seconds) for each instance. If an exact algorithm did not finish within the cutoff time, then we reported the best solution found. The remaining three heuristic algorithms ran 10 trials with a cutoff time of 30 seconds for each instance. We reported the largest value among the best-known objective values obtained in these trials. The setting is consistent with that in (Hosseinian and Butenko 2019, Nogueira et al. 2021).
In Table 2, we summarize the computational results on SET-1. In the last two columns, we show the number of instances where the corresponding algorithm performs best (#Best) and worst (#Worst) among all the five algorithms, in terms of the best known objective value and runtime, respectively. Note that the runtime of RLS refers to the wall-clock time, including both pre-processing and local search. We observe that, in all instances, the best-known objective values obtained by local search algorithms are comparable to those obtained by exact algorithms. In particular, if an exact algorithm achieves the optimal solution within 3 hours, local search algorithms can also find an optimal solution (although without proof). However, when neither of the exact algorithms finds the optimal solution, local search algorithms always produce a better one. Besides, the exact LA-B&B dominates CB&B whereas the heuristic ALS and RLS dominate ILS-VND in terms of both solution quality and runtime. Therefore, we only consider LA-B&B, ALS, and RLS for even larger instances in SET-2 and SET-3. Detailed results on this set are left in Table LABEL:tab:_Set1_detail in Appendix B.
Table 2: Summary of comparative results on SET-1 instances
Algorithm Indicator: net benefit Indicator: time(s)
#Total #Best #Worst #Best #Worst
Exact CB&B 216 126 88 78 138
LA-B&B 216 155 3 216 0
Local Search ILS-VND 216 215 0 164 52
ALS 216 216 0 216 0
RLS 216 216 0 216 0
Table 3: Comparative results of RLS on SET-2 instances
Instance $|V|$ $|E_{p}|$ $|E_{r}|$ LA-B&B ALS RLS
[…]
bio-yeast 1,458 980 968 66,545 300 68,572 9 68,574 $<$1 0 $<$1
soc-wiki-Vote 889 727 2,187 35,973 300 37,358 8 37,358 $<$1 0 $<$1
vc-exact_131 2,980 1,191 4,169 107,812 300 109,392 14 109,641 $<$1 0 $<$1
web-edu 3,031 4,813 1,661 109,856 300 113,781 16 114,140 $<$1 98 $<$1
MANN_a81 3,321 3,118 3,362 114,616 300 116,907 14 117,327 $<$1 0 $<$1
tech-routers-rf 2,113 1,539 5,093 93,690 300 97,799 11 97,799 $<$1 0 $<$1
vc-exact_039 6,795 2,574 8,046 254,941 300 258,137 9 259,542 $<$1 0 $<$1
ca-GrQc 4,158 3,179 10,243 167,846 300 174,654 15 174,782 11 282 $<$1
vc-exact_038 786 10,321 3,703 10,371 300 12,014 12 12,015 4 697 $<$1
vc-exact_078 11,349 4289 13,450 102,312 300 435,412 $<$1 438,671 $<$1 0 $<$1
vc-exact_087 13,590 5126 16,114 71,804 300 517,500 2 518,500 $<$1 0 $<$1
vc-exact_107 6,402 15,646 5,594 88,602 300 400,223 1 407,878 $<$1 0 $<$1
vc-exact_151 15,783 12,144 12,519 54,114 300 522,462 1 530,138 $<$1 0 $<$1
vc-exact_167 15,783 18,235 6,428 55,190 300 457,272 2 465,178 $<$1 0 $<$1
bio-dmela 7,393 18,897 6,672 280,897 300 302,449 10 302,992 $<$1 44 $<$1
vc-exact_011 9,877 19,102 6,871 190,729 300 301,789 7 303,793 $<$1 172 $<$1
web-spam 4,767 8,089 28,386 180,446 300 190,344 14 190,459 13 487 $<$1
vc-exact_026 6,140 27,159 9,608 177,898 300 201,316 17 201,503 $<$1 300 $<$1
vc-exact_001 6,160 19,655 20,552 190,729 300 215,668 13 215,941 $<$1 134 $<$1
vc-exact_024 7,620 11,367 35,926 226,540 300 235,440 14 236,018 12 1,255 $<$1
C1000.9 1,000 11,989 37,432 6,102 300 7,229 12 7,237 10 1,000 $<$1
tech-WHOIS 7,476 42,181 14,762 309,903 300 325,327 14 325,581 13 338 $<$1
p-hat700-3 700 30,398 31,242 2,937 300 4,473 $<$1 4,473 $<$1 700 $<$1
vc-exact_008 7,537 35,678 37,155 71,804 300 246,771 14 247,091 $<$1 574 $<$1
keller5 776 17,800 56,910 2,788 300 3,877 1 3,877 2 776 $<$1
vc-exact_194 1,150 39,520 41,331 4,273 300 5,419 $<$1 5,419 $<$1 1,150 $<$1
hamming10-4 1,024 66,369 23,231 2,871 300 3,660 $<$1 3,660 $<$1 1,024 $<$1
brock800_2 800 26,629 84,805 1,996 300 2,473 $<$1 2,473 $<$1 800 $<$1
brock800_4 800 26,831 85,126 1,968 300 2,501 $<$1 2,501 $<$1 800 $<$1
ca-HepPh 11,204 57,721 59,898 153,907 300 370,779 6 372,079 17 2,165 $<$1
p-hat700-2 700 90,788 32,134 2127 300 2,971 $<$1 2,971 $<$1 700 $<$1
vc-exact_196 1534 93,273 32,809 4273 300 5,321 $<$1 5,321 $<$1 1,534 $<$1
p-hat700-1 700 135,804 47,847 844 124 844 $<$1 844 $<$1 700 $<$1
ca-AstroPh 17,903 47,135 149,837 44,296 300 602,320 2 603,945 14 4,985 1
C2000.9 2,000 147,548 51,920 5477 300 6396 3 6396 3 2,000 $<$1
DSJC1000.5 1,000 59,694 189,980 1492 300 1861 1 1861 $<$1 1,000 $<$1
socfb-MIT 6,402 123,064 128,166 110,677 300 129,807 17 129,872 16 5,494 $<$1
p-hat1500-3 1,500 66,206 210,800 3856 300 7,501 14 7,501 16 1,500 $<$1
socfb-UCSB37 14,917 356,918 125,297 219,965 300 254,965 16 255,898 22 13,655 3
socfb-Duke14 9,885 374,731 131,706 125,940 300 151,169 15 151,326 18 9,003 4
p-hat1500-2 1,500 411,372 143,918 2,444 300 4,253 $<$1 4,253 $<$1 1,500 $<$1
socfb-Stanford3 11,586 136,823 431,486 215,381 300 251,350 12 251,833 14 9,380 2
socfb-UConn 17,206 448,038 156,829 170,548 300 274,637 12 275,034 20 16,025 5
p-hat1500-1 1,500 411,744 427,583 1,121 300 1,287 $<$1 1,287 1 1,500 $<$1
C2000.5 2,000 489,894 509,270 1,570 300 1,849 2 1,849 6 2,000 1
keller6 3,361 759,816 266,766 3,545 300 5,236 10 5,236 16 3,361 $<$1
C4000.5 4,000 1,957,884 2,039,848 1,601 300 1,948 20 1,950 22 4,000 9
# of best / # of instances 1/47 18/47 47/47
average running time(s) 296.3 7.9 5.9 1.4
Table 4: Comparative results of RLS on SET-3 instances
Instance $|V|$ $|E_{p}|$ $|E_{r}|$ LA-B&B ALS RLS
[…]
foodweb-wet 128 1,509 566 2,464 54.84 2,464 $<$1 2,464 $<$1 125 $<$1
foodweb-dry 128 1,571 535 2,226 27.87 2,226 $<$1 2,226 $<$1 125 $<$1
USAir97 332 1,589 537 9,507 2000 10,308 $<$1 10,308 $<$1 0 $<$1
powergrid 4,941 1,727 4,867 190,537 2000 193,095 59 193,339 $<$1 0 $<$1
CondMat 23,133 46,769 46,670 95,130 2000 705,901 14 711,334 45 1,648 $<$1
Email 265,009 91,093 273,388 OOM OOM 13,054,427 1 0 1
Epinion 75,879 202,653 203,087 7,400 2000 3,038,533 18 3,052,059 1 253 1
Dblp 317,080 787,268 262,598 OOM OOM 9,734,653 64 11,851 11
cnr-2000 325,557 685,686 2,053,283 OOM OOM 13,368,670 482 59,052 399
WikiTalk 2,394,385 3,494,873 1,164,692 OOM OOM 118,708,326 333 8 333
BerkStan 685,230 1,662,833 4,986,637 OOM OOM 26,031,704 1112 114,191 1031
As-Skitter 1,696,415 5,550,661 5,544,637 OOM OOM 66,582,254 723 31,173 636
soc-pokec 1,632,803 11,152,945 11,149,019 OOM OOM 43,999,402 934 881,319 134
LiveJ 4,846,609 10,710,579 32,140,658 OOM OOM 168,431,056 904 401,307 469
uk2002 18,483,186 65,451,024 196,336,234 OOM OOM 687,089,901 7678 4,806,660 6243
# of best / # of instances 3/15 3/15 15/15
average running time(s) 1347.1 15.7 818.7 617.5
[…]
The comparative results on SET-2 and SET-3 are presented in Tables 3 and 4. The first four columns list the instance information. The remaining columns show the best-known objective value (BKV), the running time of the exact algorithm (time), the mean running time that the heuristic algorithm first hits the BKV (htime), the kernel graph size ($|V_{\text{ker}}|$) and the time of generating a ${}_{\text{ker}}$). In the last two rows, we show the statistical results. Note that the htime of RLS has included the pre-processing time for building graph kernels.
Table 3 shows that RLS outperforms other methods on all SET-2 instances in terms of both best-known objective value and runtime. It is known that RLS proves optimality if all vertices are reduced (i.e. $|V_{kernel}|=0$), hence 11 instances have been solved to optimality by RLS. In contrast, LA-B&B can only solve 1 out of these 47 instances. Moreover, RLS and ALS hit the BKV in just a few seconds.
Table 5: Comparative results of algorithms with or without reduction rules
Instance $|V|$ $|V_{ker}|$ LA-B&B RLS
without reduction with reduction without reduction with reduction
BKV time(s) BKV time(s) BKV htime(s) BKV htime(s)
bio-yeast 1,458 0 66,545 300 68,574 0 68,574 1 68,574 $<$1
soc-wiki-Vote 889 0 35,973 300 37,358 0 37,358 1 37,358 $<$1
vc-exact_131 2,980 0 107,812 300 109,641 0 109,567 9 109,641 $<$1
web-edu 3,031 98 109,856 300 113,781 15 114,080 12 114,140 $<$1
MANN_a81 3,321 0 114,616 300 117,327 0 117,275 14 117,327 $<$1
tech-routers-rf 2,113 0 93,690 300 97,799 0 97,799 1 97,799 $<$1
vc-exact_039 6,795 0 254,941 300 259,542 0 259,475 13 259,542 $<$1
ca-GrQc 4,158 282 167,846 300 174,654 15 174,768 7 174,782 11
vc-exact_038 786 697 10,371 300 12,014 12 12,015 13 12,015 4
vc-exact_078 11,349 0 102,312 300 438,671 1 438,306 16 438,671 $<$1
vc-exact_087 13,590 0 71,804 300 518,500 0 518,069 15 518,500 $<$1
vc-exact_107 6,402 0 88,602 300 407,878 0 406,379 14 407,878 $<$1
vc-exact_151 15,783 0 54,114 300 530,138 0 529,003 18 530,138 $<$1
vc-exact_167 15,783 0 55,190 300 465,178 0 463,446 12 465,178 $<$1
bio-dmela 7,393 44 280,897 300 302,449 10 302,793 18 302,992 $<$1
vc-exact_011 9,877 172 190,729 300 301789 7 303,352 9 303,793 $<$1
web-spam 4,767 487 180,446 300 190,344 14 190,460 18 190,459 13
vc-exact_026 6,140 300 177,898 300 201,316 17 201,458 13 201,503 $<$1
vc-exact_001 6,160 134 190,729 300 215,668 13 215,934 13 215,941 $<$1
vc-exact_024 7,620 1,255 226,540 300 235,440 14 235,911 14 236,018 12
C1000.9 1,000 1,000 6,102 300 6,102 300 7,235 16 7,237 10
tech-WHOIS 7,476 338 309,903 300 325,327 15 325,480 16 325,581 13
p-hat700-3 700 700 2,937 300 2,937 300 4,473 1 4,473 $<$1
vc-exact_008 7,537 574 71,804 300 246,771 14 247,036 16 247,091 $<$1
keller5 776 776 2,788 300 2,788 300 3,877 2 3,877 2
vc-exact_194 1,150 1,150 4,273 300 4,273 300 5419 2 5,419 $<$1
hamming10-4 1,024 1,024 2,871 300 2,871 300 3660 1 3660 $<$1
brock800_2 800 800 1,996 300 1,996 300 2473 1 2473 $<$1
brock800_4 800 800 1,968 300 1,968 300 2501 1 2501 $<$1
ca-HepPh 11,204 2,165 153,907 300 372,079 6 372,079 14 372,079 17
p-hat700-2 700 700 2,127 300 2,127 300 2,971 1 2,971 $<$1
vc-exact_196 1,534 1,534 4,273 300 4,273 300 5,321 2 5,321 $<$1
p-hat700-1 700 700 844 124 844 124 844 1 844 $<$1
ca-AstroPh 17,903 4,985 44,296 300 602,320 2 603,732 15 603,945 14
C2000.9 2,000 2,000 5,477 300 5,477 300 6,396 3 6,396 3
DSJC1000.5 1,000 1,000 1,492 300 1,492 300 1,861 1 1,861 $<$1
socfb-MIT 6,402 5,494 110,677 300 129,807 17 129,831 16 129,872 16
p-hat1500-3 1,500 1,500 3,856 300 3,856 300 7,500 15 7,501 16
socfb-UCSB37 14,917 13,655 219,965 300 254,965 16 255,887 21 255,898 22
socfb-Duke14 9,885 9,003 125,940 300 151,169 15 151,321 14 151,326 18
p-hat1500-2 1,500 1,500 2444 300 2,444 300 4,253 1 4,253 $<$1
socfb-Stanford3 11,586 9,380 215,381 300 251,350 12 251,788 19 251,833 14
socfb-UConn 17,206 16,025 170,548 300 274,637 12 275,006 16 275,034 20
p-hat1500-1 1500 1,500 1,121 300 1,121 300 1,287 2 1,287 1
C2000.5 2,000 2,000 1,570 300 1,570 300 1,849 5 1,849 6
keller6 3,361 3,361 3,545 300 3,545 300 5,229 19 5,236 16
C4000.5 4,000 4,000 1,601 300 1,601 300 1,951 15 1,951 22
foodweb-wet 128 125 2,464 55 2,464 27 2,464 $<$1 2,464 $<$1
foodweb-dry 128 125 2,226 28 2,226 25 2,226 $<$1 2,226 $<$1
USAir97 332 0 9,507 2000 10,308 0 10,308 $<$1 10,308 $<$1
powergrid 4,941 0 190,537 2000 193,339 0 193,194 52 193,339 $<$1
CondMat 23,133 1,648 95,130 2000 708,907 2000 709,927 54 711,334 45
Email 265,009 0 OOM 13,054,427 1 13,054,390 69 13,054,427 1
Epinion 75,879 253 7,400 2000 3,051,256 2000 3,050,415 67 3,052,059 1
Dblp 317,080 11,851 OOM 9,715,860 2000 9,577,609 46 9,734,653 64
cnr-2000 325,557 59,052 OOM 11,347,241 2000 13,361,167 84 13,368,670 482
WikiTalk 2,394,385 8 OOM 118,708,326 $<$1 118,672,055 56 118,708,326 333
BerkStan 685,230 114,191 OOM 22,591,394 2000 25,813,846 2000 26,031,704 1112
As-Skitter 1,696,415 31,173 OOM 65,907,243 2000 59,886,304 2000 66,582,254 723
soc-pokec 1,632,803 881,319 OOM OOM 36,540,838 350 43,999,402 934
LiveJ 4,846,609 401,307 OOM OOM 148,758,359 318 168,431,056 904
uk2002 18,483,186 4,806,660 OOM OOM 633,365,020 1806 687,089,901 7678
# of best / # of instances 3/62 24/62 24/62 62/62
average running time(s) 415.2 296.7 119.4 202.6
Table 4 compares algorithms on SET-3 instances. OOM indicates that the algorithm runs out of memory. Both ALS and LA-B&B quickly become infeasible as instance size increases. In contrast, RLS computes feasible solutions even for large instances such as uk2002 . For instances where the benchmarks are feasible, RLS consistently achieves the best net benefit. For the infeasible instance Email, RLS even finds the optimal solution.
6.4 Evaluation of Reduction Rules
To evaluate the effectiveness of the reduction rules, we conduct the following two experiments.
Comparative results with vs. without reduction. This experiment evaluates algorithms with and without data reduction. For LA-B&B with reduction, we use kernel graphs as input. For RLS without reduction, all reductions are skipped, and random initialization is applied. ALS cannot handle kernel graphs due to unsupported negative vertex profits and edge penalties.
The comparative results are shown in Table 5. In general, the reduction rules can help both algorithms to find better solutions. Specifically, with the reduction technique, LA-B&B is able to solve some instances which it cannot solve without reduction, e.g. WikiTalk (1.6 million vertices, 11.1 million edges). And for some large instances that LA-B&B fails due to out-of-memory, it is able to report a feasible solution or even solve them. As for RLS, with these reduction rules, it consistently finds better results across all instances in a shorter time than without them.
Figure 5: The variation of kernel sizes when ablating different reductions. A cell value indicates the difference in kernel size between without using reductions and with using reductions. A higher cell value indicates a larger effect.
Figure 6: The variation of pre-processing time for getting a kernel. A cell value indicates the difference in processing time(s) between without using the corresponding reductions and with using the corresponding reductions. A higher cell value indicates a larger effect.
Ablation study of reduction rules. In this experiment, we compare kernel sizes (measured by the number of vertices in the graph) and the pre-processing time (i.e., the runtime of Algorithm 2) after ablating different reduction rules. Normally, a smaller kernel indicates a smaller search space. Experimental results are shown in Figures 5 and 6.
Figure 5 shows that the impacts on the kernel size vary among these reduction rules. When comparing individual reductions, $R10$ and $R12$ are more powerful than others in decreasing the kernel size. When comparing a group of reductions, the degree-based rules ($R7$$R11$) are the most powerful in decreasing the kernel size. The neighborhood-based rules ($R3$$R6$) are generally more effective than vertex-pair reduction rules ($R12$$R14$) except for Dblp and WikiTalk. Additionally, although $R1$ and $R2$ do not involve vertex removals, they are also powerful in decreasing the kernel size. In particular, incorporating $R2$ into RLS can further reduce more than 3000 vertices for the BerkStan instance.
It is intuitive that ablating a reduction rule can decrease the overall pre-processing time, but may at the cost of resulting in a larger kernel. However, Figure 6 shows that ablating reduction rules $R3$$R6$ or $R7$$R11$ can even increase the pre-processing time. This is due to the fact that other reductions cost more time when these reduction rules are absent. We also observe that for the larger four instances As-Skitter, Berkstan, WikiTalk and cnr-2000, the inclusion of reductions shows a significant improvement over the pre-processing time.
7 Conclusion and Future Works
In this paper, we proposed a novel approach to the Generalized Independent Set (GIS) problem using data reduction techniques to tackle the complexity of large graphs. We developed 14 reduction rules, categorized into edge transformation, neighborhood-based, low-degree, and vertex-pair reductions, and designed a reduction-driven local search (RLS) algorithm based on them. Experiments on benchmark instances showed that the RLS algorithm outperformed state-of-the-art methods, especially on large graphs with tens of thousands of vertices. Future work includes exploring advanced reduction techniques for improved efficiency and adapting our algorithms to more real-world scenarios like dynamic graphs.
Appendix
A Technical Proofs
Proof of R1.
Consider whether $u$ and $v$ are included in a generalized independent set $S$. There are two cases: either $u\notin S$ or $v\notin S$, and $\{u,v\}\subseteq S$. When either $u\notin S$ or $v\notin S$, the edge $(u,v)$ does not appear in $S$, and thus removing $(u,v)$ has no effect on the net benefit $nb(S)$. If both $u\in S$ and $v\in S$, the edge $(u,v)$ has a zero penalty, which also has no effect on the net benefit $nb(S)$. ∎
Proof of R2.
Assume, for the sake of contradiction, that there exists $I\in MGIS(G)$ that $u\in I$ and $v\in I$. We have $nb(I)-nb(I\setminus\{u\})=w(u)-\sum_{x\in N_{r}(u)\cap S}p(u,x)\leq w(u)-p(u,v)+\sum_{x\in N_{r}(u)}max(0,-p(u,x))=\tilde{w}(u)-p(u,v)<0$. In conclusion, $I\setminus\{u\}$ has a larger net benefit, which contradicts the optimality of $I$. Therefore, $u$ and $v$ cannot be simultaneously included in any $S\in MGIS(G)$, which is equivalent to the case where $p(u,v)=M$. ∎
The proof of R3 directly follows from the Reduction 2 of Lamm et al. (2019).
Proof of R4.
We will prove that $u$ is in some set in $MGIS(G)$. Assume, for the sake of contradiction, that $u$ is not included in any set in $MGIS(G)$. Let $I$ be an arbitrary set in $MGIS(G)$. Let $S:=N_{p}(u)\cap I$ and $S^{\prime}:=N_{r}(u)\cap I$. Now, we remove $S$ from $I$ and add $u$ to $I$ to obtain $I^{\prime}$. Consider the difference between two net benefits $nb(I^{\prime})-nb(I)\geq w(u)-\sum_{v\in S}\tilde{w}(v)-\sum_{v\in S^{\prime}}p(u,v)\geq 0$. Therefore, $I^{\prime}\in MGIS(G)$, which contradicts the assumption. So $u$ is in at least one set in $MGIS(G)$. And it holds that $\alpha(G)=\alpha(G^{\prime})+w(u)$. ∎
Proof of R5.
We prove that $u$ is not in any set in $MGIS(G)$. Given an arbitrary $I\in MGIS(G)$, there are two cases. When none of the neighbors of $u$ is included in $I$, $I$ does not include $u$ because $w(u)<0$. When any subset of neighbors of $u$ is included in $I$, $I$ also does not include $u$ because $w(u)-\sum_{v\in S}p(u,v)<0$ for any subset $S\subseteq N(u)$. Since $I$ is an arbitrary set in $MGIS(G)$, $u$ will not appear in any $I\in MGIS(G)$ and we can directly remove $u$ from $G$. ∎
Proof of R6.
Observe that there exists a vertex set $I\in MGIS(G)$ that $I\cap N_{p}[u]\neq\varnothing$, because if $I\cap N_{p}[u]=\varnothing$ then $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$ implies $I\cup\{u\}\in MGIS(G)$.
Whenever a vertex $x$ in $N_{p}(u)$ is included in $I$, other vertices in the clique $N_{p}[u]\setminus\{x\}$ must not be included in $I$. Then $I\cup\{u\}\setminus\{x\}\in MGIS(G)$ because $nb(I\cup\{u\}\setminus\{x\})\geq nb(I)+w(u)-w(x)+\sum_{v\in N_{r}(x)}p(v,x)-\sum_{v\in N_{r}(u)}max(0,p(u,v))\geq nb(I)+w(u)-\sum_{v\in N_{r}(u)}max(0,p(u,v))-\max(0,\tilde{w}(x))\geq nb(I)$. In conclusion, $u$ is in some set in $MGIS(G)$. ∎
Proof of R7.
Given an arbitrary $I\in MGIS(G)$.
When $w(u)\geq p(u,v)$ and $w(u)\geq 0$. If $u\notin I$ and $v\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})=nb(I)+w(u)-p(u,v)\geq nb(I)$. If $u\notin I$ and $v\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$. In summary, $u$ is included in some set in $MGIS(G)$.
When $w(u)<0$ and $w(u)\geq p(u,v)$. If $v\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})=nb(I)+w(u)-p(u,v)\geq nb(I)$ for $u\notin I$. If $v\notin I$, then $nb(I\cup\{u\})<nb(I\setminus\{u\})$. In other word, there exists a set $S\in MGIS(G)$ that $u\in S$ if and only if there exists a set $S^{\prime}\in MGIS(G^{\prime})$ that $v\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,v$ that could appear in a set in $MGIS(G)$.
When $w(u)\geq 0$, $w(u)<p(u,v)$ and $w(u)\geq\tilde{w}(v)$. If $u\notin I,v\in I$, then $nb(I\cup\{u\})\setminus\{v\})-nb(I)=w(u)-w(v)-\sum_{x\in N_{r}(v)\cap I}max(0,-p(v,x))=w(u)-\tilde{w}(v)\geq 0$, thus $I\cup\{u\}\in MGIS(G)$. Otherwise, if $u\notin I,v\notin I$, then $nb(I\cup\{u\})-nb(I)=w(u)\geq 0$, thus $I\cup\{u\}\in MGIS(G)$. In summary, $u$ is included in some vertex set $I^{\prime}\in MGIS(G)$ and $v$ is not in $I^{\prime}$.
When $w(u)\geq 0$, $w(u)<p(u,v)$ and $w(u)<\tilde{w}(v)$. $u$ and $v$ cannot be both included in $I$. Because if $u,v\in I$, then $nb(I\setminus\{u\})=nb(I)+p(u,v)-w(u)>nb(I)$ contradicts to the optimality of $I$. Furthermore, if $v\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})-nb(I\setminus\{u\})=w(u)\geq 0$. We conclude that there exists a vertex set in $MGIS(G)$ that includes $u$ iff there exists a vertex set in $MGIS(G^{\prime})$ that does not contain $v$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,v$ that could appear in a set in $MGIS(G)$.
Finally, when $w(u)<0$ and $w(u)<p(u,v)$. We have $nb(I\cup\{u\})<nb(I\setminus\{u\})$, therefore $u$ is not in any set in $MGIS(G)$ and we can directly remove $u$. ∎
Proof of R8.
Given an arbitrary $I\in MGIS(G)$. Because that $p(x,y)=M$, it holds that $x$ and $y$ cannot be both included in $I$.
Case 1: $w(u)\geq\max(p(u,x),p(u,y))$.
When $w(u)\geq 0$. If both $x$ and $y$ are not included in $I$ and $u\notin I$, then $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$. If $u\notin I$ and either $x\in I$ or $y\in I$, $nb(I\cup\{u\})\geq nb(I)+w(u)-\max(p(u,x),p(u,y))\geq nb(I)$. In summary, there exists a vertex set in $MGIS(G)$ that includes $u$.
When $w(u)<0$. If both $x$ and $y$ are not included in $I$, then $u\notin I$ because $nb(I\cup\{u\})=nb(I\setminus\{u\})+w(u)<nb(I\setminus\{u\})$. If either $x\in I$ or $y\in I$ and $u\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})\geq nb(I)+w(u)-\max(p(u,x),p(u,y))\geq nb(I)$. We conclude that there exists a vertex set in $MGIS(G)$ that includes $u$ iff there exists a vertex set in $MGIS(G^{\prime})$ that includes either $x$ or $y$. In other word, we can defer the decision of $u$ after determining $x$ and $y$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
Case 2: $p(u,y)\leq w(u)<p(u,x)$. In this case, if $y$ is in $I$, then $x\notin I$ and $I\cup\{u\}\in MGIS(G)$.
When $w(u)\geq 0$ and $w(u)\geq\tilde{w}(x)$. If $x$ is in $I$, then $y,u$ are not in $I$ and $I\cup\{u\}\setminus\{x\}\in MGIS(G)$ because $nb(I\cup\{u\}\setminus\{x\})\geq nb(I)+w(u)-\tilde{w}(x)\geq nb(I)$. If neither $x$ nor $y$ is included in $I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$ for $u\notin I$. In summary, there exists a vertex set $I^{\prime}\in MGIS(G)$ that includes $u$ and not includes $x$.
When $w(u)\geq 0$ and $w(u)<\tilde{w}(x)$. There are five cases: $x\in I$ and $u,y\notin I$, $y\in I$ and $u,x\notin I$, $u\in I$ and $x,y\notin I$, $u,y\in I$ and $x\notin I$, and $u,x\in I$ and $y\notin I$. But $u,x$ cannot appear in same $I$, because otherwise $nb(I\setminus\{u\})=nb(I)+p(u,x)-w(u)>0$ contradicts the optimality of $I$. And when $y\in I$ and $u,x\notin I$, then $I\cup\{u\}\in MGIS(G)$. In summary, there exists a vertex set $I^{\prime}\in MGIS(G)$ such that either $x\in I^{\prime}$, or both $u,y\in I^{\prime}$, or $u\in I^{\prime}$. Since $u$ only connects to $x$ and $y$ in $G$, we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that does not include $x$. In other word, we can defer $u$ after the decision of $x$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
When $w(u)<0$. If $y\notin I$, then $u\notin I$ because otherwise $nb(I\setminus\{u\})>nb(I)$. Since $u$ only connects to $x$ and $y$ in $G$, we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that includes $y$. In other word, we can defer $u$ after the decision of $y$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
Case 3: $w(u)<p(u,x)$ and $w(u)<p(u,y)$. When $w(u)\geq 0$, the three vertices form a triangle with permanent edges. The result follows from theorem 3.3 of Gu et al. (2021). In particular, (1) if $w(u)\geq 0$ and $w(u)\geq\tilde{w}(x)$, then $u$ must be in some set in $MGIS(G)$. (2) If $w(u)\geq 0$ and $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$, then $y$ must not be in some set in $MGIS(G)$. And there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that does not include $x$. (3) If $w(u)\geq 0$ and $w(u)<\tilde{w}(y)$, then there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that does not include both $x$ and $y$. (4) When $w(u)<0$, the result follows from R 5. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all conditions and all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$. ∎
Proof of R9.
Given an arbitrary $I\in MGIS(G)$. Notice that $x$ and $y$ can be both included in $I$.
Case 1: $w(u)\geq p(u,x)$. In this case, if one of $x,y$ is in $I$, say $x$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})-nb(I)\geq w(u)-p(u,x)\geq 0$ when $u\notin I$.
When $w(u)\geq 0$ and $w(u)\geq p(u,x)+p(u,y)$. If either $x\in I$ or $y\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})-nb(I)\geq w(u)-\max(p(u,x)+p(u,y),p(u,x))\geq 0$ for $u\notin I$. If none of $x,y$ is in $I$, then $I\cup\{u\}\in MGIS(G)$ because $w(u)\geq 0$. Therefore, $u$ must belong to some set in $MGIS$.
When $w(u)\geq 0$ and $w(u)<p(u,x)+p(u,y)$. If both $x,y$ are in $I$, then $nb(I\cup\{u\})<nb(I)$. If either $x\notin I$ or $y\notin I$, then $nb(I\cup\{u\})\geq nb(I)$. In other word, including $u$ has negative profit only if both $x$ and $y$ are included in the $MGIS$. we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that either $x\notin S^{\prime}$ or $y\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
When $w(u)<0$, it holds that $w(u)\geq p(u,x)+p(u,y)$ because $w(u)\geq p(u,x)$ and $p(u,y)<0$. If either $x\in I$ or $y\in I$, then $nb(I\cup\{u\})\geq nb(I)$. If both $x\notin I$ and $y\notin I$, then $nb(I\cup\{u\})<nb(I)$. In other word, including $u$ has positive profit if and only if either $x$ or $y$ is included. we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that either $x\in S^{\prime}$ or $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
Case 2: $p(u,y)\leq w(u)<p(u,x)$. In this case, we consider whether the three vertices $u,x,y$ can be included in $I$.
When $w(u)\geq 0$ and $w(u)\geq p(u,x)+p(u,y)$, the all three vertices may be included in $I$. All combinations of $x,y,u$ may appear in $I$. Except that if $x$ is included and $y$ is not included in $I$, then $u\notin I$ because $nb(I\cup\{u\})=nb(I\setminus\{u\})+w(u)-p(u,x)<nb(I\setminus\{u\})$. In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that either $x\notin S^{\prime}$ or $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
When $w(u)\geq 0$ and $w(u)<p(u,x)+p(u,y)$, the three vertices can not be included in $I$. Moreover, $u$ and $x$ cannot be both included by $I$ because otherwise $nb(I\setminus u)>nb(I)$ contradicts the optimality of $I$. If $x\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})\geq nb(I\setminus\{u\})$. In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
When $w(u)<0$ and $w(u)\geq p(u,x)+p(u,y)$, the three vertices can be included in $I$. If $y\notin I$, then $nb(I\cup\{u\})\leq nb(I\setminus\{u\})$ regardless of whether $x\in I$. If $y\in I$, then $nb(I\cup\{u\})\geq nb(I\setminus\{u\})$ regardless of whether $x\in I$ In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
When $w(u)<0$ and $w(u)<p(u,x)+p(u,y)$, the three vertices can not be included in $I$. $u$ can appear in $I$ iff $y$ is included and $x$ is not included in $I$. In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$ and $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
Case 3: $w(u)<p(u,y)$. When $w(u)\geq 0$. $u$ can appear in $I$ iff neither $x$ nor $y$ is included in $I$. In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$ and $y\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
When $w(u)<0$ and $w(u)<p(u,x)+p(u,y)$. $u$ cannot appear in $I$ because otherwise $nb(I\setminus\{u\})>nb(I)$ a contradiction. Therefore, we can remove $u$ directly.
When $w(u)<0$ and $w(u)\geq p(u,x)+p(u,y)$. $u$ can appear in $I$ iff both $x$ and $y$ are included in $I$. In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\in S^{\prime}$ and $y\in S^{\prime}$. It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$. ∎
Proof of R10.
Observe that there exists a vertex set $I\in MGIS(G)$ that includes at least one of $x,u$ must be included in $I$. Because if none of them is included in some $I\in MGIS(G)$, then $nb(I\cup\{u\})-nb(I)\geq w(u)-\sum_{v\in N_{r}(u)}\max(0,p(u,v))\geq 0$ and therefore $I\cup\{u\}\in MGIS(G)$.
Given an arbitrary $I\in MGIS(G)$. Furthermore, $x$ and $u$ cannot be both included in $I$ since $p(u,x)=M$. In summary, there are two possible states in $I$: either $x\in I,u\notin I$ or $u\in I,x\notin I$. Therefore, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,u$ that could appear in a set in $MGIS(G)$. ∎
Proof of R11.
Observe that there exists a vertex set $I\in MGIS(G)$ that includes at least one of $u,x,y$. If none of them is included in $I$, then $nb(I\cup\{u\})-nb(I)\geq w(u)-\sum_{v\in N_{r}(u)}\max(0,p(u,v))\geq 0$ In addition, at most one of $u,x,y$ can be included in $I$ because they are permanent neighbors of each other.
Given an arbitrary $I\in MGIS(G)$. When $w(u)\geq\tilde{w}(x)$. If one of $x,y$ is included in $I$,say $x$, then $nb(I\cup\{u\}\setminus\{x\})\geq nb(I)$. In summary, $u$ must be included in some set in $MGIS$.
When $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$. If $y$ is included in $I$, then $u\notin I$ and $nb(I\cup\{u\}\setminus\{y\})\geq nb(I)$. Therefore, there is a vertex set $I^{\prime}\in MGIS(G)$ that $y\notin I^{\prime}$ and exactly one of $u,x$ is included in $I^{\prime}$. We conclude that there exists a vertex set $S\in MGIS(G)$ that $u\in S$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
When $w(u)<\tilde{w}(y)$. $u$ can appear in $I$ iff both $x,y$ are not included in $I$. In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$ and $y\notin S^{\prime}$. It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$. ∎
Proof of R12.
Given an arbitrary $I\in MGIS(G)$.
When $w(u)\geq\tilde{w}(v)+w^{+}(N(u)\setminus N_{p}[v])$. If $v\in I$, then $u\notin I$. We can replace $N(u)\cap I$ with $u$ to get a larger or equal net benefit because $nb(I\cup\{u\}\setminus N(u))-nb(I)\geq w(u)-\tilde{w}(v)+w^{+}(N(u)\setminus N_{p}[v])\geq w(u)-\tilde{w}(v)+w^{+}(N(u)\cap I)\geq 0$. Therefore, $I\cup\{u\}\setminus N(u)\in MGIS(G)$. Since $v\in N_{p}(u)$, there must be some set in $MGIS(G)$ that does not include $v$.
When $w(u)\geq\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}[v])+\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p(u,x))$. If $v\in I$, then $u\notin I$. We can replace $N_{p}(u)\cap I$ with $u$ to get a larger net benefit because $nb(I\cup\{u\}\setminus N_{p}(u))-nb(I)\geq w(u)-\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}(v))+\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p(u,x))\geq w(u)-\tilde{w}(v)+w^{+}(N(u)\cap I)+\sum_{x\in N_{r}(u)\cap I}\max(0,p(u,x))\geq 0$. Since $v\in N_{p}(u)$, there must be some set in $MGIS(G)$ that does not include $v$. ∎
Proof of R13.
Assume, for the sake of contradiction, that neither $u$ nor $v$ is included in any set in $MGIS(G)$. If for any $I\in MGIS(G)$, $N(v)\cap I=\varnothing$, then $I\cup\{v\}\in MGIS(G)$ because $nb(I\cup\{v\})\geq nb(I)$, a contradiction. Note that $u\notin N(v)\cap I$. If there exists a set $I\in MGIS(G)$ such that $N(v)\cap I\neq\varnothing$, then $w(v)\geq w^{+}(N(v))-max(0,w(u))\geq w^{+}(S)$. $nb(I\cup\{v\}\setminus N(v))\geq nb(I)$, a contradiction.
Therefore, there exists a set in $MGIS(G)$ that includes at least one of $u,v$. We conclude that their common permanent neighbors must not be included in some set in $MGIS(G)$. ∎
Proof of R14.
Given an arbitrary $I\in MGIS(G)$. If $u$ is included in $I$, then $N_{p}(v)\cap I=\varnothing$. And $I\cup\{v\}\in MGIS(G)$ because $nb(I\cup\{v\})-nb(I\setminus\{v\})=w(v)-\sum_{x\in N_{r}(v)}\max(0,p(v,x))\geq 0$. The same argument holds when $v$ is included in $I$. Therefore, we can fold $u$ and $v$ into a vertex $v^{\prime}$. There exists $S\in MGIS(G)$ that $u\in S$ and $v\in S$ iff there exists $S^{\prime}\in MGIS(G^{\prime})$ that $v^{\prime}\in S^{\prime}$. ∎
B Comparative Details on 216 Instances of SET-1
Table 6: Comparative results on SET-1 instances
Instance $|V|$ $|E_{p}|$ $|E_{r}|$ CB&B LA-B&B ILS-VND ALS RLS
BKV time(s) BKV time(s) BKV htime(s) BKV htime(s) BKV htime(s)
$p=0.25$, Setting 1
brock200_2_A 200 7416 2460 986 280 986 2 986 $<$1 986 $<$1 986 $<$1
brock400_2_A 400 44847 14939 765 441 765 1 765 $<$1 765 $<$1 765 $<$1
C125.9_A 125 5250 1713 454 $<$1 454 $<$1 454 $<$1 454 $<$1 454 $<$1
C250.9_A 250 20940 7044 581 10 581 $<$1 581 $<$1 581 $<$1 581 $<$1
gen200_p0.9_55_A 200 13434 4176 535 4 535 $<$1 535 $<$1 535 $<$1 535 $<$1
gen400_p0.9_75_A 400 53671 18149 628 82 628 $<$1 628 3 628 $<$1 628 $<$1
hamming8-4_A 256 15580 5284 1094 240 1094 $<$1 1094 $<$1 1094 $<$1 1094 $<$1
keller4_A 171 7057 2378 941 15 941 $<$1 941 $<$1 941 $<$1 941 $<$1
MANN_a27_A 378 53050 17501 533 28 533 $<$1 533 $<$1 533 $<$1 533 $<$1
p_hat300-1_A 300 8239 2694 2308 10800 2680 10800 2744 $<$1 2744 $<$1 2744 $<$1
p_hat300-2_A 300 16403 5525 2076 5182 2076 608 2076 $<$1 2076 $<$1 2076 $<$1
p_hat300-3_A 300 25132 8258 739 100 739 $<$1 739 $<$1 739 $<$1 739 $<$1
brock200_2_B 200 7416 2460 962 276 962 2 962 $<$1 962 $<$1 962 $<$1
brock400_2_B 400 44847 14939 741 445 741 1 741 $<$1 741 $<$1 741 $<$1
C125.9_B 125 5250 1713 437 1 437 $<$1 437 $<$1 437 $<$1 437 $<$1
C250.9_B 250 20940 7044 549 11 549 $<$1 549 $<$1 549 $<$1 549 $<$1
gen200_p0.9_55_B 200 13434 4176 510 5 510 $<$1 510 $<$1 510 $<$1 510 $<$1
gen400_p0.9_75_B 400 53671 18149 595 80 595 $<$1 595 $<$1 595 $<$1 595 $<$1
hamming8-4_B 256 15580 5284 1094 234 1094 $<$1 1094 $<$1 1094 $<$1 1094 $<$1
keller4_B 171 7057 2378 941 15 941 $<$1 941 $<$1 941 $<$1 941 $<$1
MANN_a27_B 378 53050 17501 503 31 503 $<$1 503 $<$1 503 $<$1 503 $<$1
p_hat300-1_B 300 8239 2694 2259 10800 2643 10800 2712 $<$1 2712 $<$1 2712 $<$1
p_hat300-2_B 300 16403 5525 2062 5194 2062 680 2062 $<$1 2062 $<$1 2062 $<$1
p_hat300-3_B 300 25132 8258 713 102 713 $<$1 713 $<$1 713 $<$1 713 $<$1
brock200_2_C 200 7416 2460 932 280 932 2 932 $<$1 932 $<$1 932 $<$1
brock400_2_C 400 44847 14939 698 455 698 2 698 $<$1 698 $<$1 698 $<$1
C125.9_C 125 5250 1713 403 1 403 $<$1 403 $<$1 403 $<$1 403 $<$1
C250.9_C 250 20940 7044 502 12 502 $<$1 502 $<$1 502 $<$1 502 $<$1
gen200_p0.9_55_C 200 13434 4176 467 4 467 $<$1 467 $<$1 467 $<$1 467 $<$1
gen400_p0.9_75_C 400 53671 18149 533 81 533 $<$1 533 $<$1 533 $<$1 533 $<$1
hamming8-4_C 256 15580 5284 1094 229 1094 1 1094 $<$1 1094 $<$1 1094 $<$1
keller4_C 171 7057 2378 941 14 941 $<$1 941 $<$1 941 $<$1 941 $<$1
MANN_a27_C 378 53050 17501 443 30 443 $<$1 443 $<$1 443 $<$1 443 $<$1
p_hat300-1_C 300 8239 2694 2205 10800 2566 10800 2649 $<$1 2649 $<$1 2649 $<$1
p_hat300-2_C 300 16403 5525 2033 5360 2033 778 2033 $<$1 2033 $<$1 2033 $<$1
p_hat300-3_C 300 25132 8258 688 111 688 $<$1 688 $<$1 688 $<$1 688 $<$1
$p=0.25$, Setting 2
brock200_2_A 200 7416 2460 1489 452 1489 5 1489 $<$1 1489 $<$1 1489 $<$1
brock400_2_A 400 44847 14939 1084 593 1084 2 1084 2 1084 $<$1 1084 $<$1
C125.9_A 125 5250 1713 685 1 685 $<$1 685 $<$1 685 $<$1 685 $<$1
C250.9_A 250 20940 7044 785 12 785 $<$1 785 $<$1 785 $<$1 785 $<$1
gen200_p0.9_55_A 200 13434 4176 778 5 778 $<$1 778 $<$1 778 $<$1 778 $<$1
gen400_p0.9_75_A 400 53671 18149 882 94 882 $<$1 882 $<$1 882 $<$1 882 $<$1
hamming8-4_A 256 15580 5284 1790 331 1790 2 1790 $<$1 1790 $<$1 1790 $<$1
keller4_A 171 7057 2378 1500 23 1500 $<$1 1500 $<$1 1500 $<$1 1500 $<$1
MANN_a27_A 378 53050 17501 683 32 683 $<$1 683 $<$1 683 $<$1 683 $<$1
p_hat300-1_A 300 8239 2694 3981 10800 4182 10800 4674 1 4674 $<$1 4674 $<$1
p_hat300-2_A 300 16403 5525 2994 8743 2994 9402 2994 $<$1 2994 $<$1 2994 $<$1
p_hat300-3_A 300 25132 8258 1180 137 1180 1 1180 $<$1 1180 $<$1 1180 $<$1
brock200_2_B 200 7416 2460 739 456 739 5 739 $<$1 739 $<$1 739 $<$1
brock400_2_B 400 44847 14939 534 611 534 2 534 2 534 $<$1 534 $<$1
C125.9_B 125 5250 1713 335 1 335 $<$1 335 $<$1 335 $<$1 335 $<$1
C250.9_B 250 20940 7044 385 13 385 $<$1 385 $<$1 385 $<$1 385 $<$1
gen200_p0.9_55_B 200 13434 4176 378 5 378 $<$1 378 $<$1 378 $<$1 378 $<$1
gen400_p0.9_75_B 400 53671 18149 432 96 432 $<$1 432 $<$1 432 $<$1 432 $<$1
hamming8-4_B 256 15580 5284 890 333 890 2 890 $<$1 890 $<$1 890 $<$1
keller4_B 171 7057 2378 750 22 750 $<$1 750 $<$1 750 $<$1 750 $<$1
MANN_a27_B 378 53050 17501 333 34 333 $<$1 333 $<$1 333 $<$1 333 $<$1
p_hat300-1_B 300 8239 2694 1981 10800 2082 10800 2324 2 2324 $<$1 2324 $<$1
p_hat300-2_B 300 16403 5525 1494 8947 1494 10800 1494 $<$1 1494 $<$1 1494 $<$1
p_hat300-3_B 300 25132 8258 580 134 580 1 580 1 580 $<$1 580 $<$1
brock200_2_C 200 7416 2460 364 460 364 5 364 $<$1 364 $<$1 364 $<$1
brock400_2_C 400 44847 14939 259 646 259 2 259 1 259 $<$1 259 $<$1
C125.9_C 125 5250 1713 160 1 160 $<$1 160 $<$1 160 $<$1 160 $<$1
C250.9_C 250 20940 7044 185 13 185 $<$1 185 $<$1 185 $<$1 185 $<$1
gen200_p0.9_55_C 200 13434 4176 178 5 178 $<$1 178 $<$1 178 $<$1 178 $<$1
gen400_p0.9_75_C 400 53671 18149 207 98 207 $<$1 207 $<$1 207 $<$1 207 $<$1
hamming8-4_C 256 15580 5284 440 355 440 2 440 $<$1 440 $<$1 440 $<$1
keller4_C 171 7057 2378 375 23 375 $<$1 375 $<$1 375 $<$1 375 $<$1
MANN_a27_C 378 53050 17501 158 33 158 $<$1 158 $<$1 158 $<$1 158 $<$1
p_hat300-1_C 300 8239 2694 981 10800 1032 10800 1149 1 1149 $<$1 1149 $<$1
p_hat300-2_C 300 16403 5525 744 10800 744 10658 744 $<$1 744 $<$1 744 $<$1
p_hat300-3_C 300 25132 8258 280 142 280 $<$1 280 $<$1 280 $<$1 280 $<$1
$p=0.50$, Setting 1
brock200_2_A 200 4916 4960 1298 8390 1298 93 1298 $<$1 1298 $<$1 1298 $<$1
brock400_2_A 400 29711 30075 1103 10800 1123 355 1123 $<$1 1123 $<$1 1123 $<$1
C125.9_A 125 3500 3463 627 5 627 $<$1 627 $<$1 627 $<$1 627 $<$1
C250.9_A 250 14017 13967 817 266 817 1 817 $<$1 817 $<$1 817 $<$1
gen200_p0.9_55_A 200 8908 9002 785 68 785 $<$1 785 $<$1 785 $<$1 785 $<$1
gen400_p0.9_75_A 400 35823 35997 895 5699 895 29 895 $<$1 895 $<$1 895 $<$1
hamming8-4_A 256 10329 10535 1301 7608 1301 49 1301 $<$1 1301 $<$1 1301 $<$1
keller4_A 171 4738 4697 1118 153 1118 2 1118 $<$1 1118 $<$1 1118 $<$1
MANN_a27_A 378 35345 35206 812 1217 812 6 812 $<$1 812 $<$1 812 $<$1
p_hat300-1_A 300 5505 5428 2568 10800 3026 10800 3129 $<$1 3129 $<$1 3129 $<$1
p_hat300-2_A 300 11051 10877 2063 10800 2477 10800 2477 $<$1 2477 $<$1 2477 $<$1
p_hat300-3_A 300 16820 16570 1029 3748 1029 38 1029 $<$1 1029 $<$1 1029 $<$1
brock200_2_B 200 4916 4960 1224 9589 1224 125 1224 $<$1 1224 $<$1 1224 $<$1
brock400_2_B 400 29711 30075 1010 10800 1035 445 1035 $<$1 1035 $<$1 1035 $<$1
C125.9_B 125 3500 3463 582 5 582 $<$1 582 $<$1 582 $<$1 582 $<$1
C250.9_B 250 14017 13967 744 292 744 2 744 $<$1 744 $<$1 744 $<$1
gen200_p0.9_55_B 200 8908 9002 716 72 716 $<$1 716 $<$1 716 $<$1 716 $<$1
gen400_p0.9_75_B 400 35823 35997 805 6602 805 34 805 $<$1 805 $<$1 805 $<$1
hamming8-4_B 256 10329 10535 1255 7780 1255 62 1255 $<$1 1255 $<$1 1255 $<$1
keller4_B 171 4738 4697 1094 156 1094 2 1094 $<$1 1094 $<$1 1094 $<$1
MANN_a27_B 378 35345 35206 707 1417 707 7 707 1 707 $<$1 707 $<$1
p_hat300-1_B 300 5505 5428 2492 10800 2950 10800 3023 $<$1 3023 $<$1 3023 $<$1
p_hat300-2_B 300 11051 10877 1979 10800 2405 10800 2405 $<$1 2405 $<$1 2405 $<$1
p_hat300-3_B 300 16820 16570 967 4067 967 46 967 $<$1 967 $<$1 967 $<$1
brock200_2_C 200 4916 4960 1101 10800 1101 204 1101 $<$1 1101 $<$1 1101 $<$1
brock400_2_C 400 29711 30075 861 10800 892 691 892 $<$1 892 $<$1 892 $<$1
C125.9_C 125 3500 3463 506 6 506 $<$1 506 $<$1 506 $<$1 506 $<$1
C250.9_C 250 14017 13967 623 334 623 2 623 $<$1 623 $<$1 623 $<$1
gen200_p0.9_55_C 200 8908 9002 597 91 597 $<$1 597 $<$1 597 $<$1 597 $<$1
gen400_p0.9_75_C 400 35823 35997 651 7858 651 50 651 $<$1 651 $<$1 651 $<$1
hamming8-4_C 256 10329 10535 1184 7447 1184 85 1184 $<$1 1184 $<$1 1184 $<$1
keller4_C 171 4738 4697 1049 145 1049 3 1049 $<$1 1049 $<$1 1049 $<$1
MANN_a27_C 378 35345 35206 552 1849 552 10 552 1 552 $<$1 552 $<$1
p_hat300-1_C 300 5505 5428 2303 10800 2818 10800 2897 $<$1 2897 $<$1 2897 $<$1
p_hat300-2_C 300 11051 10877 1852 10800 2263 10800 2263 $<$1 2263 $<$1 2263 $<$1
p_hat300-3_C 300 16820 16570 851 4709 851 69 851 $<$1 851 $<$1 851 $<$1
$p=0.50$, Setting 2
brock200_2_A 200 4916 4960 2034 10800 2034 677 2034 $<$1 2034 $<$1 2034 $<$1
brock400_2_A 400 29711 30075 1628 10800 1630 1323 1630 11 1630 $<$1 1630 $<$1
C125.9_A 125 3500 3463 1152 7 1152 $<$1 1152 $<$1 1152 $<$1 1152 $<$1
C250.9_A 250 14017 13967 1236 343 1236 3 1236 3 1236 $<$1 1236 $<$1
gen200_p0.9_55_A 200 8908 9002 1151 113 1151 $<$1 1151 3 1151 $<$1 1151 $<$1
gen400_p0.9_75_A 400 35823 35997 1335 8467 1335 69 1335 2 1335 $<$1 1335 $<$1
hamming8-4_A 256 10329 10535 2155 10800 2155 166 2155 1 2155 $<$1 2155 $<$1
keller4_A 171 4738 4697 1759 353 1759 5 1759 $<$1 1759 $<$1 1759 $<$1
MANN_a27_A 378 35345 35206 1226 1950 1226 10 1226 12 1226 $<$1 1226 $<$1
p_hat300-1_A 300 5505 5428 4740 10800 5637 10800 5637 $<$1 5637 $<$1 5637 $<$1
p_hat300-2_A 300 11051 10877 3342 10800 3750 10800 3943 $<$1 3943 $<$1 3943 $<$1
p_hat300-3_A 300 16820 16570 1658 7243 1658 157 1658 $<$1 1658 $<$1 1658 $<$1
brock200_2_B 200 4916 4960 984 10800 984 757 984 $<$1 984 $<$1 984 $<$1
brock400_2_B 400 29711 30075 765 10800 780 1385 780 10 780 $<$1 780 $<$1
C125.9_B 125 3500 3463 552 7 552 $<$1 552 $<$1 552 $<$1 552 $<$1
C250.9_B 250 14017 13967 586 391 586 3 586 4 586 $<$1 586 $<$1
gen200_p0.9_55_B 200 8908 9002 551 121 551 $<$1 551 1 551 $<$1 551 $<$1
gen400_p0.9_75_B 400 35823 35997 635 9395 635 72 635 $<$1 635 $<$1 635 $<$1
hamming8-4_B 256 10329 10535 1055 10800 1055 163 1055 $<$1 1055 $<$1 1055 $<$1
keller4_B 171 4738 4697 859 367 859 5 859 $<$1 859 $<$1 859 $<$1
MANN_a27_B 378 35345 35206 576 2161 576 11 576 13 576 $<$1 576 $<$1
p_hat300-1_B 300 5505 5428 2340 10800 2325 10800 2787 $<$1 2787 $<$1 2787 $<$1
p_hat300-2_B 300 11051 10877 1642 10800 1850 10800 1943 $<$1 1943 $<$1 1943 $<$1
p_hat300-3_B 300 16820 16570 808 7790 808 161 808 $<$1 808 $<$1 808 $<$1
brock200_2_C 200 4916 4960 459 10800 459 934 459 $<$1 459 $<$1 459 $<$1
brock400_2_C 400 29711 30075 341 10800 355 1711 355 12 355 $<$1 355 $<$1
C125.9_C 125 3500 3463 252 8 252 $<$1 252 $<$1 252 $<$1 252 $<$1
C250.9_C 250 14017 13967 261 501 261 3 261 3 261 $<$1 261 $<$1
gen200_p0.9_55_C 200 8908 9002 251 146 251 1 251 2 251 $<$1 251 $<$1
gen400_p0.9_75_C 400 35823 35997 285 10800 285 83 285 2 285 $<$1 285 $<$1
hamming8-4_C 256 10329 10535 505 10800 505 170 505 $<$1 505 $<$1 505 $<$1
keller4_C 171 4738 4697 409 389 409 5 409 $<$1 409 $<$1 409 $<$1
MANN_a27_C 378 35345 35206 251 2666 251 13 251 10 251 $<$1 251 $<$1
p_hat300-1_C 300 5505 5428 1138 10800 1125 10800 1362 $<$1 1362 $<$1 1362 $<$1
p_hat300-2_C 300 11051 10877 788 10800 900 10800 943 $<$1 943 $<$1 943 $<$1
p_hat300-3_C 300 16820 16570 383 9014 383 178 383 $<$1 383 $<$1 383 $<$1
$p=0.75$, Setting 1
brock200_2_A 200 2438 7438 1615 10800 1851 10800 1885 $<$1 1885 $<$1 1885 $<$1
brock400_2_A 400 14751 45035 1403 10800 1667 10800 1728 5 1728 $<$1 1728 $<$1
C125.9_A 125 1733 5230 1023 247 1023 4 1023 $<$1 1023 $<$1 1023 $<$1
C250.9_A 250 7073 20911 1200 10800 1236 5143 1236 2 1236 $<$1 1236 $<$1
gen200_p0.9_55_A 200 4425 13485 1186 10800 1206 601 1206 $<$1 1206 $<$1 1206 $<$1
gen400_p0.9_75_A 400 18063 53757 1155 10800 1490 10800 1490 $<$1 1490 $<$1 1490 $<$1
hamming8-4_A 256 5173 15691 1549 10800 1723 10800 1759 $<$1 1759 $<$1 1759 $<$1
keller4_A 171 2400 7035 1400 10800 1434 1573 1434 $<$1 1434 $<$1 1434 $<$1
MANN_a27_A 378 17580 52971 1313 10800 1323 10800 1323 $<$1 1323 $<$1 1323 $<$1
p_hat300-1_A 300 2734 8199 2504 10800 3932 10800 4164 $<$1 4164 $<$1 4164 $<$1
p_hat300-2_A 300 5603 16325 2120 10800 2902 10800 2990 1 2990 $<$1 2990 $<$1
p_hat300-3_A 300 8388 25002 1350 10800 1561 10800 1564 $<$1 1564 $<$1 1564 $<$1
brock200_2_B 200 2438 7438 1416 10800 1602 10800 1641 $<$1 1641 $<$1 1641 $<$1
brock400_2_B 400 14751 45035 1129 10800 1341 10800 1386 2 1386 $<$1 1386 $<$1
C125.9_B 125 1733 5230 856 395 856 8 856 $<$1 856 $<$1 856 $<$1
C250.9_B 250 7073 20911 960 10800 1001 10800 1001 1 1001 $<$1 1001 $<$1
gen200_p0.9_55_B 200 4425 13485 961 10800 983 1585 983 $<$1 983 $<$1 983 $<$1
gen400_p0.9_75_B 400 18063 53757 920 10800 1105 10800 1120 $<$1 1120 $<$1 1120 $<$1
hamming8-4_B 256 5173 15691 1356 10800 1486 10800 1579 $<$1 1579 $<$1 1579 $<$1
keller4_B 171 2400 7035 1230 10800 1268 4468 1268 $<$1 1268 $<$1 1268 $<$1
MANN_a27_B 378 17580 52971 922 10800 1011 10800 1021 $<$1 1021 $<$1 1021 $<$1
p_hat300-1_B 300 2734 8199 2306 10800 3547 10800 3886 $<$1 3886 $<$1 3886 $<$1
p_hat300-2_B 300 5603 16325 1762 10800 2588 10800 2782 $<$1 2782 $<$1 2782 $<$1
p_hat300-3_B 300 8388 25002 1144 10800 1209 10800 1299 $<$1 1299 $<$1 1299 $<$1
brock200_2_C 200 2438 7438 1158 10800 1303 10800 1321 $<$1 1321 $<$1 1321 $<$1
brock400_2_C 400 14751 45035 793 10800 961 10800 1033 $<$1 1033 $<$1 1033 $<$1
C125.9_C 125 1733 5230 644 10800 644 45 644 $<$1 644 $<$1 644 $<$1
C250.9_C 250 7073 20911 688 10800 734 10800 734 1 734 $<$1 734 $<$1
gen200_p0.9_55_C 200 4425 13485 707 10800 727 9879 727 $<$1 727 $<$1 727 $<$1
gen400_p0.9_75_C 400 18063 53757 613 10800 756 10800 772 1 772 $<$1 772 $<$1
hamming8-4_C 256 5173 15691 1120 10800 1248 10800 1378 $<$1 1378 $<$1 1378 $<$1
keller4_C 171 2400 7035 1060 10800 1109 10800 1109 $<$1 1109 $<$1 1109 $<$1
MANN_a27_C 378 17580 52971 583 10800 648 10800 651 2 651 $<$1 651 $<$1
p_hat300-1_C 300 2734 8199 1866 10800 3191 10800 3480 $<$1 3480 $<$1 3480 $<$1
p_hat300-2_C 300 5603 16325 1511 10800 2249 10800 2473 $<$1 2473 $<$1 2473 $<$1
p_hat300-3_C 300 8388 25002 886 10800 919 10800 1004 $<$1 1004 $<$1 1004 $<$1
$p=0.75$, Setting 2
brock200_2_A 200 2438 7438 2999 10800 3078 10800 3326 8 3326 $<$1 3326 $<$1
brock400_2_A 400 14751 45035 2603 10800 2787 10800 2941 12 2941 1 2941 3
C125.9_A 125 1733 5230 1837 365 1837 62.71 1837 $<$1 1837 $<$1 1837 $<$1
C250.9_A 250 7073 20911 2011 10800 2171 10800 2171 5 2171 $<$1 2171 $<$1
gen200_p0.9_55_A 200 4425 13485 2094 10800 2096 7832 2096 $<$1 2096 $<$1 2096 $<$1
gen400_p0.9_75_A 400 18063 53757 2164 10800 2310 10800 2404 19 2404 $<$1 2404 1
hamming8-4_A 256 5173 15691 2813 10800 3124 10800 3124 1 3124 $<$1 3124 $<$1
keller4_A 171 2400 7035 2532 10800 2690 10800 2690 $<$1 2690 $<$1 2690 $<$1
MANN_a27_A 378 17580 52971 1976 10800 2206 10800 2208 $<$1 2208 $<$1 2208 $<$1
p_hat300-1_A 300 2734 8199 5566 10800 $6$ 10800 7899 25 7899 $<$1 7899 $<$1
p_hat300-2_A 300 5603 16325 4102 10800 4714 10800 5343 $<$1 5343 $<$1 5343 $<$1
p_hat300-3_A 300 8388 25002 2477 10800 2658 10800 2838 2 2838 $<$1 2838 $<$1
brock200_2_B 200 2438 7438 1398 10800 1427 10800 1533 $<$1 1533 $<$1 1533 $<$1
brock400_2_B 400 14751 45035 1161 10800 1210 10800 1291 16 1291 1 1291 4
C125.9_B 125 1733 5230 837 523 837 84 837 $<$1 837 $<$1 837 $<$1
C250.9_B 250 7073 20911 911 10800 969 10800 971 3 971 $<$1 971 $<$1
gen200_p0.9_55_B 200 4425 13485 944 10800 946 10800 946 $<$1 946 $<$1 946 $<$1
gen400_p0.9_75_B 400 18063 53757 936 10800 1010 10800 1052 11 1054 1 1054 1
hamming8-4_B 256 5173 15691 1309 10800 1474 10800 1474 $<$1 1474 $<$1 1474 $<$1
keller4_B 171 2400 7035 1190 10800 1240 10800 1254 $<$1 1254 $<$1 1254 $<$1
MANN_a27_B 378 17580 52971 875 10800 956 10800 958 11 958 $<$1 958 $<$1
p_hat300-1_B 300 2734 8199 2666 10800 3216 10800 3818 2 3818 $<$1 3818 $<$1
p_hat300-2_B 300 5603 16325 1952 10800 2177 10800 2543 $<$1 2543 $<$1 2543 $<$1
p_hat300-3_B 300 8388 25002 1101 10800 1208 10800 1288 2 1288 $<$1 1288 $<$1
brock200_2_C 200 2438 7438 607 10800 607 10800 658 $<$1 658 $<$1 658 $<$1
brock400_2_C 400 14751 45035 461 10800 465 10800 500 3 500 $<$1 500 $<$1
C125.9_C 125 1733 5230 337 1519 337 204 337 $<$1 337 $<$1 337 $<$1
C250.9_C 250 7073 20911 359 10800 370 10800 372 5 372 $<$1 372 $<$1
gen200_p0.9_55_C 200 4425 13485 352 10800 371 10800 371 $<$1 371 $<$1 371 $<$1
gen400_p0.9_75_C 400 18063 53757 356 10800 379 10800 395 16 395 $<$1 395 $<$1
hamming8-4_C 256 5173 15691 556 10800 649 10800 652 $<$1 652 $<$1 652 $<$1
keller4_C 171 2400 7035 510 10800 554 10800 558 $<$1 558 $<$1 558 $<$1
MANN_a27_C 378 17580 52971 325 10800 335 10800 335 11 335 $<$1 335 $<$1
p_hat300-1_C 300 2734 8199 1205 10800 1491 10800 1793 1 1793 $<$1 1793 $<$1
p_hat300-2_C 300 5603 16325 854 10800 981 10800 1159 $<$1 1159 $<$1 1159 $<$1
p_hat300-3_C 300 8388 25002 453 10800 494 10800 529 1 529 $<$1 529 $<$1

### v2

@@
 The Generalized Independent Set (GIS) problem extends the classical maximum independent set problem by incorporating profits for vertices and penalties for edges.
-This generalized problem has been identified in diverse applications in fields such as forest harvest planning, competitive facility location, social network analysis, and even machine learning.
+This generalized problem has been identified in diverse applications in fields such as forest harvesting, competitive facility location, social network analysis, and even machine learning.
 However, solving the GIS problem in large-scale, real-world networks remains computationally challenging.
@@
 We then present a reduction-driven local search (RLS) algorithm that integrates these reduction rules into the pre-processing, the initial solution generation, and the local search components in a computationally efficient way.
-The RLS is empirically evaluated on 278 graphs arising from different application scenarios.
-The results indicates that the RLS is highly competitive – For most graphs, it achieves significantly superior solutions compared to other known solvers, and it effectively provides solutions for graphs exceeding 260 million edges, a task at which every other known method fails.
-Analysis also reveals that the data reduction plays a key role in achieving such a competitive performance.
+The RLS is empirically evaluated on 278 graphs drawn from different application scenarios.
+The results indicate that the RLS is highly competitive – For most graphs, it achieves significantly superior solutions compared to other known solvers, and it effectively provides solutions for graphs exceeding 260 million edges, a task at which every other known method fails.
+Analysis also reveals that data reduction plays a key role in achieving such a competitive performance.
 Keywords: Generalized independent set; Data reduction; Local search; Large graph
@@
 Each removable edge $e\in E_{r}$ is associated with a penalty $p(e)\in\mathbb{R}$.
-The generalized independent set of the graph $G$ is a vertex set $I\subseteq V$ such that no permanent edge exists between any two vertices of $I$ in $G$.
+A generalized independent set of the graph $G$ is a vertex set $I\subseteq V$ such that no permanent edge exists between any two vertices of $I$ in $G$.
 Then, the GIS problem asks for a generalized independent set $I$ of maximum net benefit, which is the total vertex profits of $I$ minus the total penalties of removable edges whose end vertices are in $I$.
 An example is shown in Figure 1.
-Figure 1: A toy example of the GIS problem. The set of red vertices is a generalized independent set of net benefit of $2-1+6+5=12$.
-When all the edges are permanent edges, the GIS problem equals the well-known maximum independent set problem (MIS) which asks for a maximum set of non-adjacent vertices from a given graph.
-As the GIS allows a more flexible definition of vertex profit and edge penalties, it captures a broader range of real-world scenarios than the classical independent set problem.
-We leave detailed examples of its applicability in diverse domains in Section 2.1.
+Figure 1: A toy example of the GIS problem. The set of red vertices is a generalized independent set of net benefit of (2+6+5)1=12(2+6+5)-1=12.
+When all the edges are permanent and each vertex has a profit of 1, the GIS problem becomes the well-known maximum independent set problem (MIS), which asks for a maximum-cardinality set of non-adjacent vertices in a given graph.
+As the GIS problem allows a more flexible definition of vertex profit and edge penalties, it captures a broader range of real-world scenarios than the classical independent set problem.
+We give detailed examples of applications in diverse domains in Section 2.1.
 However, since MIS is NP-hard, W[1]-hard and difficult to approximate (Cygan et al.
-2015), the equivalence indicates that the GIS is also computationally hard.
+2015), the equivalence above indicates that the GIS problem is also computationally hard.
 Existing works have proposed various approaches to solve the GIS problem practically.
@@
 (2021)).
-For example, the state-of-the-art ALS algorithm proposed by Zheng et al.
+For example, the most recent algorithm, named Adaptive Local Search (ALS), proposed by Zheng et al.
 (2024) achieves near-optimal solutions on all standard benchmark instances and scales to problems with up to 18,000 vertices.
-Nevertheless, we observe that these heuristic algorithms struggle to maintain solution quality in even larger instances.
-The data reduction, which reduces input size while preserving problem optimality, has recently emerged as a promising technique for improving algorithm scalability on large instances.
+Nevertheless, we observed that these heuristic algorithms struggle to maintain solution quality in even larger instances.
+Data reduction, which reduces input size while preserving problem optimality, has recently emerged as a promising technique for improving algorithm scalability on large instances.
 The technique has been successfully applied to problems such as independent set, vertex cover, and maximum clique (Großmann et al.
@@
 2019).
-For the GIS problem, research on data reduction is still in its early stages.
-There are three main challenges in developing data reduction techniques for the GIS problem: (1) The complex interplay among vertex profits, edge penalties, and the underlying graph structure, (2) Translating reduction rules into efficient, practical algorithms, often requiring incremental implementations (Akiba and Iwata 2016, Chang et al.
+For the GIS problem, research on data reduction is still at its early stages.
+There are three main challenges in developing data reduction techniques for the GIS problem: (1) The complex interplay among vertex profits, edge penalties, and the underlying graph structure, (2) Translating reduction rules into efficient, practical algorithms, which often requires incrementally applying the reduction rules (Akiba and Iwata 2016, Chang et al.
 2017), (3) Integrating reductions into local search algorithms while balancing search time and efficiency.
@@
     We propose an efficient reduction-driven local search algorithm, named RLS, to solve the GIS problem.
-We incorporate our proposed reduction rules into three main components of RLS: the pre-processing, which incrementally reduces the size of original input instance, the initial solution generation, which interplays the vertex selection and data reduction, and the neighborhood search, which uses a novel reduction-based move operator to guide the search trajectory.
+We incorporate our proposed reduction rules into three main components of the RLS: the pre-processing, which incrementally reduces the size of the original input instance, the initial solution generation, which interplays the vertex selection and data reduction, and the neighborhood search, which uses a novel reduction-based move operator to guide the search trajectory.
   3. 3.
-    We evaluate the proposed RLS algorithm on a total of 278 instances.
-On small instances, RLS consistently achieves the known optimal solutions with significantly shorter running time than exact algorithms.
-On large instances, RLS exhibits superior scalability over existing heuristic algorithms in solution quality.
-Notably, RLS successfully handles instances with over 260 million edges, where existing heuristic methods fail.
-Our experimental results also demonstrate that data reduction techniques significantly improve the computational efficiency of RLS.
-The rest of the paper is organized as follows.
-Section 2 reviews applications and related works, followed by Section 3 which introduces the notations and definitions.
-In Section 4, we present our reduction rules for GIS.
+    We evaluate the proposed the RLS algorithm on a total of 278 instances.
+On small instances with up to 400 vertices and 80,000 edges, the RLS consistently achieves the known optimal solutions with significantly shorter running time than exact algorithms.
+On large instances, the RLS exhibits superior scalability over existing heuristic algorithms in solution quality.
+Notably, the RLS successfully handles instances with over 18 million vertices and over 260 million edges, where existing heuristic methods fail.
+Our experimental results also demonstrate that data reduction techniques significantly improve the computational efficiency of the RLS.
+The rest of the paper is organized as follows: Section 2 reviews applications and related works, followed by Section 3 which introduces the notations and definitions.
+In Section 4, we present our reduction rules for the GIS problem.
 Then, in Section 5, we describe the RLS algorithm.
@@
 2 Background
+In this section, we provide some application examples of the GIS model and review existing GIS algorithms.
 2.1 Applications Related to GIS
 In the following, we highlight several important applications of the GIS problem:
-Forest harvest planning.
-The planning of forest harvest involves selecting forest areas (cells) for timber, aiming to maximize benefits like timber value while minimizing ecological damage from excessive cutting.
+Forest harvesting.
+The planning of forest harvesting involves selecting forest areas (cells) for timber, aiming to maximize benefits like timber value while minimizing ecological damage from excessive cutting.
 As a GIS problem, forest cells are vertices weighted by harvesting benefits, and edges represent penalties for harvesting adjacent cells (Hochbaum and Pathria (1997)).
@@
 The facility location problem, modeled with vertices as facilities and edges as geographic relationships, can be framed as a GIS problem to maximize utilities (Hochbaum (2004)).
-Similar spatial conflicts appear in airplane seating (Pavlik et al.
+Similar spatial conflicts exist in airplane seating (Pavlik et al.
 2021) and label placement (Mauri et al.
@@
 Social network analysis.
-The GIS problem has broad applications in social network analysis.
+Potentially, the GIS problem has broad applications in social network analysis.
 A maximum independent set corresponds to the complement of a minimum vertex cover and is equivalent to a maximum clique in the graph’s complement (Sanchis and Jagota (1996)).
@@
 Exact Algorithms.
-In pursuit of optimal solutions, Hochbaum (2004) introduced an integer programming model that considers both nodes and edges as variables, with each edge associated with a specific constraint.
-However, this model relies on binary variables for removable edges, resulting in large formulations; for example, a graph with 400 vertices requires 80,000 variables, making direct linear programming infeasible.
-To address this, Colombi et al.
+In pursuit of optimal solutions, Hochbaum (2004) introduced an integer linear programming model that assigns variables to each node and edge, and associates each edge with a specific constraint.
+As a result, the integer linear program for a graph of 2000 vertices (e.g.
+C2000.5 in the DIMACS competition set) has approximately 500,000 variables and 500,000 constraints, which is quite large for modern ILP solvers like CPLEX.
+(It fails to even report a solution within 3 hours for this instance.) Then, Colombi et al.
 (2017) reformulated the GIS problem as an unconstrained binary quadratic program (UBQP), and developed a branch-and-cut algorithm, along with additional valid inequalities.
 Note that they also generated a total of 216 random instances which often serve as benchmark datasets in the later work.
-Hosseinian and Butenko (2019) employed another linear program to bound the optimal solution, which is later incorporated into a branch-and-bound framework.
+Hosseinian and Butenko (2019) employed another quadratic program to bound the optimal solution, which was later incorporated into a branch-and-bound framework.
 Recently, Zheng et al.
 (2024) used a Lagrangian relaxation in their branch-and-bound framework to estimate tight upper bounds.
-They showed that their algorithm is able to solve the GIS instances with up to 400 vertices within $3$ hours.
+Their experiments showed that their algorithm is able to solve the GIS instances with up to 400 vertices within $3$ hours.
 Heuristic Algorithms.
 Kochenberger et al.
-(2007) also developed a UBQP formulation for the GIS problem solved via tabu search.
+(2007) also developed a UBQP formulation of the GIS problem and solved it using tabu search.
 Colombi et al.
@@
 More recently, Zheng et al.
-(2024) proposed a more refined local search heuristic, namely ALS, which explores three move operations – adding one vertex to the solution, dropping one vertex from the solution, and replacing a vertex in the solution with another one not in the solution.
-The ALS algorithm reported state-of-the-art results on the set of 216 benchmark instances proposed by Colombi et al.
+(2024) proposed a more refined local search heuristic“AL” to explores three move operations – adding one vertex to the solution, dropping one vertex from the solution, and replacing a vertex in the solution with another one not in the solution.
+The ALS algorithm has been reported to achieve state-of-the-art performance on the standard benchmark set of 216 instances proposed by Colombi et al.
 (2017).
-Their algorithm is able to find high-quality solutions on large instances with up to $17903$ vertices within $5$ minutes.
-However, its performance on even larger graphs, such as the uk2002 graph with 18 million vertices, remains unknown.
+In particular, it is able to find high-quality solutions for instances with up to $17,903$ vertices within $5$ minutes.
+However, its performance on graphs with more than 20,000 vertices has not been systematically evaluated.
+Our experiments show that ALS quickly becomes impractical on such large-scale instances.
 Data Reduction Techniques.
-The data reduction techniques are originally introduced in the framework of parameterized complexity analysis by the name of kernelization (Cygan et al.
+Data reduction techniques for integer programming models can be traced back several decades, referred to as variable fixing and coefficient reduction (crowder1983solving).
+Then, the parameterized complexity community has also adopted this technique, known as kernelization, to achieve strong theoretical bounds on the time complexity of many hard combinatorial graph problems (Cygan et al.
 2015).
@@
 (2016) and Chang et al.
-(2017) combine reductions with local search and heuristics to achieve near-optimal performance.
+(2017) integrate reductions with heuristics to achieve near-optimal performance.
 These techniques enable handling graphs with millions of vertices and edges (Lamm et al.
@@
 However, as for the GIS problem, no existing work focuses on reduction rules.
-And extending existing reductions to vertex- and edge-weighted graphs for the GIS problem remains challenging.
+Extending existing reductions for the independent set problem to the GIS problem, where both vertices and edges are weighted, remains challenging.
 3 Preliminary
-By the definition, the input instance of the GIS problem is an (undirected weighted) graph $G=(V,E=E_{p}\cup E_{r})$ where each vertex $v\in V$ is associated with a profit $w(v)\in\mathbb{R}$ and each removable edge $e\in E_{r}$ is associated with a penalty $p(e)$.
-In order to reformulate the problem in a uniform way, we extend the definition by associating each edge $e\in E_{p}$ also with a penalty $p(e)=M$ where $M\in R$ is a sufficiently large value, e.g., the absolute sum of all vertex profits and removable edge penalties, $\sum_{v\in V}|w(v)|+\sum_{e\in E_{r}}|p(e)|$).
-For a set $S\subseteq V$, we formally define $nb(S)$ as the net benefit of $S$, i.e.
-$nb(S)=\sum_{v\in S}w(v)-\sum_{\{u,v\}\subseteq S,(u,v)\in E}p(u,v).$
-Then, the GIS problem is equal to finding a set $S\subseteq V$ that maximizes the net benefit of $S$ from a given graph $G$.
-The optimal objective value of the GIS problem in $G$ is denoted by $\alpha(G)$, i.e.
-$\alpha(G)=\max_{S\subseteq V}nb(S)$
-We also denote $MGIS(G)$ as the set of all vertex sets $S$ with optimal objective value.
-It is clear that every $S\in MGIS(S)$ is a generalized independent set which is defined in the introductory section.
-We say a generalized independent set $S\subseteq V$ is maximal if there is no other generalized independent set $S^{\prime}$ such that $S\subset S^{\prime}$ and $nb(S^{\prime})\geq nb(S)$.
-For convenience, we import more notations in Table 1.
-Note that we use $\tilde{w}(v)$ to represent the profit of $v$ plus the non-negative penalties of edges that are incident to $v$.
-The $w^{+}(S)$ is an upper bound on the net benefit of any subset of $S$.
-For example, in Figure 1, $\tilde{w}(u)=2+1=3$.
-When $S$ denotes the set of red vertices in Figure 1, $w^{+}(S)=3+6+5=14$.
-In the rest of the paper, we use notation $\coloneqq$ to represent assignment and $=$ to indicate equivalence.
+According to the definition, an input instance of the GIS problem is an (undirected weighted) graph $G=(V,E=E_{p}\cup E_{r})$ where each vertex $v\in V$ is associated with a profit $w(v)\in\mathbb{R}$ and each removable edge $e\in E_{r}$ is associated with a penalty $p(e)\in\mathbb{R}$.
+For a vertex set $S\subseteq V$, we define $nb(S)$ as the net benefit of $S$, i.e.,
+$nb(S)=\sum_{v\in S}w(v)-\sum_{u\in S,v\in S,e=\{u,v\}\in E_{r}}p(e).$
+Then, a generalized independent set of $G$ is a vertex subset $S\subseteq V$ such that $G[S]$ contains no permanent edge.
+The GIS problem can be formulated as finding a generalized independent set such that its net benefit is maximized.
+We denote the optimal objective value of the GIS problem in $G$ as $\alpha(G)$, i.e.,
+$\alpha(G)=\max_{S\subseteq V}nb(S)\quad s.t.\quad E_{p}\cap\{\{u,v\}\mid u,v\in S\}=\varnothing$
+We also denote $MGIS(G)$ as the family of generalized independent sets $S\subseteq V$ that achieve the optimal objective value $\alpha(G)$. Each $S\in MGIS(G)$ corresponds to an optimal solution.
 Table 1: Table of Notations
-$G=(V,E)$ a graph $G$ with vertex set $V$ and edge set $E$
-$E_{p},E_{r}$ the set of pernament edges and the set of removable edges respectively
-$w(v),p(u,v)$ the profit of including a vertex $v$, and the penalty of including both vertices $u$ and $v$
-$N(v),N_{p}(v),N_{r}(v)$ the sets of all neighbors, permanent neighbors and removable neighbors of $v$, respectively
+$G=(V,E=E_{p}\cup E_{r})$
+a graph $G$ with vertex set $V$ and edge set $E$,
+$E_{p},E_{r}$ represent the set of permanent edges and the set of removable edges, respectively
+$w(v),p(\{u,v\})$ the profit of including a vertex $v$, and the penalty of including both vertices $u$ and $v$
+$N(v),N_{p}(v),N_{r}(v)$
+subsets of $V$, representing the sets of vertices connected to $v$, vertices connected to $v$ by
+a permanent edge and vertices connected to $v$ by a removable edge, respectively
 $|N(v)|,|N_{p}(v)|,|N_{r}(v)|$ the degree, permanent degree, and removable degree of $v$, respectively
 $N[v],N_{p}[v],N_{r}[v]$
-the sets of all neighbors, permanent neighbors and removable neighbors of $v$ including $v$
-itself, respectively
+subsets of $V$, representing the sets $N(v)\cup\{v\}$, $N_{p}(v)\cup\{v\}$, $N_{r}(v)\cup\{v\}$, respectively
 $\Delta$ the maximum degree of vertices in $G$
 $\tilde{w}(v)$
-$\tilde{w}(v)=w(v)+\sum_{x\in N_{r}(v)}\max(0,-p(v,x))$
-$S,G\setminus S$ a vertex subset of $G$, and the graph obtained by removing $S$ and all incident edges
+$\tilde{w}(v)=w(v)+\sum_{x\in N_{r}(v)}\max(0,-p(\{v,x\}))$
+S,G[S],E(G[S])S,{G[S],E(G[S])} a vertex subset of $G$, the subgraph induced by $S$, and the edge set of the induced subgraph
 $w(S),w^{+}(S)$ $w(S)=\sum_{v\in S}w(v)$, $w^{+}(S)=\sum_{v\in S}\max(0,\tilde{w}(v))$
-$nb(S)$ the net benefit of vertex set $S$, $nb(S)=\sum_{v\in S}w(v)-\sum_{\{u,v\}\subseteq S}p(u,v)$
-$\alpha(G),MGIS(G)$ the largest net benefit value, and the set of all vertex sets whose net benefit is $\alpha(G)$
-Integer Programming. The GIS problem can be easily formulated as (Colombi et al. 2017):
-$\alpha(G)=\max\quad$ $\sum_{i\in V}w(i)x_{i}-\sum_{\{i,j\}\in E_{r}}p(i,j)y_{ij}$ (1)
-s.t. $x_{i}+x_{j}\leq 1$ $(i,j)\in E_{p}$ (2)
-$x_{i}+x_{j}-y_{ij}\leq 1$ $(i,j)\in E_{r}$ (3)
+$nb(S)$ the net benefit of a vertex set $S$, $nb(S)=\sum_{v\in S}w(v)-\sum_{u\in S,v\in S,e=\{u,v\}\in E_{r}}p(e)$
+$\alpha(G),MGIS(G)$
+the largest net benefit value achievable without including any permanent edge,
+and the family of all subsets of $V$ attaining $\alpha(G)$ value under the same constraint
+For convenience, we report more notations in Table 1.
+Note that we use $\tilde{w}(v)$ to represent the profit of $v$ plus the absolute sum of non-positive penalties of edges that are incident to $v$.
+Given that $S\subseteq V$ is a generalized independent set, $w^{+}(S)$ is an upper bound on the net benefit of $S$.
+For example, in Figure 1, $\tilde{w}(u)=2+2.5=4.5$.
+When $S$ denotes the set of red vertices in Figure 1, $w^{+}(S)=4.5+6+5=15.5$.
+In the rest of the paper, we use the notation $\coloneqq$ to represent assignment and == to indicate equivalence.
+Integer Programming. The GIS problem can be conveniently formulated as follows (Colombi et al. 2017):
+$\alpha(G)=\max\quad$ $\sum_{i\in V}w(i)x_{i}-\sum_{\{i,j\}\in E_{r}}p(\{i,j\})y_{ij}$ (1)
+s.t. $x_{i}+x_{j}\leq 1$ $\{i,j\}\in E_{p}$ (2)
+$x_{i}+x_{j}-y_{ij}\leq 1$ $\{i,j\}\in E_{r}$ (3)
+${x_{i}\geq y_{ij},x_{j}\geq y_{ij}}$ $\{i,j\}\in E_{r}$ (4)
 $x_{i}\in\{0,1\}$ $i\in V$ (5)
-$y_{ij}\in\{0,1\}$ $(i,j)\in E_{r}$ (5)
-where binary variables $x_{i}$ indicate whether node $i$ is selected, and binary variables $y_{ij}$ indicate whether the edge is included.
+$y_{ij}\in\{0,1\}$ $\{i,j\}\in E_{r}$ (6)
+where binary variables $x_{i}$ indicate whether the vertex $i$ is selected, and binary variables $y_{ij}$ indicate whether the removable edge $\{i,j\}$ is included.
 4 Reduction Rules
 In this section, we develop reduction rules that identify reducible graph structures without compromising optimality.
-We rely on four operations to reduce the current instance to another one.
-(1) Edge transformation, which identifies removable edges that can be deleted or transferred to permanent edges; (2) Vertex inclusion, which identifies vertices that are in some set $I\in MGIS(G)$; (3) Vertex exclusion, which identifies vertices that are not in some set $I\in MGIS(G)$.
+We rely on four operations to reduce the current instance to a simplified instance.
+(1) Edge transformation, which identifies removable edges that can be deleted or transformed to permanent edges; (2) Vertex inclusion, which identifies vertices that are in at least one set $I\in MGIS(G)$; (3) Vertex exclusion, which identifies vertices that do not belong to a set $I\in MGIS(G)$.
 (4) Folding, which identifies vertices that can be merged into a single vertex.
-Based on the four operations, we will introduce a total of 14 reduction rules.
+Based on the four operations, we introduce a total of 14 reduction rules as follows.
+For simplicity, we denote the graph obtained after applying a reduction rule by $G^{\prime}=(V^{\prime},E^{\prime}=E_{p}^{\prime}\cup E_{r}^{\prime})$.
+Unless otherwise specified, we assume that $w^{\prime}(v)$, the profit of a vertex $v\in V^{\prime}$,keeps its original profit $w(v)$ if $v$ also exists in $V$, and $p^{\prime}(e)$, the penalty of a removable edge $e\in E_{r}^{\prime}$, keeps its original penalty $p(e)$ if $e$ also exists in $E_{r}$.
 4.1 Edge Transformation
@@
 R 1 (Edge Removal Reduction).
-If there exists an edge $(u,v)\in E$ such that $p(u,v)=0$, then $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}$ is obtained by removing the edge $(u,v)$ from $G$.
-Technical proofs of the correctness of this reduction rule, as well as the following reduction rules, are deferred in Appendix A.
-R1 implies that removing an edge $(u,v)$ that $p(u,v)=0$ does not change optimality.
+If there exists an edge $e=\{u,v\}\in E_{r}$ such that $p(e)=0$, then $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}$ is obtained by only removing the edge $e$ from $G$.
+The proofs of the correctness of this reduction rule, and the subsequent reduction rules, are deferred to Appendix A. R1 implies removing an edge $e$ with $p(e)=0$ does not change optimality.
 R 2 (Edge-penalty Reduction).
-If there is an edge $(u,v)$ such that $p(u,v)<M$ and $p(u,v)>min(\tilde{w}(u),\tilde{w}(v))$, then $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}$ is obtained by setting $p(u,v):=M$ in $G$.
-Compared to permanent edges, removable edges are more challenging to reduce because of the difficulties of identifying whether both endpoints of a removable edge are in some $S\in MGIS(G)$.
-R2 allows us to transform removable edges into permanent edges without sacrificing optimality, underpinning many further reduction rules.
+If there is an edge $e=\{u,v\}\in E_{r}$ such that $p(e)>min(\tilde{w}(u),\tilde{w}(v))$, then $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}$ is obtained by only moving $e$ from $E_{r}$ to $E_{p}$ in $G$.
+Compared to a permanent edge, a removable edge is more challenging to reduce because of the difficulties of identifying whether both endpoints of a removable edge are in a set $S\in MGIS(G)$.
+R2 allows us to transform a removable edge into a permanent edge without sacrificing optimality, underpinning many further reduction rules.
 4.2 Neighborhood-based Reductions
@@
 R 3 (Neighborhood Weight Reduction).
-If there is a vertex $u\in V$ that $w(u)\geq w^{+}(N(u))$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus N_{p}[u]$, and $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$.
-The neighborhood weight reduction generalizes the neighborhood reduction rule in (Lamm et al.
-2019) for solving the MIS problem.
-We further extend R3 by considering more neighboring structures and obtain R4-R6.
+If there is a vertex $u\in V$ such that $w(u)\geq w^{+}(N(u))$, then $u$ is in at least one set $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus N_{p}[u]]$, and $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$ for each $v\in N_{r}(u)$.
+We extend R3 by considering more neighboring structures to obtain R4-R6.
 R 4 (Neighborhood Penalty Reduction).
-If there exists a vertex $u\in V$ such that $w(u)\geq w^{+}(N_{p}(u))+\sum_{x\in N_{r}(u)}\max(0,p(u,x))$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus N_{p}[u]$ and $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$.
+If there exists a vertex $u\in V$ such that $w(u)\geq w^{+}(N_{p}(u))+\sum_{x\in N_{r}(u)}\max(0,p(\{u,x\}))$, then $u$ is in at least one set $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus N_{p}[u]]$, and $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$ for each $v\in N_{r}(u)$.
 R 5 (Negative Profit Reduction).
-If there exists a vertex $u\in V$ such that $w(u)<0$ and $\sum_{v\in N(u)}\min(0,p(u,v))>w(u)$, then $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$.
+If there exists a vertex $u\in V$ such that $w(u)<0$ and $\sum_{v\in N_{r}(u)}\min(0,p(\{u,v\}))>w(u)$, then $u$ is not in any $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$.
+Next, we show the clique reduction rule. A clique of $G$ is a subset of vertices in which every two vertices are connected by a permanent edge.
 R 6 (Clique Reduction).
-If there exists a vertex $u\in V$ such that $N_{p}(u)$ is a clique and $w(u)\geq\sum_{v\in N_{r}(u)}\max(0,p(u,v))+\max(0,\max_{v\in N_{p}(u)}\tilde{w}(v))$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus N_{p}[u]$ and $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$.
+If there exists a vertex $u\in V$ such that $N_{p}(u)$ is a clique of $G$ and $w(u)\geq\sum_{v\in N_{r}(u)}\max(0,p(\{u,v\}))+\max(0,\max_{v\in N_{p}(u)}\tilde{w}(v))$, then $u$ is in at least one set $I\in MGIS(G)$, $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus N_{p}[u]]$, and $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$ for each $v\in N_{r}(u)$.
 4.3 Low-degree Reductions
@@
 Let $u$ be a vertex with a single neighbor $v$.
-Case 1: $w(u)\geq p(u,v)$.
-(1) If $w(u)\geq 0$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G\setminus\{u\}$ and $w(v)\coloneqq w(v)-p(u,v)$.
-(2) If $w(u)<0$, then $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G\setminus\{u\}$, and $w(v)\coloneqq w(v)-p(u,v)+w(u)$.
-Case 2: $w(u)<p(u,v)$.
-(1) If $w(u)\geq 0$ and $w(u)\geq\tilde{w}(v)$, then $\alpha(G)=\alpha(G\setminus\{u,v\})+w(u)$.
-(2) If $w(u)\geq 0$ and $w(u)<\tilde{w}(v)$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G\setminus\{u\}$ and $w(v)\coloneqq w(v)-w(u)$.
-(3) If $w(u)<0$, then $\alpha(G)=\alpha(G\setminus\{u\})$.
+Case 1: $\{u,v\}\in E_{r}$ and $w(u)\geq p(\{u,v\})$.
+(1) If $w(u)\geq 0$, then $u$ is in at least one set $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G[V\setminus\{u\}]$ and $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$.
+(2) If $w(u)<0$, then $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G[V\setminus\{u\}]$ and $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})+w(u)$.
+$u$ is in at least one set $I\in MGIS(G)$ iff $v$ is in at least one set $I\in MGIS(G^{\prime})$.
+Case 2: $\{u,v\}\in E_{p}$ or $\{u,v\}\in E_{r}$ but $w(u)<p(\{u,v\})$.
+(1) If $w(u)\geq 0$ and $w(u)\geq\tilde{w}(v)$, then $u$ is in at least one set $I\in MGIS(G)$, $v\notin I$, and $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u,v\}]$.
+(2) If $w(u)\geq 0$ and $w(u)<\tilde{w}(v)$, then $u$ and $v$ cannot be both included in the same $I\in MGIS(G)$, we have $\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G[V\setminus\{u\}]$ and $w^{\prime}(v)\coloneqq w(v)-w(u)$.
+(3) If $w(u)<0$, then $u$ is not in any $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$.
 Note that R7 covers all degree-one vertices. In other words, any vertex of degree one can be removed using this rule.
 4.3.2 Degree-two Vertices Reductions
+For ease of describing $R8$ and $R9$, we assume $p(\{u,v\})=+\infty$ for each $\{u,v\}\in E_{p}$. This assumption does not affect the optimal objective value in the graph.
 R 8 (Permanently Connected Neighbors Reduction).
-Let $u$ be a vertex with exactly two neighbors $x,y$ and $p(x,y)=M$. The following table shows the reduction rules for reducing $u$ under different conditions.
+Let $u$ be a vertex with exactly two neighbors $x,y$ and $\{x,y\}\in E_{p}$. The following table shows the rules for reducing $u$ under different conditions.
 conditions reduction
-Case 1: $w(u)\geq\max(p(u,x),p(u,y))$ $w(u)\geq 0$
-R8.1.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G\setminus\{u\}$,
-$w(x)\coloneqq w(x)-p(u,x)$, $w(y)\coloneqq w(y)-p(u,y)$.
+Case 1: $w(u)\geq\max(p(\{u,x\}),p(\{u,y\}))$ $w(u)\geq 0$
+R8.1.1: $u$ is in at least one set $I\in MGIS(G)$,
+$\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G[V\setminus\{u\}]$,
+$w^{\prime}(x)\coloneqq w(x)-p(\{u,x\})$, $w^{\prime}(y)\coloneqq w(y)-p(\{u,y\})$.
 $w(u)<0$
-R8.1.2: $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G\setminus\{u\}$,
-$w(x)\coloneqq w(x)+w(u)-p(u,x)$, $w(y)\coloneqq w(y)+w(u)-p(u,y)$.
-Case 2: $p(u,x)>w(u)\geq p(u,y)$ w.l.o.g, assume $p(u,x)>p(u,y)$
+R8.1.2: $u$ is in at least one set $I\in MGIS(G)$ iff $x\in I$ or $y\in I$,
+$\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G[V\setminus\{u\}]$,
+$w^{\prime}(x)\coloneqq w(x)+w(u)-p(\{u,x\})$, $w^{\prime}(y)\coloneqq w(y)+w(u)-p(\{u,y\})$.
+Case 2: $p(\{u,x\})>w(u)\geq p(\{u,y\})$ w.l.o.g, assume $p(\{u,x\})>p(\{u,y\})$
 $w(u)\geq 0$ and
 $w(u)\geq\tilde{w}(x)$
-R8.2.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u,x\}$
-and $w(y)\coloneqq w(y)-p(u,y)$.
+R8.2.1: $u$ is in at least one set $I\in MGIS(G)$ and $x\notin I$,
+$\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u,x\}]$,
+and $w^{\prime}(y)\coloneqq w(y)-p(\{u,y\})$.
 $w(u)\geq 0$ and
 $w(u)<\tilde{w}(x)$
-R8.2.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$
-$w(x)\coloneqq w(x)-w(u)$, $w(y)\coloneqq w(y)-p(u,y)$.
+R8.2.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$,
+$w^{\prime}(x)\coloneqq w(x)-w(u)$, $w^{\prime}(y)\coloneqq w(y)-p(\{u,y\})$.
+$u$ is in at least one set $I\in MGIS(G)$ iff $x$ is not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
 $w(u)<0$
-R8.2.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$
-$w(y)\coloneqq w(y)+w(u)-p(u,y)$.
-Case 3: $\min(p(u,x),p(u,y))>w(u)$ w.l.o.g, assume $\tilde{w}(x)\geq\tilde{w}(y)$
+R8.2.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$,
+$w^{\prime}(y)\coloneqq w(y)+w(u)-p(\{u,y\})$.
+$u$ is in at least one set $I\in MGIS(G)$ iff $y$ is in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
+Case 3: $\min(p(\{u,x\}),p(\{u,y\}))>w(u)$ w.l.o.g, assume $\tilde{w}(x)\geq\tilde{w}(y)$
 $w(u)\geq 0$ and
 $w(u)\geq\tilde{w}(x)$
-R8.3.1: $\alpha(G)=\alpha(G\setminus\{u,x,y\})+w(u)$
+R8.3.1: $u$ is in at least one set $I\in MGIS(G)$,
+$\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u,x,y\}]$
 $w(u)\geq 0$ and
 $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$
-R8.3.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u,y\}$
-$w(x)\coloneqq w(x)-w(u)$.
+R8.3.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u,y\}]$,
+$w^{\prime}(x)\coloneqq w(x)-w(u)$.
+$u$ is in at least one set $I\in MGIS(G)$ iff $x$ is not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
 $w(u)\geq 0$
 and $w(u)<\tilde{w}(y)$
-R8.3.3: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$
-$w(x)\coloneqq w(x)-w(u)$, $w(y)\coloneqq w(y)-w(u)$.
-$w(u)<0$ R8.3.4: $\alpha(G)=\alpha(G\setminus\{u\})$.
-Figure 2: Examples of R8.
-Suppose that vertex $u$ is adjacent to $x$ and $y$ and, $(x,y)\in E_{p}$.
-$G_{R}$ represents the residual graph $G\setminus\{u,x,y\}$.
-$\alpha$ and $\alpha^{\prime}$ indicate the optimal objective value in the original graph and the graph after reduction, respectively.
+R8.3.3: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$,
+$w^{\prime}(x)\coloneqq w(x)-w(u)$, $w^{\prime}(y)\coloneqq w(y)-w(u)$.
+$u$ is in at least one set $I\in MGIS(G)$ iff both $x$ and $y$ are not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
+$w(u)<0$
+R8.3.4: $u$ is not in any $I\in MGIS(G)$, $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$.
 Examples of $R8$ are presented in Figure 2.
@@
 Then by $R8.1.1$, we can remove $u$ from the original graph $G$ (left) to get the reduced graph $G^{\prime}$(right) and $\alpha(G)=\alpha(G^{\prime})+10$.
-To describe the $R9$ reduction, we further define $p(u,v)=0$ for $(u,v)\notin E$. Note that this does not alter the objective value in the graph.
+For ease of describing $R9$-$R11$, we further assume $p(\{u,v\})=0$ for $\{u,v\}\notin E$.
+Note that this does not alter the objective value in the graph.
+Then, for every pair of vertices $\{u,v\}$, there is a corresponding penalty $p(\{u,v\})$.
 R 9 (Non-permanently Connected Neighbors Reduction).
-Let $u$ be a vertex with two neighbors $x,y$ and $p(x,y)<M$. W.l.o.g, assume $p(u,x)\geq p(u,y)$. The following table summarizes the reduction rules under different conditions.
+Let $u$ be a vertex with two neighbors $x$ and $y$.
+Assume $p(\{u,x\})\geq p(\{u,y\})$ without loss of generality.
+If $\{x,y\}\in E_{r}$ or $\{x,y\}\notin E$, the following table summarizes the reduction rules under different conditions.
 conditions reduction
-Case 1: $w(u)\geq p(u,x)$
+Case 1: $w(u)\geq p(\{u,x\})$
 $w(u)\geq 0$ and
-$w(u)\geq p(u,x)+p(u,y)$
-R9.1.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
-$w(x)\coloneqq w(x)-p(u,x)$, $w(y)\coloneqq w(y)-p(u,y)$
+$w(u)\geq p(\{u,x\})+p(\{u,y\})$
+R9.1.1: $u$ is in at least one set $I\in MGIS(G)$,
+$\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$,
+$w^{\prime}(x)\coloneqq w(x)-p(\{u,x\})$, $w^{\prime}(y)\coloneqq w(y)-p(\{u,y\})$
 $w(u)\geq 0$ and
-$w(u)<p(u,x)+p(u,y)$
-R9.1.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(x)\coloneqq w(x)-p(u,x)$,
-$w(y)\coloneqq w(y)-p(u,y)$, $p(x,y)\coloneqq p(x,y)-p(u,x)-p(u,y)+w(u)$
+$w(u)<p(\{u,x\})+p(\{u,y\})$
+R9.1.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
+$w^{\prime}(x)\coloneqq w(x)-p(\{u,x\})$, $w^{\prime}(y)\coloneqq w(y)-p(\{u,y\})$, $p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})-p(\{u,x\})-p(\{u,y\})+w(u)$
+$u$ is in at least one set $I\in MGIS(G)$ iff either $x$ or $y$ is not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
 $w(u)<0$
-R9.1.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(x)\coloneqq w(x)+w(u)-p(u,x)$,
-$w(y)\coloneqq w(y)+w(u)-p(u,y)$, $p(x,y)\coloneqq p(x,y)+w(u)$.
-Case 2: $p(u,x)>w(u)\geq p(u,y)$
+R9.1.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
+$w^{\prime}(x)\coloneqq w(x)+w(u)-p(\{u,x\})$, $w^{\prime}(y)\coloneqq w(y)+w(u)-p(\{u,y\})$, $p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})+w(u)$.
+$u$ is in at least one set $I\in MGIS(G)$ iff either $x$ or $y$ is in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
+Case 2: $p(\{u,x\})>w(u)\geq p(\{u,y\})$
 $w(u)\geq 0$ and
-$w(u)\geq p(u,x)+p(u,y)$
-R9.2.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(y)\coloneqq w(y)-p(u,y)$,
-$w(x)\coloneqq w(x)-w(u)$, $p(x,y)\coloneqq p(x,y)+p(u,x)-w(u)$.
+$w(u)\geq p(\{u,x\})+p(\{u,y\})$
+R9.2.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
+$w^{\prime}(y)\coloneqq w(y)-p(\{u,y\})$, $w^{\prime}(x)\coloneqq w(x)-w(u)$, $p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})+p(\{u,x\})-w(u)$.
+$u$ is in at least one set $I\in MGIS(G)$ iff either $x$ is not in or $y$ is in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
 $w(u)\geq 0$ and
-$w(u)<p(u,x)+p(u,y)$
-R9.2.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
-$w(y)\coloneqq w(y)-p(u,y)$, $w(x)\coloneqq w(x)-w(u)$, $p(x,y)\coloneqq p(x,y)-p(u,y)$.
+$w(u)<p(\{u,x\})+p(\{u,y\})$
+R9.2.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
+$w^{\prime}(y)\coloneqq w(y)-p(\{u,y\})$, $w^{\prime}(x)\coloneqq w(x)-w(u)$, $p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})-p(\{u,y\})$.
+$u$ is in at least one set $I\in MGIS(G)$ iff $x$ is not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
 $w(u)<0$ and
-$w(u)\geq p(u,x)+p(u,y)$
-R9.2.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
-$w(y)\coloneqq w(y)+w(u)-p(u,y)$, $p(x,y)\coloneqq p(x,y)+p(u,x)$.
+$w(u)\geq p(\{u,x\})+p(\{u,y\})$
+R9.2.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
+$w^{\prime}(y)\coloneqq w(y)+w(u)-p(\{u,y\})$, $p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})+p(\{u,x\})$.
+$u$ is in at least one set $I\in MGIS(G)$ iff $y$ is in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
 $w(u)<0$ and
-$w(u)<p(u,x)+p(u,y)$
-R9.2.4: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
-$w(y)\coloneqq w(y)+w(u)-p(u,y)$, $p(x,y)\coloneqq p(x,y)+w(u)-p(u,y)$.
-Case 3: $p(u,y)>w(u)$ $w(u)\geq 0$
-R9.3.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
-$w(x)\coloneqq w(x)-w(u)$, $w(y)\coloneqq w(y)-w(u)$, $p(x,y)\coloneqq p(x,y)-w(u)$.
+$w(u)<p(\{u,x\})+p(\{u,y\})$
+R9.2.4: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
+$w^{\prime}(y)\coloneqq w(y)+w(u)-p(\{u,y\})$, $p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})+w(u)-p(\{u,y\})$.
+$u$ is in at least one set $I\in MGIS(G)$ iff $x$ is not in and $y$ is in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
+Case 3: $p(\{u,y\})>w(u)$ $w(u)\geq 0$
+R9.3.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
+$w^{\prime}(x)\coloneqq w(x)-w(u)$, $w^{\prime}(y)\coloneqq w(y)-w(u)$, $p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})-w(u)$.
+$u$ is in at least one set $I\in MGIS(G)$ iff both $x$ and $y$ are not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
 $w(u)<0$ and
-$w(u)<p(u,x)+p(u,y)$
-R9.3.2: $\alpha(G)=\alpha(G\setminus\{u\})$.
+$w(u)<p(\{u,x\})+p(\{u,y\})$
+R9.3.2: $u$ is not in any $I\in MGIS(G)$, $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$.
 $w(u)<0$ and
-$w(u)\geq p(u,x)+p(u,y)$
-R9.3.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
-$p(x,y)\coloneqq p(x,y)-w(u)+p(u,x)+p(u,y)$.
-Figure 3: Examples of R9. Suppose that vertex $u$ is adjacent to vertices $x$ and $y$, and $(x,y)\notin E_{p}$.
+$w(u)\geq p(\{u,x\})+p(\{u,y\})$
+R9.3.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
+$p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})-w(u)+p(\{u,x\})+p(\{u,y\})$.
+$u$ is in at least one set $I\in MGIS(G)$ iff both $x$ and $y$ are in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
+Figure 2: Examples of $R8$.
+Suppose that vertex $u$ is adjacent to $x$ and $y$, and $\{x,y\}\in E_{p}$.
+$G_{R}$ represents the residual graph $G[V\setminus\{u,x,y\}]$.
+$\alpha$ and $\alpha^{\prime}$ indicate the optimal objective value in the original graph and the graph after reduction, respectively.
+Figure 3: Examples of $R9$. Suppose that vertex $u$ is adjacent to vertices $x$ and $y$, and $\{x,y\}\in E_{r}$ or $\{x,y\}\notin E$.
 Examples of reduction rule $R9$ are presented in Figure 3.
@@
 4.3.3 Low Permanent-degree Reductions
-We further introduce two reduction rules to reduce high-degree vertices that have only one or two permanent neighbors.
+We further introduce two reduction rules to reduce high-degree vertices that are connected by only one or two vertices by permanent edges.
 R 10 (Permanent Degree-one Vertices Reduction).
-If there exists a vertex $u$ that has exactly one permanent neighbor $x$ and $w(u)\geq\sum_{v\in N_{r}(u)}max(0,p(u,v))$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(x)\coloneqq w(x)-w(u)$, $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$, and $p(x,v)\coloneqq p(x,v)-p(u,v)$ for each $v\in N_{r}(u)\setminus N_{p}(x)$.
+If there exists a vertex $u$ such that $u$ has exactly one neighbor $x\in N_{p}(u)$ and $w(u)\geq\sum_{v\in N_{r}(u)}max(0,p(\{u,v\}))$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,v\}\mid v\in N_{r}(u)\setminus N_{p}(x)\})$, $w^{\prime}(x)\coloneqq w(x)-w(u)$, $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$ for each $v\in N_{r}(u)$, and $p^{\prime}(\{x,v\})\coloneqq p(\{x,v\})-p(\{u,v\})$ for each $v\in N_{r}(u)\setminus N_{p}(x)$.
+$u$ is in at least one set $I\in MGIS(G)$ iff $x$ is not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
 R 11 (Permanent Degree-two Vertices Reduction).
-Let $u$ be a vertex with exactly two permanent neighbors $x,y$, that $p(x,y)=M$, and $w(u)\geq\sum_{v\in N_{r}(u)}max(0,p(u,v))$. W.l.o.g, assume $\tilde{w}(x)\geq\tilde{w}(y)$.
-Case 1: If $w(u)\geq\tilde{w}(x)$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u,x,y\}$, and $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$.
-Case 2: If $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u,y\}$, $w(x)\coloneqq w(x)-w(u)$, $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$, and $p(x,v)\coloneqq p(x,v)-p(u,v)$ for each $v\in N_{r}(u)\setminus N_{p}(x)$.
-Case 3: If $w(u)<\tilde{w}(y)$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(x)\coloneqq w(x)-w(u)$, $w(y)\coloneqq w(y)-w(u)$, $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$, $p(x,v)\coloneqq p(x,v)-p(u,v)$ for each $v\in N_{r}(u)\setminus N_{p}(x)$, and $p(y,v)\coloneqq p(y,v)-p(u,v)$ for each $v\in N_{r}(u)\setminus N_{p}(y)$.
+Let $u$ be a vertex with exactly two neighbors $x,y\in N_{p}(u)$, $\{x,y\}\in E_{p}$, and $w(u)\geq\sum_{v\in N_{r}(u)}max(0,p(\{u,v\}))$. W.l.o.g, assume $\tilde{w}(x)\geq\tilde{w}(y)$.
+Case 1: If $w(u)\geq\tilde{w}(x)$, then $u$ must be included in at least one set $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u,x,y\}]$, and $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$ for each $v\in N_{r}(u)$.
+Case 2: If $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$, then $y$ is not in at least one set $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq(V\setminus\{u,y\},E(G[V\setminus\{u,y\}])\cup\{\{x,v\}\mid v\in N_{r}(u)\setminus N_{p}(x)\})$, $w^{\prime}(x)\coloneqq w(x)-w(u)$, $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$ for each $v\in N_{r}(u)$, and $p^{\prime}(\{x,v\})\coloneqq p(\{x,v\})-p(\{u,v\})$ for each $v\in N_{r}(u)\setminus N_{p}(x)$.
+$u$ is in at least one set $I\in MGIS(G)$ iff $x$ is not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
+Case 3: If $w(u)<\tilde{w}(y)$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,v\}\mid v\in N_{r}(u)\setminus N_{p}(x)\}\cup\{\{y,v\}\mid v\in N_{r}(u)\setminus N_{p}(y)\})$, $w^{\prime}(x)\coloneqq w(x)-w(u)$, $w^{\prime}(y)\coloneqq w(y)-w(u)$, $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$ for each $v\in N_{r}(u)$, $p^{\prime}(\{x,v\})\coloneqq p(\{x,v\})-p(\{u,v\})$ for each $v\in N_{r}(u)\setminus N_{p}(x)$, and $p^{\prime}(\{y,v\})\coloneqq p(\{y,v\})-p(\{u,v\})$ for each $v\in N_{r}(u)\setminus N_{p}(y)$.
+$u$ is in at least one set $I\in MGIS(G)$ iff both $x$ and $y$ are not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
 4.4 Vertex-pair Reductions
@@
 R 12 (Permanent Edge Reduction).
-If there is a permanent edge $(u,v)$ such that $w(u)\geq\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}[v])+\min(w^{+}(N_{r}(u)\setminus N_{p}(v)),\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p(u,x)))$, then $\alpha(G)=\alpha(G\setminus\{v\})$.
+If there is an edge $\{u,v\}$ such that $\{u,v\}\in E_{p}$ and $w(u)\geq\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}[v])+\min(w^{+}(N_{r}(u)\setminus N_{p}(v)),\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p\{u,x\}))$, then $v$ is not in at least one $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G[V\setminus\{v\}]$.
 R 13 (Common Neighbors Reduction).
-If there is a permanent edge $(u,v)$ such that $w(v)\geq w^{+}(N(v))-\max(0,w(u))$, then $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G\setminus(N_{p}(u)\cap N_{p}(v))$.
+If there is an edge $\{u,v\}$ such that $\{u,v\}\in E_{p}$ and $w(v)\geq w^{+}(N(v))-\max(0,w(u))$, then $N_{p}(u)\cap N_{p}(v)$ is not a subset of at least one $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G[V\setminus(N_{p}(u)\cap N_{p}(v))]$.
 R 14 (Twin Vertices Reduction).
-If there exist two vertices $u,v$ such that $N_{p}(u)=N_{p}(v)$, $w(u)\geq\sum_{x\in N_{r}(u)}\max(0,p(u,x))$ and $w(v)\geq\sum_{x\in N_{r}(v)}\max(0,p(v,x))$, then $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}$ is obtained by folding $u$ and $v$ into a vertex $v^{\prime}$, $w(v^{\prime})\coloneqq w(u)+w(v)-p(u,v)$, $p(u,x)\coloneqq M$ for each $x\in N_{p}(u)$, and $p(v^{\prime},x)\coloneqq p(u,x)+p(v,x)$ for each $x\in N_{r}(v)\cup N_{r}(u)$.
+If there exist two vertices $u,v$ such that $N_{p}(u)=N_{p}(v)$, $w(u)\geq\sum_{x\in N_{r}(u)}\max(0,p(\{u,x\}))$ and $w(v)\geq\sum_{x\in N_{r}(v)}\max(0,p(\{v,x\}))$, then $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}$ is obtained by folding $u$ and $v$ into a vertex $v^{\prime}$, $w(v^{\prime})\coloneqq w(u)+w(v)-p(\{u,v\})$, add edge $\{v^{\prime},x\}$ as a permanent edge in $G^{\prime}$ for each $x\in N_{p}(u)$, and add edge $\{v^{\prime},x\}$ as a removable edge in $G^{\prime}$ with $p^{\prime}(\{v^{\prime},x\})\coloneqq p(\{u,x\})+p(\{v,x\})$ for each $x\in N_{r}(v)\cup N_{r}(u)$.
+Both $u$ and $v$ belong to at least one $I\in MGIS(G)$ iff $v^{\prime}\in MGIS(G^{\prime})$.
 5 Reduction-driven Local Search
@@
 Output: the best solution $S^{*}$ found
-$S_{init}\leftarrow\varnothing$; /* initial an empty solution */
-$G,S_{init}\leftarrow\textit{Complete\_Reduction}(G,S_{init})$; /* Pre-process to obtain a kernel graph via $R1$-$R14$/
+$G,S_{init}\leftarrow\textit{Complete\_Reduction}(G)$; /* Pre-process to obtain a kernel graph and partial solution via $R1$-$R14$ */
 1 while cutoff time is not satisfied do
-       /* construct an initial solution using $R1$-$R14$/
+ /* construct an initial solution using $R1$-$R14$ */
 2 $S\leftarrow S_{init}\cup\textit{Random\_Peeling}(G)$;
@@
 The RLS begins by pre-processing the original graph $G$.
-It uses the Complete_Reduction procedure which applies all the reduction rules to obtain a kernel graph.
+It uses the Complete_Reduction procedure which applies all the reduction rules to obtain a kernel graph, that is, a graph where no further reduction rules can be applied.
 Due to the correctness of the reduction rules, solving the GIS problem in the original graph is equivalent to solving it in this kernel graph.
-The $S_{init}$ is a partial vertex set that must be in the optimal solution.
-The remaining steps of RLS follow the traditional iterated local search algorithm (Glover and Laguna (1998), Lourenço et al.
-(2019)): At each iteration, it supplements $S_{init}$ with a random generalized independent set obtained by Random_Peeling, then uses Neighborhood_Search_with_Partial_Reduction to improve the current solution (line 1)-1) .
+The $S_{init}$ is a partial vertex set that must be a subset of at least one optimal solution.
+The remaining steps of the RLS follow the traditional iterated local search algorithm (Glover and Laguna (1998), Lourenço et al.
+(2019)): At each iteration, it supplements $S_{init}$ with a random generalized independent set obtained by Random_Peeling, then uses Neighborhood_Search_with_Partial_Reduction to improve the current solution (line 1-1) .
 The current solution is denoted by $S$ and the best-known solution visited by the iterative search is kept in $S^{*}$ (line 1).
 The iteration continues until the cutoff time limit is met.
-Compared with traditional local search, RLS seamlessly integrates reduction techniques into the pre-processing, the initial solution building, and the neighborhood searching procedures.
-Input: a graph $G=(V,E)$ with vertex profits and edge penalties, initial solution $S\subseteq V$
+Compared with traditional local search, the RLS seamlessly integrates reduction techniques into the pre-processing, the initial solution building, and the neighborhood searching procedures.
+Input: a graph $G=(V,E)$ with vertex profits and edge penalties
 Output: updated graph $G$ and partial solution $S$
+$S\leftarrow\varnothing$; /* initial an empty solution */
 /* initial candidate sets for 14 reduction rules */
@@
 7    for each $elem\in\text{Cand}_{i}$ do
-8                   if the condition of $Ri$ is satisfied by checking $elem$ then
-9                         change the structure, profits and penalties of $G$ and set $S$ according to $Ri$;
-10                         update the Candj if $Rj$ reachable from $Ri$ in Figure 4;
+8       if the condition of $Ri$ is satisfied by checking elemelem then
+9            change the structure, profits and penalties of $G$ and set $S$ according to RiRi;
+10            update the Candj if RjRj reachable from $Ri$ in Figure 4;
 11
-12                  Remove $elem$ from Candi;
+12        Remove elemelem from Candi;
 return $G,S$;
@@
 5.1 Complete Reduction as Pre-processing
-The Complete_Reduction algorithm exhaustively applies the 14 reduction rules in Section 4.
+The Complete_Reduction algorithm exhaustively applies the 14 reduction rules described in Section 4.
 A naive implementation is to check every rule independently, process the graph when applicable, and stop when no rules apply.
-The times of checking all these rules is $O(14(|V|+|E|))$ as it requires scanning the entire graph for each rule.
+The times for checking all these rules are bounded by $O(|V|+|E|)$ as it requires scanning the entire graph for each rule.
 Given the large size of real-world graphs, this is inefficient in practice.
-We propose a more efficient implementation which incrementally detects applicable rules.
+We propose a more efficient implementation that incrementally detects applicable rules.
 Specifically, we observe that rules are interdependent, with one application triggering others.
-For example, if only the condition of $R1$ holds in graph $G$ (i.e., there exists an edge $(u,v)$ with $p(u,v)=0$), the edge $(u,v)$ is removed, making R3-15 potentially applicable but not the other rules.
+For example, if only the condition of $R1$ is satisfied in graph $G$ (i.e., there exists an edge $\{u,v\}\in E_{r}$ with $p(\{u,v\})=0$), then the edge $\{u,v\}$ is removed, potentially enabling the application of R3R3-$14$, while the remaining rules remain inapplicable.
 Hence, we make an in-depth analysis of this dependency relationship for all these rules, and we show the result in Figure 4.
@@
 It takes a graph $G$ as its input and outputs a kernel graph and a set $S\subseteq V$.
-In detail, it initializes a special candidate set Candi which contains either vertices, edges, or vertex pairs, for a rule $Ri$ where $i=1,...,14$.
+In detail, it initializes a special candidate set Candi, which contains either vertices, edges, or vertex pairs, for a rule $Ri$ where i=1,,14i=1,...,14.
 Lines 1-4 show how Candi is initialized.
 When an element $elem\in\text{Cand}_{i}$ satisfies the condition of $Ri$, the current graph $G$ is changed and the current set $S$ is updated according to $Ri$.
-Due to the change of graph $G$, if another reduction rule $Rj$ can be applied (i.e.
-$Rj$ is reachable from $Ri$ in the dependency graph in Figure 4), then the vertices, edges, or vertex pairs which are “neighbor elements” of $elem$ are moved into Candj.
-Consequently, the algorithm checks whether each reduction rule is satisfied by tracking the candidate set in an incremental manner.
+Along with the change of graph $G$, if another reduction rule $Rj$ can be applied (i.e., $Rj$ is reachable from $Ri$ in the dependency graph in Figure 4), then the vertices, edges, or vertex pairs which are “neighbor elements” of elemelem are moved into Candj.
+The process continues recursively until all candidate sets are empty.
+Therefore, the algorithm checks whether each reduction rule is satisfied by tracking the candidate set in an incremental manner.
+In our implementations, $R1$ and $R2$ are prioritized because the removal of edges may enable more reduction rules.
+$R3$-$R6$ are assigned secondary priority, as they directly identify vertices that are in at least one set in MGIS(G), which may further shrink the candidate set.
+$R7$-$R11$ are of the third priority because they can identify vertices that are in at least one set in MGIS(G) in some cases.
+$R12$-$R14$ are put at the last because they are not able to identify such vertices.
 Figure 4: The dependency graph of the reduction rules.
@@
 A directed edge on the right side indicates what reduction rules can be satisfied due to this change.
-5.2 Random Peeling for Initial Solutions
-Random-peeling is a simple random procedure for generating an initial solution.
-It starts with an empty vertex set $S$.
-Then when $V\setminus N_{p}(S)$ is not empty (meaning that $S$ can possibly be a larger generalized independent set), a vertex $u$ is picked randomly from $V\setminus N_{p}(S)$ and $u$ is moved into $S$.
-Due to the removal of $u$ from $G$, the Complete_Reduction is used again to further shrink the remaining graph $G$ (line 3).
-When $V\setminus N_{p}(S)$ becomes empty, we further extend $S$ by randomly picking a vertex $v$ such that $nb(S\cup\{v\})>nb(S)$.
-The $S$ is returned when it becomes a maximal generalized independent set.
 Input: a graph $G=(V,E)$ with vertex profits and edge penalties
@@
 4
-5      $G,S\leftarrow\textit{Complete\_Reduction}(G\setminus\{u\},S)$;
+5 $G,S\leftarrow\textit{Complete\_Reduction}(G[V\setminus\{u\}],S)$;
 6extend $S$ to be a maximal generalized independent set;
@@
 Algorithm 3 The Random_Peeling algorithm
+5.2 Random Peeling for Initial Solutions
+Random-peeling is a simple random procedure for generating an initial solution.
+It starts with an empty vertex set $S$.
+When $V\setminus N_{p}[S]$ is not empty (meaning that $S$ can possibly be a larger generalized independent set), a vertex $u$ is picked randomly from $V\setminus N_{p}[S]$ and $u$ is moved into $S$.
+After the removal of $u$ from $G$, the Complete_Reduction is used again to further shrink the remaining graph $G$ (line 3).
+When $V\setminus N_{p}[S]$ becomes empty, we further extend $S$ by randomly picking a vertex $v$ from the randomly removed vertices such that $v\notin N_{p}[S]$ and $nb(S\cup\{v\})>nb(S)$.
+The $S$ is returned when there is no other generalized independent set $S^{\prime}$ such that $S\subset S^{\prime}$ and $nb(S^{\prime})\geq nb(S)$.
+Clearly $S$ is a maximal generalized independent set.
 5.3 Neighborhood Search
-To improve the incumbent solution, we use the iterated tabu search with three move operators.
-Move operators.
-In RLS, there are three move operators that update the current solution: ADD, SWAP, and the novel REDUCE operator based on two reduction rules.
-Given the current solution $S\subseteq V$, the ADD$(v)$ operator displaces a vertex $v$ (if any) from $V\setminus N_{p}[S]$ to the $S$ and the SWAP$(u,v)$ operator adds a vertex $v\in N_{p}(S)$ if $|N_{p}(v)\cap S|=1$ and removes the only vertex $u\in N_{p}(v)\cap S$ from the current solution $S$.
-Indeed, ADD and SWAP are often used in existing local search methods such as Zhou and Hao (2017), Jin and Hao (2015).
-The REDUCE$(V\setminus N_{p}[S])$ operator applies $R3$ and $R4$ to include candidate vertices that are definitely in a set in $MGIS(G)$.
-Input: a graph $G=(V,E)$ with vertex profits and edge penalties, initial solution $S$, search depth $L\in\mathbb{N}$, perturbation parameter $\epsilon\in(0,1)$
+To improve the current solution, we use an iterated tabu search with three move operators.
+[…]
+Input: a graph $G=(V,E)$ with vertex profits and edge penalties, initial solution $S$, search tolerance $L\in\mathbb{N}$, perturbation parameter $\epsilon\in(0,1)$
 Output: a solution $S$ to the GIS problem
@@
 2 while True do
-3       $M_{1}\leftarrow V\setminus N_{p}[S]$, $M_{2}\leftarrow\{v\in N_{p}(S)\mid u\in S,(u,v)\in E_{p},|N_{p}(v)\cap S|=1\}\setminus tabu\_list$
+3 $M_{1}\leftarrow V\setminus N_{p}[S]$, $M_{2}\leftarrow\{v\in N_{p}(S)\mid u\in S,\{u,v\}\in E_{p},|N_{p}(v)\cap S|=1\}\setminus tabu\_list$
 4 $v\leftarrow$ a vertex with the largest move gain from $M_{1}\cup M_{2}$;
-       if $v\in M_{1}$ then  $S\leftarrow S\cup\{v\}$, $M_{1}\leftarrow M_{1}\setminus\{v\}$ /* add $v$ into $S$/
+ if $v\in M_{1}$ then $S\leftarrow S\cup\{v\}$, $M_{1}\leftarrow M_{1}\setminus\{v\}$ /* add $v$ into $S$ */
 5 else if $v\in M_{2}$ then
-             $u\leftarrow N_{p}(v)\cap S$, $S\leftarrow S\cup\{v\}\setminus\{u\}$; /* Swap $v$ and $u$/
-6             add $u$ into $tabu\_list$ with tabu tenure $T_{u}=10+random(0,|M_{2}|)$
-       $REDUCE(M_{1})$ and reassign $S$; /* apply $R3$ and $R4$/
+    $u\leftarrow N_{p}(v)\cap S$, $S\leftarrow S\cup\{v\}\setminus\{u\}$; /* Swap $v$ and $u$ */
+6      add $u$ into tabu_listtabu\_list with tabu tenure $T_{u}=10+random(0,|M_{2}|)$
+ $REDUCE(M_{1})$ and reassign $S$; /* apply $R3$ and $R4$ */
 7 if $nb(S)>nb(S^{*})$ then $S^{*}\leftarrow S$, $l\leftarrow 0$ else $l\leftarrow l+1$
@@
 Fast Calculation of Move Gain.
-For each vertex $u\in V$, we maintain $B_{u}=w(u)-\sum_{v\in S,(u,v)\in E}p(u,v)$.
-Then, the move gain of $ADD(v)$, which is the gain of the net benefit after applying a $ADD(v)$, is $B_{v}$, and the move gain of $SWAP(u,v)$ is $B_{v}-B_{u}$.
-After an $ADD(v)$ move, for each $u\in N_{r}(v)$, we update $B_{u}$ as $B_{u}\coloneqq B_{u}-p(u,v)$ After a $SWAP(u,v)$ move, for each $x\in N_{r}(u)$, $B_{x}\coloneqq B_{x}+p(x,u)$ and for each $y\in N_{r}(v)$, $B_{y}\coloneqq B_{y}-p(y,v)$.
-The time complexity of performing $ADD$ or $SWAP$ is $O(\Delta)$, while the REDUCE operator operates exactly.
+For each vertex $u\in V$, we maintain $B_{u}=w(u)-\sum_{v\in S,\{u,v\}\in E_{r}}p(\{u,v\})$.
+Then, the move gain of ADD(v)ADD(v)—the gain of the net benefit after applying a $ADD(v)$, is $B_{v}$, and the move gain of $SWAP(u,v)$ is $B_{v}-B_{u}$.
+After an $ADD(v)$ move, for each $u\in N_{r}(v)$, we update $B_{u}$ as $B_{u}\coloneqq B_{u}-p(\{u,v\})$.
+After a $SWAP(u,v)$ move, for each $x\in N_{r}(u)$, $B_{x}\coloneqq B_{x}+p(\{x,u\})$ and for each $y\in N_{r}(v)$, $B_{y}\coloneqq B_{y}-p(\{y,v\})$.
+The time complexity of performing $ADD$ or $SWAP$ is $O(\Delta)$, and of REDUCE is $O(|E|)$.
 The overall complexity of applying all three operators is $O(|E|)$.
-Notably, the REDUCE operator is not specific to the GIS problem and can enhance the efficiency of local search in other optimization problems.
-Neighborhood Search with Tabu and Perturbation. The neighborhood search in RLS iteratively performs the three basic operators above to improve the current solution $S$, as shown in Algorithm 4.
+Notably, the REDUCE operator is not limited to the GIS problem and can enhance the efficiency of local search in other optimization problems.
+Neighborhood Search with Tabu and Perturbation. The neighborhood search in the RLS iteratively performs the three basic operators above to improve the current solution $S$, as shown in Algorithm 4.
 At each iteration in lines 4-4, it first builds $M_{1}$, the set of vertices that can be added into $S$ by $ADD$, and $M_{2}$, the set of vertices that can be swapped into $S$ by $SWAP$.
-By evaluating $M_{1}$ and $M_{2}$, the $ADD$ or $SWAP$ with the largest move gain is applied, as presented in lines 4-8.
-After performing the ADD or SWAP operator, we additionally use the REDUCE operator to update $M_{1}$.
-This decreases the search space of the $ADD$ operator because $V\setminus N_{p}(S)$ is much larger than $S$ in practice.
+By evaluating $M_{1}$ and $M_{2}$, $ADD$ or $SWAP$ with the largest move gain is applied, as presented in lines 4-8.
+After performing $ADD$ or $SWAP$, we additionally use the REDUCE operator to add more vertices to $S$.
+This narrows the search space $M_{1}=V\setminus N_{p}[S]$ of $ADD$.
 To avoid short-term search cycles, a tabu list is used to exclude recently visited vertices for successive iterations (line 4).
@@
 Specifically, when we detect that no operators can be applied or the best solution remains unchanged after a certain number of iterations, we remove a portion of low-contribution vertices.
-The whole neighborhood search terminates when the consecutive number of iterations without improvement reaches the search depth parameter $L$ (line 4).
+The whole neighborhood search terminates when the consecutive number of iterations without improvement reaches the search tolerance parameter $L$ (line 4).
 6 Computational Experiments
 This section is dedicated to a comprehensive evaluation of the proposed reduction rules and heuristic approaches.
-In order to evaluate the algorithm performance, we compare both solution quality and runtime between RLS and other benchmark algorithms; to evaluate the effect of reduction rules, we compare the algorithms with reductions and without reductions; we also conduct an ablation study to evaluate the effect of each reduction rule.
+To evaluate the algorithm performance, we compare both solution quality and runtime between the RLS and other benchmark algorithms.
+To evaluate the effect of the reduction rules, we compare the algorithms with and without reductions; we also conduct an ablation study to evaluate the effect of each reduction rule.
 6.1 Datasets
-The proposed reduction rules and the RLS algorithm are evaluated on three sets and a total of $278=216+47+15$ instances, covering various application scenarios.
+The proposed reduction rules and the RLS algorithm are evaluated on three sets of $216+47+15=278$ instances, covering various application scenarios.
 SET-1.
@@
 (2024).
-This set of instances consists of 12 graphs with 125 to 400 vertices and 6963 to 71820 edges (available at https://or-dii.unibs.it/instances/Instances_GISP.zip).
+This set of instances consists of 12 graphs with 125 to 400 vertices and 6,963 to 71,820 edges (available at https://or-dii.unibs.it/instances/Instances_GISP.zip).
 SET-2.
@@
 To further evaluate the algorithms on large instances and diverse real-world scenarios, we introduce 15 instances with up to 18 million vertices and 261 million edges.
-The first four smaller instances represent ecosystem and road/facility networks, while the remaining 11 larger instances are drawn from hyperlink, citation, and social networks.
+The first 4 smaller instances represent ecosystem and road/facility networks, while the remaining 11 larger instances are drawn from hyperlinks, citations, and social networks.
 The instances were generated using the same method for generating SET-2 (available at https://github.com/PlutoAiyi/RLS/tree/main).
 6.2 Experimental Settings
-All experiments are conducted on a machine with an Intel(R) Xeon(R) Platinum 8360Y CPU @ 2.40GHz and 2T RAM.
-Benchmark Algorithms. To evaluate the effectiveness of our proposed reduction rules and RLS algorithm, we compare them with the best-performing algorithms for the GIS problem.
+All experiments are conducted on a machine with an Intel(R) Xeon(R) Platinum 8360Y CPU @ 2.40GHz and 2TB RAM.
+Benchmark Algorithms. To evaluate the effectiveness of our proposed reduction rules and the RLS algorithm, we compare them with the best-performing algorithms for the GIS problem.
   • CB&B: The exact branch-and-bound algorithm using the UBQP formulation proposed by Hosseinian and Butenko (2019).
   • LA-B&B: The exact branch-and-bound algorithm using an Lagrange relaxation techniques by Zheng et al. (2024).
+  • CPLEX: The mixed integer programming solver with version 22.12 (released on December 2024).
   • ILS-VND: The iterative local search algorithm proposed by Nogueira et al. (2021).
-  • ALS: The adaptive multi-neighborhood local search algorithm proposed by Zheng et al. (2024)
+  • ALS: The adaptive local search algorithm proposed by Zheng et al. (2024)
   • RLS: Our proposed reduction-driven local search algorithm. The algorithm is implemented in C++.
-The source codes of ILS-VND, LA-B&B and ALS were made available by the authors. All source codes are available at https://github.com/PlutoAiyi/RLS.
+The source codes of CB&B, LA-B&B, ILS-VND and ALS were made available by the authors. All source codes are available at https://github.com/PlutoAiyi/RLS.
 Parameter Settings.
-The RLS algorithm requires two parameters: the search depth $L$ and the perturbation parameter $\epsilon\in(0,1)$.
+The RLS algorithm requires two parameters: the search tolerance $L$ and the perturbation parameter $\epsilon\in(0,1)$.
 We set $L=10|V|$ where $|V|$ is the number of vertices of the input graph and $\epsilon=0.2$.
-This setting is obtained using the general IRACE automatic parameter configuration tool (López-Ibáñez et al.
+This setting was obtained using the general IRACE automatic parameter configuration tool (López-Ibáñez et al.
 2016).
-For large instances like soc-pokec, LiveJ and uk2002, we decrease the search depth as $L=0.5|V|$ to encourage explorations within the cutoff time.
+For large instances like soc-pokec, LiveJ and uk2002, we decrease the search tolerance as $L=0.5|V|$ to encourage explorations within the cutoff time.
 6.3 Performance Comparison with Existing Algorithms
 We first compared five algorithms on the SET-1 instances.
-The exact algorithms CB&B and LA-B&B ran with a cutoff time of 3 hours (10800 seconds) for each instance.
-If an exact algorithm did not finish within the cutoff time, then we reported the best solution found.
+The two branch & bound algorithms CB&B and LA-B&B ran with a cutoff time of 3 hours (10,800 seconds) for each instance.
+If an exact algorithm fails to finish within the cutoff time, then we report the best solution found so far.
 The remaining three heuristic algorithms ran 10 trials with a cutoff time of 30 seconds for each instance.
@@
 In Table 2, we summarize the computational results on SET-1.
-In the last two columns, we show the number of instances where the corresponding algorithm performs best (#Best) and worst (#Worst) among all the five algorithms, in terms of the best known objective value and runtime, respectively.
-Note that the runtime of RLS refers to the wall-clock time, including both pre-processing and local search.
+In the last two columns, we report the number of instances where each algorithm achieves the best (#Best) and worst (#Worst) performance, with respect to the best-known objective value and runtime, respectively.
+Note that the runtime of the RLS refers to the wall-clock time, including both pre-processing and local search.
 We observe that, in all instances, the best-known objective values obtained by local search algorithms are comparable to those obtained by exact algorithms.
-In particular, if an exact algorithm achieves the optimal solution within 3 hours, local search algorithms can also find an optimal solution (although without proof).
-However, when neither of the exact algorithms finds the optimal solution, local search algorithms always produce a better one.
-Besides, the exact LA-B&B dominates CB&B whereas the heuristic ALS and RLS dominate ILS-VND in terms of both solution quality and runtime.
-Therefore, we only consider LA-B&B, ALS, and RLS for even larger instances in SET-2 and SET-3.
-Detailed results on this set are left in Table LABEL:tab:_Set1_detail in Appendix B.
+In particular, if an exact algorithm obtains the optimal solution within 3 hours, local search algorithms are also able to find an optimal solution.
+However, when neither of the exact algorithms finds the optimal solution, local search algorithms always produce a better solution.
+Besides, the exact LA-B&B dominates CB&B whereas the heuristics ALS and the RLS dominate ILS-VND in terms of both solution quality and runtime.
+Therefore, we only consider LA-B&B, ALS, and the RLS for even larger instances in SET-2 and SET-3.
+Detailed results on SET-1 are shown in Table LABEL:tab:_Set1_detail in Appendix B.
 Table 2: Summary of comparative results on SET-1 instances
@@
 #Total #Best #Worst #Best #Worst
-Exact CB&B 216 126 88 78 138
+Branch & Bound CB&B 216 126 88 78 138
 LA-B&B 216 155 3 216 0
@@
 Table 3: Comparative results of RLS on SET-2 instances
-Instance $|V|$ $|E_{p}|$ $|E_{r}|$ LA-B&B ALS RLS
+Instance $|V|$ $|E_{p}|$ $|E_{r}|$ CPLEX LA-B&B ALS RLS
 […]
-bio-yeast 1,458 980 968 66,545 300 68,572 9 68,574 $<$1 0 $<$1
-soc-wiki-Vote 889 727 2,187 35,973 300 37,358 8 37,358 $<$1 0 $<$1
-vc-exact_131 2,980 1,191 4,169 107,812 300 109,392 14 109,641 $<$1 0 $<$1
-web-edu 3,031 4,813 1,661 109,856 300 113,781 16 114,140 $<$1 98 $<$1
-MANN_a81 3,321 3,118 3,362 114,616 300 116,907 14 117,327 $<$1 0 $<$1
-tech-routers-rf 2,113 1,539 5,093 93,690 300 97,799 11 97,799 $<$1 0 $<$1
-vc-exact_039 6,795 2,574 8,046 254,941 300 258,137 9 259,542 $<$1 0 $<$1
-ca-GrQc 4,158 3,179 10,243 167,846 300 174,654 15 174,782 11 282 $<$1
-vc-exact_038 786 10,321 3,703 10,371 300 12,014 12 12,015 4 697 $<$1
-vc-exact_078 11,349 4289 13,450 102,312 300 435,412 $<$1 438,671 $<$1 0 $<$1
-vc-exact_087 13,590 5126 16,114 71,804 300 517,500 2 518,500 $<$1 0 $<$1
-vc-exact_107 6,402 15,646 5,594 88,602 300 400,223 1 407,878 $<$1 0 $<$1
-vc-exact_151 15,783 12,144 12,519 54,114 300 522,462 1 530,138 $<$1 0 $<$1
-vc-exact_167 15,783 18,235 6,428 55,190 300 457,272 2 465,178 $<$1 0 $<$1
-bio-dmela 7,393 18,897 6,672 280,897 300 302,449 10 302,992 $<$1 44 $<$1
-vc-exact_011 9,877 19,102 6,871 190,729 300 301,789 7 303,793 $<$1 172 $<$1
-web-spam 4,767 8,089 28,386 180,446 300 190,344 14 190,459 13 487 $<$1
-vc-exact_026 6,140 27,159 9,608 177,898 300 201,316 17 201,503 $<$1 300 $<$1
-vc-exact_001 6,160 19,655 20,552 190,729 300 215,668 13 215,941 $<$1 134 $<$1
-vc-exact_024 7,620 11,367 35,926 226,540 300 235,440 14 236,018 12 1,255 $<$1
-C1000.9 1,000 11,989 37,432 6,102 300 7,229 12 7,237 10 1,000 $<$1
-tech-WHOIS 7,476 42,181 14,762 309,903 300 325,327 14 325,581 13 338 $<$1
-p-hat700-3 700 30,398 31,242 2,937 300 4,473 $<$1 4,473 $<$1 700 $<$1
-vc-exact_008 7,537 35,678 37,155 71,804 300 246,771 14 247,091 $<$1 574 $<$1
-keller5 776 17,800 56,910 2,788 300 3,877 1 3,877 2 776 $<$1
-vc-exact_194 1,150 39,520 41,331 4,273 300 5,419 $<$1 5,419 $<$1 1,150 $<$1
-hamming10-4 1,024 66,369 23,231 2,871 300 3,660 $<$1 3,660 $<$1 1,024 $<$1
-brock800_2 800 26,629 84,805 1,996 300 2,473 $<$1 2,473 $<$1 800 $<$1
-brock800_4 800 26,831 85,126 1,968 300 2,501 $<$1 2,501 $<$1 800 $<$1
-ca-HepPh 11,204 57,721 59,898 153,907 300 370,779 6 372,079 17 2,165 $<$1
-p-hat700-2 700 90,788 32,134 2127 300 2,971 $<$1 2,971 $<$1 700 $<$1
-vc-exact_196 1534 93,273 32,809 4273 300 5,321 $<$1 5,321 $<$1 1,534 $<$1
-p-hat700-1 700 135,804 47,847 844 124 844 $<$1 844 $<$1 700 $<$1
-ca-AstroPh 17,903 47,135 149,837 44,296 300 602,320 2 603,945 14 4,985 1
-C2000.9 2,000 147,548 51,920 5477 300 6396 3 6396 3 2,000 $<$1
-DSJC1000.5 1,000 59,694 189,980 1492 300 1861 1 1861 $<$1 1,000 $<$1
-socfb-MIT 6,402 123,064 128,166 110,677 300 129,807 17 129,872 16 5,494 $<$1
-p-hat1500-3 1,500 66,206 210,800 3856 300 7,501 14 7,501 16 1,500 $<$1
-socfb-UCSB37 14,917 356,918 125,297 219,965 300 254,965 16 255,898 22 13,655 3
-socfb-Duke14 9,885 374,731 131,706 125,940 300 151,169 15 151,326 18 9,003 4
-p-hat1500-2 1,500 411,372 143,918 2,444 300 4,253 $<$1 4,253 $<$1 1,500 $<$1
-socfb-Stanford3 11,586 136,823 431,486 215,381 300 251,350 12 251,833 14 9,380 2
-socfb-UConn 17,206 448,038 156,829 170,548 300 274,637 12 275,034 20 16,025 5
-p-hat1500-1 1,500 411,744 427,583 1,121 300 1,287 $<$1 1,287 1 1,500 $<$1
-C2000.5 2,000 489,894 509,270 1,570 300 1,849 2 1,849 6 2,000 1
-keller6 3,361 759,816 266,766 3,545 300 5,236 10 5,236 16 3,361 $<$1
-C4000.5 4,000 1,957,884 2,039,848 1,601 300 1,948 20 1,950 22 4,000 9
-# of best / # of instances 1/47 18/47 47/47
-average running time(s) 296.3 7.9 5.9 1.4
+bio-yeast 1,458 980 968 68,574 1,111 <<1 66,545 1,062 2 68,572 1,113 9 68,574 1,111 <<1 0 <<1
+soc-wiki-Vote 889 727 2,187 37,358 648 <<1 35,973 615 <<1 37,358 648 8 37,358 648 <<1 0 <<1
+vc-exact_131 2,980 1,191 4,169 109,641 2,098 <<1 107,812 2,045 14 109,392 2,121 14 109,641 2,098 <<1 0 <<1
+web-edu 3,031 4,813 1,661 114,140 1,752 <<1 109,856 1,644 10 113,781 1,746 16 114,140 1,753 <<1 98 <<1
+MANN_a81 3,321 3,118 3,362 117,327 2,084 <<1 114,616 1,999 16 116,907 2,125 14 117,327 2,084 <<1 0 <<1
+tech-routers-rf 2,113 1,539 5,093 97,799 1,668 <<1 93,690 1,588 54 97,799 1,668 11 97,799 1,667 <<1 0 <<1
+vc-exact_039 6,795 2,574 8,046 259,542 5,018 <<1 254,941 4,908 253 258,137 5,164 9 259,542 5,018 <<1 0 <<1
+ca-GrQc 4,158 3,179 10,243 174,782 2,991 9 167,846 2,871 45 174,654 2,997 15 174,782 2,994 11 282 <<1
+vc-exact_038 786 10,321 3,703 12,015 208 589 10,406 168 1,994 12,014 208 12 12,015 207 4 697 <<1
+vc-exact_078 11,349 4,289 13,450 438,671 8,330 <<1 431,371 8,128 1,074 435,412 8,671 <<1 438,671 8,330 <<1 0 <<1
+vc-exact_087 13,590 5,126 16,114 518,500 10,025 2 509,734 9,809 2,319 517,500 10,083 2 518,500 10,026 <<1 0 <<1
+vc-exact_107 6,402 15,646 5,594 407,878 6,763 <<1 396,456 6,488 10,646 400,223 6,846 1 407,878 6,763 <<1 0 <<1
+vc-exact_151 15,783 12,144 12,519 530,138 9,526 2 516,445 9,135 2,402 522,462 9,764 1 530,138 9,526 <<1 0 <<1
+vc-exact_167 15,783 18,235 6,428 465,178 7,797 1 449,579 7,438 10,469 457,272 7,923 2 465,178 7,797 <<1 0 <<1
+bio-dmela 7,393 18,897 6,672 302,992 4,952 <<1 280,897 4,474 170 302,449 4,959 10 302,992 4,951 <<1 44 <<1
+vc-exact_011 9,877 19,102 6,871 303,793 5,369 1 286,137 4,950 4,818 301,789 5,487 7 303,793 5,387 <<1 172 <<1
+web-spam 4,767 8,089 28,386 190,507 3,285 4.61 180,446 3,130 53 190,344 3,299 14 190,459 3,289 13 487 <<1
+vc-exact_026 6,140 27,159 9,608 201,509 3,887 1 177,902 3,372 668 201,316 3,922 17 201,503 3,900 <<1 300 <<1
+vc-exact_001 6,160 19,655 20,552 215,941 4,129 2 194,984 3,709 9,575 215,668 4,175 13 215,941 4,128 <<1 134 <<1
+vc-exact_024 7,620 11,367 35,926 236,109 4,632 10,783 226,540 4,438 244 235,440 4,735 14 236,018 4,659 12 1,255 <<1
+C1000.9 1,000 11,989 37,432 5,263 99 10,800 6,104 116 430 7,229 126 12 7,239 125 10 1,000 <<1
+tech-WHOIS 7,476 42,181 14,762 325,584 5,368 5 309,903 5,050 218 325,327 5,367 14 325,581 5,371 13 338 <<1
+p-hat700-3 700 30,398 31,242 4,025 68 10,800 3,030 52 7,436 4,473 78 <<1 4,473 78 <<1 700 <<1
+vc-exact_008 7,537 35,678 37,155 247,091 4,800 3 210,126 4,089 9,093 246,771 4,839 14 247,091 4,810 <<1 574 <<1
+keller5 776 17,800 56,910 2,373 58 10,800 2,860 54 6,327 3,877 69 1 3,877 69 2 776 <<1
+vc-exact_194 1,150 39,520 41,331 3,734 77 10,708 4,386 73 1,324 5,419 86 <<1 5,419 86 <<1 1,150 <<1
+hamming10-4 1,024 66,369 23,231 3,256 52 8,108 3,073 45 2,069 3,660 54 <<1 3,660 54 <<1 1,024 <<1
+brock800_2 800 26,629 84,805 1,177 49 10,800 1,996 36 10 2,473 44 <<1 2,473 43 <<1 800 <<1
+brock800_4 800 26,831 85,126 1,583 48 10,800 1,979 39 7,215 2,501 43 <<1 2,501 42 <<1 800 <<1
+ca-HepPh 11,204 57,721 59,898 372,096 6,094 10,800 348,738 5571 422 370,779 6,083 6 372,079 6,098 17 2,165 <<1
+p-hat700-2 700 90,788 32,134 2,870 41 10,140 2,345 37 2,078 2,971 42 <<1 2,971 42 <<1 700 <<1
+vc-exact_196 1,534 93,273 32,809 3,916 62 10,800 4,273 57 66 5,321 76 <<1 5,321 76 <<1 1,534 <<1
+p-hat700-1 700 135,804 47,847 730 11 10,800 844 11 5 844 11 <<1 844 11 <<1 700 <<1
+ca-AstroPh 17,903 47,135 149,837 604,288 10,468 10,800 558,736 9,752 2,544 602,320 10,521 2 604,019 10,491 14 4,985 1
+C2000.9 2,000 147,548 51,920 3,958 65 9,489 5,477 73 11 6,396 86 3 6,396 86 3 2,000 <<1
+DSJC1000.5 1,000 59,694 189,980 925 39 7,645 1,515 30 694 1,861 32 1 1,861 32 <<1 1,000 <<1
+socfb-MIT 6,402 123,064 128,166 7,098 124 10,800 110,677 1,779 20 129,807 2,142 17 129,917 2,132 16 5,494 <<1
+p-hat1500-3 1,500 66,206 210,800 4,097 132 10,800 3,877 83 704 7,501 148 14 7,501 148 16 1,500 <<1
+socfb-UCSB37 14,917 356,918 125,297 16,708 287 10,800 219,965 3,286 256 254,965 4,013 16 255,924 3,994 22 13,655 3
+socfb-Duke14 9,885 374,731 131,706 11,565 198 10,800 125,940 1,887 39 151,169 2,415 15 151,365 2,390 18 9,003 4
+p-hat1500-2 1,500 411,372 143,918 3,542 53 10,800 2,611 43 2,150 4,253 68 <<1 4,253 68 <<1 1,500 <<1
+socfb-Stanford3 11,586 136,823 431,486 7,250 122 10,800 215,381 3,932 265 251,350 4,441 12 251,932 4,436 14 9,380 2
+socfb-UConn 17,206 448,038 156,829 16,996 279 10,800 231,718 3,471 322 274,637 4,329 12 275,320 4,299 18 16,025 5
+p-hat1500-1 1,500 411,744 427,583 0 0 10,800 1,184 18 5,499 1,287 19 <<1 1,287 19 1 1,500 <<1
+C2000.5 2,000 489,894 509,270 0 0 10,800 1,581 24 865 1,849 26 2 1,849 26 6 2,000 1
+keller6 3,361 759,816 266,766 0 0 10,800 3,545 55 1,561 5,236 73 10 5,236 73 16 3,361 <<1
+C4000.5 4,000 1,957,884 2,039,848 0 0 10,800 1,611 25 1,659 1,948 28 20 1,952 28 22 4,000 9
+# of best / # of instances 24/47 1/47 18/47 41/47
+average running time(s) 5,402.1 2,087.4 7.9 5.9 1.4
+Next, we compared the RLS with LA-B&B, ALS and CPLEX on SET-2 and SET-3.
+Each of the two heuristic ALS and the RLS were run 10 trials and reported its best known objective value.
+Each trial was given a cutoff time of 30 seconds.
+In total, each heuristic algorithm ran 300 seconds per instance in SET-2.
+The settings are the same as in Zheng et al.
+(2024).
+The cutoff time of exact algorithms was directly set to 3 hours (10,800 seconds) per instance in SET-2.
+However, for SET-3, the heuristic trial time was increased to 200 seconds.
+For large instances like soc-pokec and LiveJ, the cutoff time of ALS and the RLS were given 1,000 seconds per trial.
+For the larger uk2002, the cutoff time of LA-B&B and CPLEX was 100,000 seconds, and the cutoff time of ALS and the RLS is 10,000 seconds per trial.
+In particular, for the three instances soc-pokec, LiveJ and uk2002, the RLS excluded $R6$ and R12R121414, as these reductions exhibited considerably slower performance compared to the other rules in the experiments.
+For LA-B&B and CPLEX, if the algorithm did not finish within the total running time, the best solution found so far was reported.
+[…]
 Table 4: Comparative results of RLS on SET-3 instances
-Instance $|V|$ $|E_{p}|$ $|E_{r}|$ LA-B&B ALS RLS
+Instance $|V|$ $|E_{p}|$ $|E_{r}|$ CPLEX LA-B&B ALS RLS
 […]
-foodweb-wet 128 1,509 566 2,464 54.84 2,464 $<$1 2,464 $<$1 125 $<$1
-foodweb-dry 128 1,571 535 2,226 27.87 2,226 $<$1 2,226 $<$1 125 $<$1
-USAir97 332 1,589 537 9,507 2000 10,308 $<$1 10,308 $<$1 0 $<$1
-powergrid 4,941 1,727 4,867 190,537 2000 193,095 59 193,339 $<$1 0 $<$1
-CondMat 23,133 46,769 46,670 95,130 2000 705,901 14 711,334 45 1,648 $<$1
-Email 265,009 91,093 273,388 OOM OOM 13,054,427 1 0 1
-Epinion 75,879 202,653 203,087 7,400 2000 3,038,533 18 3,052,059 1 253 1
-Dblp 317,080 787,268 262,598 OOM OOM 9,734,653 64 11,851 11
-cnr-2000 325,557 685,686 2,053,283 OOM OOM 13,368,670 482 59,052 399
-WikiTalk 2,394,385 3,494,873 1,164,692 OOM OOM 118,708,326 333 8 333
-BerkStan 685,230 1,662,833 4,986,637 OOM OOM 26,031,704 1112 114,191 1031
-As-Skitter 1,696,415 5,550,661 5,544,637 OOM OOM 66,582,254 723 31,173 636
-soc-pokec 1,632,803 11,152,945 11,149,019 OOM OOM 43,999,402 934 881,319 134
-LiveJ 4,846,609 10,710,579 32,140,658 OOM OOM 168,431,056 904 401,307 469
-uk2002 18,483,186 65,451,024 196,336,234 OOM OOM 687,089,901 7678 4,806,660 6243
-# of best / # of instances 3/15 3/15 15/15
-average running time(s) 1347.1 15.7 818.7 617.5
-[…]
-The comparative results on SET-2 and SET-3 are presented in Tables 3 and 4.
-The first four columns list the instance information.
-The remaining columns show the best-known objective value (BKV), the running time of the exact algorithm (time), the mean running time that the heuristic algorithm first hits the BKV (htime), the kernel graph size ($|V_{\text{ker}}|$) and the time of generating a ${}_{\text{ker}}$).
-In the last two rows, we show the statistical results.
-Note that the htime of RLS has included the pre-processing time for building graph kernels.
-Table 3 shows that RLS outperforms other methods on all SET-2 instances in terms of both best-known objective value and runtime.
-It is known that RLS proves optimality if all vertices are reduced (i.e.
-$|V_{kernel}|=0$), hence 11 instances have been solved to optimality by RLS.
+foodweb-wet 128 1,509 566 2,464 44 <<1 2,464 44 8 2,464 44 <<1 2,464 44 <<1 125 <<1
+foodweb-dry 128 1,571 535 2,226 42 <<1 2,226 42 8 2,226 42 <<1 2,226 42 <<1 125 <<1
+USAir97 332 1,589 537 10,308 194 <<1 9,507 178 407 10,308 194 <<1 10,308 194 <<1 0 <<1
+powergrid 4,941 1,727 4,867 193,339 3,710 <<1 190,537 3,684 108 193,095 3,754 59 193,339 3,537 <<1 0 <<1
+CondMat 23,133 46,769 46,670 711,517 12,821 12 685,442 12,061 5,749 705,901 1,306 14 711,334 12,870 45 1,648 <<1
+Email 265,009 91,093 273,388 13,054,427 257,435 118 OOM OOM 13,054,427 257,394 1 0 1
+Epinion 75,879 202,653 203,087 3,052,059 57,976 60 OOM 3,038,533 58,267 18 3,052,059 58,008 1 253 1
+Dblp 317,080 787,268 262,598 9,737,911 171,421 790 OOM OOM 9,734,653 171,403 64 11,851 11
+cnr-2000 325,557 685,686 2,053,283 13,367,460 260,331 10,800 OOM OOM 13,368,670 260,716 482 59,052 399
+WikiTalk 2,394,385 3,494,873 1,164,692 118,708,326 2,342,835 3,566 OOM OOM 118,708,326 2,342,861 333 8 333
+BerkStan 685,230 1,662,833 4,986,637 26,036,210 501,197 10,800 OOM OOM 26,031,704 501,410 1,112 114,191 1,031
+As-Skitter 1,696,415 5,550,661 5,544,637 66,603,360 1,264,844 10,800 OOM OOM 66,582,254 1,265,266 723 31,173 636
+soc-pokec 1,632,803 11,152,945 11,149,019 4,146,172 84,217 10,800 OOM OOM 43,999,402 833,989 934 881,319 134
+LiveJ 4,846,609 10,710,579 32,140,658 0 0 10,800 OOM OOM 168,431,056 3,258,831 904 401,307 469
+uk2002 18,483,186 65,451,024 196,336,234 0 0 100,000 OOM OOM 687,089,901 13,716,043 7,678 4,806,660 6,243
+# of best / # of instances 11/15 3/15 3/15 11/15
+average running time(s) 10,570.1 1,256 15.7 818.7 617.5
+Table 3 shows that the RLS outperforms other methods on all SET-2 instances in terms of both best-known objective value and runtime.
+It is known that the RLS proves optimality if all vertices are reduced (i.e., $|V_{kernel}|=0$), hence 11 out of 47 instances have been solved to optimality by the RLS.
 In contrast, LA-B&B can only solve 1 out of these 47 instances.
-Moreover, RLS and ALS hit the BKV in just a few seconds.
-Table 5: Comparative results of algorithms with or without reduction rules
-Instance $|V|$ $|V_{ker}|$ LA-B&B RLS
-without reduction with reduction without reduction with reduction
-BKV time(s) BKV time(s) BKV htime(s) BKV htime(s)
-bio-yeast 1,458 0 66,545 300 68,574 0 68,574 1 68,574 $<$1
-soc-wiki-Vote 889 0 35,973 300 37,358 0 37,358 1 37,358 $<$1
-vc-exact_131 2,980 0 107,812 300 109,641 0 109,567 9 109,641 $<$1
-web-edu 3,031 98 109,856 300 113,781 15 114,080 12 114,140 $<$1
-MANN_a81 3,321 0 114,616 300 117,327 0 117,275 14 117,327 $<$1
-tech-routers-rf 2,113 0 93,690 300 97,799 0 97,799 1 97,799 $<$1
-vc-exact_039 6,795 0 254,941 300 259,542 0 259,475 13 259,542 $<$1
-ca-GrQc 4,158 282 167,846 300 174,654 15 174,768 7 174,782 11
-vc-exact_038 786 697 10,371 300 12,014 12 12,015 13 12,015 4
-vc-exact_078 11,349 0 102,312 300 438,671 1 438,306 16 438,671 $<$1
-vc-exact_087 13,590 0 71,804 300 518,500 0 518,069 15 518,500 $<$1
-vc-exact_107 6,402 0 88,602 300 407,878 0 406,379 14 407,878 $<$1
-vc-exact_151 15,783 0 54,114 300 530,138 0 529,003 18 530,138 $<$1
-vc-exact_167 15,783 0 55,190 300 465,178 0 463,446 12 465,178 $<$1
-bio-dmela 7,393 44 280,897 300 302,449 10 302,793 18 302,992 $<$1
-vc-exact_011 9,877 172 190,729 300 301789 7 303,352 9 303,793 $<$1
-web-spam 4,767 487 180,446 300 190,344 14 190,460 18 190,459 13
-vc-exact_026 6,140 300 177,898 300 201,316 17 201,458 13 201,503 $<$1
-vc-exact_001 6,160 134 190,729 300 215,668 13 215,934 13 215,941 $<$1
-vc-exact_024 7,620 1,255 226,540 300 235,440 14 235,911 14 236,018 12
-C1000.9 1,000 1,000 6,102 300 6,102 300 7,235 16 7,237 10
-tech-WHOIS 7,476 338 309,903 300 325,327 15 325,480 16 325,581 13
-p-hat700-3 700 700 2,937 300 2,937 300 4,473 1 4,473 $<$1
-vc-exact_008 7,537 574 71,804 300 246,771 14 247,036 16 247,091 $<$1
-keller5 776 776 2,788 300 2,788 300 3,877 2 3,877 2
-vc-exact_194 1,150 1,150 4,273 300 4,273 300 5419 2 5,419 $<$1
-hamming10-4 1,024 1,024 2,871 300 2,871 300 3660 1 3660 $<$1
-brock800_2 800 800 1,996 300 1,996 300 2473 1 2473 $<$1
-brock800_4 800 800 1,968 300 1,968 300 2501 1 2501 $<$1
-ca-HepPh 11,204 2,165 153,907 300 372,079 6 372,079 14 372,079 17
-p-hat700-2 700 700 2,127 300 2,127 300 2,971 1 2,971 $<$1
-vc-exact_196 1,534 1,534 4,273 300 4,273 300 5,321 2 5,321 $<$1
-p-hat700-1 700 700 844 124 844 124 844 1 844 $<$1
-ca-AstroPh 17,903 4,985 44,296 300 602,320 2 603,732 15 603,945 14
-C2000.9 2,000 2,000 5,477 300 5,477 300 6,396 3 6,396 3
-DSJC1000.5 1,000 1,000 1,492 300 1,492 300 1,861 1 1,861 $<$1
-socfb-MIT 6,402 5,494 110,677 300 129,807 17 129,831 16 129,872 16
-p-hat1500-3 1,500 1,500 3,856 300 3,856 300 7,500 15 7,501 16
-socfb-UCSB37 14,917 13,655 219,965 300 254,965 16 255,887 21 255,898 22
-socfb-Duke14 9,885 9,003 125,940 300 151,169 15 151,321 14 151,326 18
-p-hat1500-2 1,500 1,500 2444 300 2,444 300 4,253 1 4,253 $<$1
-socfb-Stanford3 11,586 9,380 215,381 300 251,350 12 251,788 19 251,833 14
-socfb-UConn 17,206 16,025 170,548 300 274,637 12 275,006 16 275,034 20
-p-hat1500-1 1500 1,500 1,121 300 1,121 300 1,287 2 1,287 1
-C2000.5 2,000 2,000 1,570 300 1,570 300 1,849 5 1,849 6
-keller6 3,361 3,361 3,545 300 3,545 300 5,229 19 5,236 16
-C4000.5 4,000 4,000 1,601 300 1,601 300 1,951 15 1,951 22
-foodweb-wet 128 125 2,464 55 2,464 27 2,464 $<$1 2,464 $<$1
-foodweb-dry 128 125 2,226 28 2,226 25 2,226 $<$1 2,226 $<$1
-USAir97 332 0 9,507 2000 10,308 0 10,308 $<$1 10,308 $<$1
-powergrid 4,941 0 190,537 2000 193,339 0 193,194 52 193,339 $<$1
-CondMat 23,133 1,648 95,130 2000 708,907 2000 709,927 54 711,334 45
-Email 265,009 0 OOM 13,054,427 1 13,054,390 69 13,054,427 1
-Epinion 75,879 253 7,400 2000 3,051,256 2000 3,050,415 67 3,052,059 1
-Dblp 317,080 11,851 OOM 9,715,860 2000 9,577,609 46 9,734,653 64
-cnr-2000 325,557 59,052 OOM 11,347,241 2000 13,361,167 84 13,368,670 482
-WikiTalk 2,394,385 8 OOM 118,708,326 $<$1 118,672,055 56 118,708,326 333
-BerkStan 685,230 114,191 OOM 22,591,394 2000 25,813,846 2000 26,031,704 1112
-As-Skitter 1,696,415 31,173 OOM 65,907,243 2000 59,886,304 2000 66,582,254 723
-soc-pokec 1,632,803 881,319 OOM OOM 36,540,838 350 43,999,402 934
-LiveJ 4,846,609 401,307 OOM OOM 148,758,359 318 168,431,056 904
-uk2002 18,483,186 4,806,660 OOM OOM 633,365,020 1806 687,089,901 7678
-# of best / # of instances 3/62 24/62 24/62 62/62
-average running time(s) 415.2 296.7 119.4 202.6
+Similarly, CPLEX fails to solve instances with 40,000 variables (e.g., vc-exact_024 and C1000.9).
+Moreover, CPLEX even fails to report a feasible solution for 4 large instances.
+In contrast, the RLS and ALS hit the BKV with an average hitting time of no more than 8 seconds.
 Table 4 compares algorithms on SET-3 instances.
@@
 Both ALS and LA-B&B quickly become infeasible as instance size increases.
-In contrast, RLS computes feasible solutions even for large instances such as uk2002 .
-For instances where the benchmarks are feasible, RLS consistently achieves the best net benefit.
-For the infeasible instance Email, RLS even finds the optimal solution.
+CPLEX again fails to report a feasible solution for large instances.
+In contrast, the RLS computes feasible solutions even for large instances such as uk2002 .
+For instances where the benchmarks are feasible, the RLS consistently achieves the best net benefit.
+For the infeasible instance Email, the RLS even finds the optimal solution.
 6.4 Evaluation of Reduction Rules
-To evaluate the effectiveness of the reduction rules, we conduct the following two experiments.
+To evaluate the effectiveness of the reduction rules, we conducted the following two experiments.
+Table 5: Comparative results of algorithms with or without reduction rules
+Instance $|V|$ $|V_{ker}|$ CPLEX LA-B&B RLS
+without reduction with reduction without reduction with reduction without reduction with reduction
+BKV htime(s) BKV htime(s) BKV htime(s) BKV htime(s) BKV htime(s) BKV htime(s)
+bio-yeast 1,458 0 68,574 <<1 68,574 <<1 66,545 2 68,574 <<1 68,574 1 68,574 <<1
+soc-wiki-Vote 889 0 37,358 <1<1 37,358 <<1 35,973 <<1 37,358 <<1 37,358 1 37,358 <<1
+vc-exact_131 2,980 0 109,641 <1<1 109,641 <<1 107,812 14 109,641 <<1 109,567 9 109,641 <<1
+web-edu 3,031 98 114,140 <1<1 114,140 <1<1 109,856 10 114,140 21 114,080 12 114,140 <<1
+MANN_a81 3,321 0 117,327 <1<1 117,327 <<1 114,616 16 117,327 <<1 117,275 14 117,327 <<1
+tech-routers-rf 2,113 0 97,799 <1<1 97,799 <<1 93,690 54 97,799 <<1 97,799 1 97,799 <<1
+vc-exact_039 6,795 0 259,542 <1<1 259,542 <<1 254,941 253 259,542 <<1 259,475 13 259,542 <<1
+ca-GrQc 4,158 282 174,782 9 174,782 6 167,846 45 174,525 2,998 174,768 7 174,782 11
+vc-exact_038 786 697 12,015 592 12,015 585 10,406 1,994 10,782 28 12,015 13 12,015 4
+vc-exact_078 11,349 0 438,671 <1<1 438,671 1 431,371 1,074 438,671 1 438,306 16 438,671 <<1
+vc-exact_087 13,590 0 518,500 2 518,500 <<1 509,734 2,319 518,500 <<1 518,069 15 518,500 <<1
+vc-exact_107 6,402 0 407,878 <1<1 407,878 <<1 396,456 10,646 407,878 <<1 406,379 14 407,878 <<1
+vc-exact_151 15,783 0 530,138 2 530,138 <<1 516,445 2,402 530,138 <<1 529,003 18 530,138 <<1
+vc-exact_167 15,783 0 465,178 1 465,178 <<1 449,579 10,469 465,178 <<1 463,446 12 465,178 <<1
+bio-dmela 7,393 44 302,992 <1<1 302,992 <1<1 280,897 170 302,992 <1<1 302,793 18 302,992 <<1
+vc-exact_011 9,877 172 303,793 1 303,793 <1<1 286,137 4,818 303,589 1271 303,352 9 303,793 <<1
+web-spam 4,767 487 190,507 5 190,507 <1<1 180,446 53 189,897 <1<1 190,460 18 190,460 13
+vc-exact_026 6,140 300 201,509 1 201,509 <1<1 177,902 668 200,868 <1<1 201,458 13 201,503 <<1
+vc-exact_001 6,160 134 215,941 2 215,941 <1<1 194,984 9,575 215,941 2 215,934 13 215,941 <<1
+vc-exact_024 7,620 1,255 236,109 10,800 236,109 10,800 226,540 244 234,933 11 235,911 14 236,018 12
+C1000.9 1,000 1,000 5,263 10,800 5,263 10,800 6,104 430 6,104 430 7,235 16 7,237 10
+tech-WHOIS 7,476 338 325,584 5 325,584 <1<1 309,903 218 325,119 255 325,480 16 325,581 13
+p-hat700-3 700 700 4,025 10,800 4,025 10,800 3,030 7,436 3,030 7,436 4,473 1 4,473 <<1
+vc-exact_008 7,537 574 247,091 3 247,091 <1<1 210,126 9,093 245,289 6,627 247,036 16 247,091 <<1
+keller5 776 776 2,373 10,800 2,373 10,800 2,860 6,327 2,860 6,327 3,877 2 3,877 2
+vc-exact_194 1,150 1,150 3,734 10,800 3,734 10,800 4,386 1,324 4,386 1,324 5,419 2 5,419 <<1
+hamming10-4 1,024 1,024 3,256 10,800 3,256 10,800 3,073 2,069 3,073 2,069 3,660 1 3,660 <<1
+brock800_2 800 800 1,177 10,800 1,177 10,800 1,996 10 1,996 10 2,473 1 2,473 <<1
+brock800_4 800 800 1,583 10,800 1,583 10,800 1,979 7,215 1,979 7,215 2,501 1 2,501 <<1
+ca-HepPh 11,204 2,165 372,096 10,800 372,229 10,800 348,738 422 368,637 2 372,079 14 372,079 17
+p-hat700-2 700 700 2,870 10,800 2,870 10,800 2,345 2,078 2,345 2,078 2,971 1 2,971 <<1
+vc-exact_196 1,534 1,534 3,916 10,800 3,916 10,800 4,273 66 4,273 66 5,321 2 5,321 <<1
+p-hat700-1 700 700 730 10,800 730 10,800 844 5 844 5 844 1 844 <<1
+ca-AstroPh 17,903 4,985 604,288 10,800 604,342 10,800 558,736 2,544 593,899 37 603,732 15 603,945 14
+C2000.9 2,000 2,000 3,958 10,800 3,958 10,800 5,477 11 5,477 11 6,396 3 6,396 3
+DSJC1000.5 1,000 1,000 925 10,800 925 10,800 1,515 694 1,515 694 1,861 1 1,861 <<1
+socfb-MIT 6,402 5,494 7,098 10,800 118,114 10,800 110,677 20 113,623 10 129,831 16 129,872 16
+p-hat1500-3 1,500 1,500 4,097 10,800 4,097 10,800 3,877 704 3,877 704 7,500 15 7,501 16
+socfb-UCSB37 14,917 13,655 16,708 10,800 57,084 10,800 219,965 256 223,780 150 255,887 21 255,898 22
+socfb-Duke14 9,885 9,003 11,565 10,800 127,145 10,800 125,940 39 127,269 2,231 151,321 14 151,326 18
+p-hat1500-2 1,500 1,500 3,542 10,800 3,542 10,800 2,611 2,150 2,611 2,150 4,253 1 4,253 <<1
+socfb-Stanford3 11,586 9,380 7,250 10,800 127,708 10,800 215,381 265 223,398 7,756 251,788 19 251,833 14
+socfb-UConn 17,206 16,025 16,996 10,800 51,328 10,800 231,718 322 235,104 240 275,006 16 275,034 20
+p-hat1500-1 1,500 1,500 0 10,800 0 10,800 1,184 5,499 1,184 5,499 1,287 2 1,287 1
+C2000.5 2,000 2,000 0 10,800 0 10,800 1,581 865 1,581 865 1,849 5 1,849 6
+keller6 3,361 3,361 0 10,800 0 10,800 3,545 1,561 3,545 1,561 5,229 19 5,236 16
+C4000.5 4,000 4,000 0 10,800 0 10,800 1,611 1,659 1,611 1,659 1,951 15 1,951 22
+foodweb-wet 128 125 2,464 <1<1 2,464 <1<1 2,464 8 2,464 6 2,464 <<1 2,464 <<1
+foodweb-dry 128 125 2,226 <1<1 2,226 <1<1 2,226 8 2,226 8 2,226 <<1 2,226 <<1
+USAir97 332 0 10,308 <1<1 10,308 <1<1 9,507 407 10,308 <1<1 10,308 <<1 10,308 <<1
+powergrid 4,941 0 193,339 <1<1 193,339 <1<1 190,537 108 193,339 <1<1 193,194 52 193,339 <<1
+CondMat 23,133 1,648 711,517 12 711,517 <1<1 685,442 5,749 708,907 84 709,927 54 711,334 45
+Email 265,009 0 13,054,427 118 13,054,427 1 OOM 13,054,427 1 13,054,390 69 13,054,427 1
+Epinion 75,879 253 3,052,059 60 3,052,059 1 OOM 3,051,256 196 3,050,415 67 3,052,059 1
+Dblp 317,080 11,851 9,737,911 790 9,737,911 39 OOM 9,715,860 307 9,577,609 46 9,734,653 64
+cnr-2000 325,557 59,052 13,367,460 10,800 13,641,278 10,800 OOM 11,347,241 10,796 13,361,167 84 13,368,670 482
+WikiTalk 2,394,385 8 118,708,326 3,566 118,708,326 333 OOM 118,708,326 333 118,672,055 56 118,708,326 333
+BerkStan 685,230 114,191 26,036,210 10,800 26,055,767 10,800 OOM 22,591,394 1,031 25,813,846 200 26,031,704 1,112
+As-Skitter 1,696,415 31,173 66,603,360 10,800 66,830,857 10,800 OOM 65,907,243 10,512 59,886,304 200 66,582,254 723
+soc-pokec 1,632,803 881,319 4,146,172 10,800 27,037,851 10,800 OOM OOM 36,540,838 350 43,999,402 934
+LiveJ 4,846,609 401,307 0 10,800 163,061,840 10,800 OOM OOM 148,758,359 318 168,431,056 904
+uk2002 18,483,186 4,806,660 0 100,000 567,666,923 100,000 OOM OOM 633,365,020 1,806 687,089,901 7,678
+# of best / # of instances 31/62 36/62 3/62 21/62 22/62 51/62
+average running time(s) 7,096.5 7,028.8 2,007.5 1,441.1 61.3 202.6
 Comparative results with vs.
@@
 This experiment evaluates algorithms with and without data reduction.
-For LA-B&B with reduction, we use kernel graphs as input.
-For RLS without reduction, all reductions are skipped, and random initialization is applied.
+For LA-B&B and CPLEX with the reductions, we used kernel graphs as input.
+For the RLS without reduction, all reductions were skipped, and random initialization was applied.
 ALS cannot handle kernel graphs due to unsupported negative vertex profits and edge penalties.
 The comparative results are shown in Table 5.
-In general, the reduction rules can help both algorithms to find better solutions.
-Specifically, with the reduction technique, LA-B&B is able to solve some instances which it cannot solve without reduction, e.g.
-WikiTalk (1.6 million vertices, 11.1 million edges).
-And for some large instances that LA-B&B fails due to out-of-memory, it is able to report a feasible solution or even solve them.
-As for RLS, with these reduction rules, it consistently finds better results across all instances in a shorter time than without them.
-Figure 5: The variation of kernel sizes when ablating different reductions.
-A cell value indicates the difference in kernel size between without using reductions and with using reductions.
-A higher cell value indicates a larger effect.
-Figure 6: The variation of pre-processing time for getting a kernel.
-A cell value indicates the difference in processing time(s) between without using the corresponding reductions and with using the corresponding reductions.
-A higher cell value indicates a larger effect.
+In general, the reduction rules can help all algorithms to find better solutions.
+Specifically, with the reduction techniques, LA-B&B is able to solve some instances which it cannot solve without reduction, e.g., WikiTalk (1.6 million vertices, 11.1 million edges).
+For some large instances where original LA-B&B run out-of-memory, the new LA-B&B with the reductions can report a feasible solution or solve them to optimality within the cutoff time.
+With the reduction techniques, CPLEX can solve some instances with less time.
+For two large instances LiveJ and uk2002 where CPLEX fails, CPLEX with the reductions is able to report feasible solutions.
+As for the RLS, with these reduction rules, it consistently finds better results across all instances in a shorter time than without them.
+Figure 5: Variation of kernel sizes when ablating different reductions.
+Each cell shows the difference in kernel size between not using and using the corresponding reductions; higher values indicate larger effects.
+Figure 6: Variation of pre-processing time for getting a kernel.
+Each cell shows the difference in time (seconds) between not using and using the corresponding reductions; higher values indicate larger effects.
 Ablation study of reduction rules.
@@
 Figure 5 shows that the impacts on the kernel size vary among these reduction rules.
-When comparing individual reductions, $R10$ and $R12$ are more powerful than others in decreasing the kernel size.
-When comparing a group of reductions, the degree-based rules ($R7$$R11$) are the most powerful in decreasing the kernel size.
-The neighborhood-based rules ($R3$$R6$) are generally more effective than vertex-pair reduction rules ($R12$$R14$) except for Dblp and WikiTalk.
+When comparing individual reductions, $R10$ and $R12$ are more powerful than the other reduction rules in decreasing the kernel size.
+When comparing a group of reductions, the degree-based rules (R7R7R11R11) are the most powerful in decreasing the kernel size.
+The neighborhood-based rules (R3R3R6R6) are generally more effective than the vertex-pair reduction rules (R12R12R14R14) except for Dblp and WikiTalk.
 Additionally, although $R1$ and $R2$ do not involve vertex removals, they are also powerful in decreasing the kernel size.
-In particular, incorporating $R2$ into RLS can further reduce more than 3000 vertices for the BerkStan instance.
-It is intuitive that ablating a reduction rule can decrease the overall pre-processing time, but may at the cost of resulting in a larger kernel.
-However, Figure 6 shows that ablating reduction rules $R3$$R6$ or $R7$$R11$ can even increase the pre-processing time.
-This is due to the fact that other reductions cost more time when these reduction rules are absent.
-We also observe that for the larger four instances As-Skitter, Berkstan, WikiTalk and cnr-2000, the inclusion of reductions shows a significant improvement over the pre-processing time.
+In particular, incorporating $R2$ into the RLS can further reduce more than 3,000 vertices for the BerkStan instance.
+Intuitively, ablating a reduction rule can decrease the overall pre-processing time, but may at the cost of resulting in a larger kernel.
+However, Figure 6 shows that ablating reduction rules R3R3R6R6 or R7R7R11R11 can even increase the pre-processing time.
+This occurs because the remaining reductions become more time-consuming in the absence of these rules.
+We also observe that for the four larger instances As-Skitter, Berkstan, WikiTalk and cnr-2000, the inclusion of reductions shows a significant improvement over the pre-processing time.
 7 Conclusion and Future Works
@@
 Experiments on benchmark instances showed that the RLS algorithm outperformed state-of-the-art methods, especially on large graphs with tens of thousands of vertices.
-Future work includes exploring advanced reduction techniques for improved efficiency and adapting our algorithms to more real-world scenarios like dynamic graphs.
+Future work will focus on developing more advanced reduction techniques to further improve efficiency and extending our algorithms to real-world scenarios such as dynamic graphs with changing weights.
 Appendix
@@
 Consider whether $u$ and $v$ are included in a generalized independent set $S$.
-There are two cases: either $u\notin S$ or $v\notin S$, and $\{u,v\}\subseteq S$.
-When either $u\notin S$ or $v\notin S$, the edge $(u,v)$ does not appear in $S$, and thus removing $(u,v)$ has no effect on the net benefit $nb(S)$.
-If both $u\in S$ and $v\in S$, the edge $(u,v)$ has a zero penalty, which also has no effect on the net benefit $nb(S)$.
+There are four cases: $u\in S$ and $v\in S$, $u\in S$ and $v\notin S$, $u\notin S$ and $v\in S$, and $u\notin S$ and $v\notin S$.
+When either $u\notin S$ or $v\notin S$, the edge $e=\{u,v\}$ does not exist in $G[S]$, and thus removing $e$ has no effect on the net benefit $nb(S)$.
+If both $u\in S$ and $v\in S$, the edge $\{u,v\}$ has a zero penalty, which has no effect on the net benefit $nb(S)$ either.
 ∎
 Proof of R2.
-Assume, for the sake of contradiction, that there exists $I\in MGIS(G)$ that $u\in I$ and $v\in I$.
-We have $nb(I)-nb(I\setminus\{u\})=w(u)-\sum_{x\in N_{r}(u)\cap S}p(u,x)\leq w(u)-p(u,v)+\sum_{x\in N_{r}(u)}max(0,-p(u,x))=\tilde{w}(u)-p(u,v)<0$.
-In conclusion, $I\setminus\{u\}$ has a larger net benefit, which contradicts the optimality of $I$.
-Therefore, $u$ and $v$ cannot be simultaneously included in any $S\in MGIS(G)$, which is equivalent to the case where $p(u,v)=M$.
+Assume, for the sake of contradiction, that there exists $I\in MGIS(G)$ such that $u\in I$ and $v\in I$.
+Assume, without loss of generality, that $\tilde{w}(u)\leq\tilde{w}(v)$.
+We have $nb(I)-nb(I\setminus\{u\})=w(u)-\sum_{x\in N_{r}(u)\cap{I}}p(\{u,x\})\leq w(u)-p(\{u,v\})+\sum_{x\in N_{r}(u)}max(0,-p(\{u,x\}))=\tilde{w}(u)-p(\{u,v\})<0$.
+In conclusion, $I\setminus\{u\}$ has a larger net benefit without containing permanent edges, which contradicts the optimality of $I$.
+Therefore, $u$ and $v$ cannot be simultaneously included in any $S\in MGIS(G)$, which is equivalent to the case where $\{u,v\}\in E_{p}$.
 ∎
-The proof of R3 directly follows from the Reduction 2 of Lamm et al. (2019).
+Proof of R3.
+We prove that $u$ is in at least one set in $MGIS(G)$.
+Assume, for the sake of contradiction, that $u$ is not included in any set in $MGIS(G)$.
+Let $I$ be an arbitrary set in $MGIS(G)$.
+It is clear that at least one vertex in $N(u)$ must be included in $I$.
+Otherwise, $I\cup\{u\}$ implies with a larger net benefit without including any permanent edge, a contradiction, since $w(u)\geq w^{+}(N(u))$, $nb(I\cup\{u\}\setminus N(u))\geq nb(I)$.
+So $u$ is in at least one set in $MGIS(G)$.
+It holds that $\alpha(G)=\alpha(G^{\prime})+w(u)$.
+∎
 Proof of R4.
-We will prove that $u$ is in some set in $MGIS(G)$.
+We prove that $u$ is in at least one set in $MGIS(G)$.
 Assume, for the sake of contradiction, that $u$ is not included in any set in $MGIS(G)$.
@@
 Now, we remove $S$ from $I$ and add $u$ to $I$ to obtain $I^{\prime}$.
-Consider the difference between two net benefits $nb(I^{\prime})-nb(I)\geq w(u)-\sum_{v\in S}\tilde{w}(v)-\sum_{v\in S^{\prime}}p(u,v)\geq 0$.
+Consider the difference between two net benefits $nb(I^{\prime})-nb(I)\geq w(u)-\sum_{v\in S}\tilde{w}(v)-\sum_{v\in S^{\prime}}p(\{u,v\})\geq 0$.
+Moreover, $G[I^{\prime}]$ does not contain permanent edges.
 Therefore, $I^{\prime}\in MGIS(G)$, which contradicts the assumption.
 So $u$ is in at least one set in $MGIS(G)$.
-And it holds that $\alpha(G)=\alpha(G^{\prime})+w(u)$.
+It holds that $\alpha(G)=\alpha(G^{\prime})+w(u)$.
 ∎
@@
 Given an arbitrary $I\in MGIS(G)$, there are two cases.
-When none of the neighbors of $u$ is included in $I$, $I$ does not include $u$ because $w(u)<0$.
-When any subset of neighbors of $u$ is included in $I$, $I$ also does not include $u$ because $w(u)-\sum_{v\in S}p(u,v)<0$ for any subset $S\subseteq N(u)$.
+When none of the neighbors of $u$ is included in $I$, II does not include $u$ because $w(u)<0$.
+When any subset of neighbors of $u$ is included in $I$, II does not include $u$ either because $w(u)-\sum_{v\in S}p(\{u,v\})<0$ for any subset $S\subseteq N(u)$.
 Since $I$ is an arbitrary set in $MGIS(G)$, $u$ will not appear in any $I\in MGIS(G)$ and we can directly remove $u$ from $G$.
@@
 Proof of R6.
-Observe that there exists a vertex set $I\in MGIS(G)$ that $I\cap N_{p}[u]\neq\varnothing$, because if $I\cap N_{p}[u]=\varnothing$ then $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$ implies $I\cup\{u\}\in MGIS(G)$.
+Observe that there exists a vertex set $I\in MGIS(G)$ such that $I\cap N_{p}[u]\neq\varnothing$, because if $I\cap N_{p}[u]=\varnothing$ then $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$ implies $I\cup\{u\}\in MGIS(G)$.
 Whenever a vertex $x$ in $N_{p}(u)$ is included in $I$, other vertices in the clique $N_{p}[u]\setminus\{x\}$ must not be included in $I$.
-Then $I\cup\{u\}\setminus\{x\}\in MGIS(G)$ because $nb(I\cup\{u\}\setminus\{x\})\geq nb(I)+w(u)-w(x)+\sum_{v\in N_{r}(x)}p(v,x)-\sum_{v\in N_{r}(u)}max(0,p(u,v))\geq nb(I)+w(u)-\sum_{v\in N_{r}(u)}max(0,p(u,v))-\max(0,\tilde{w}(x))\geq nb(I)$.
-In conclusion, $u$ is in some set in $MGIS(G)$.
+Then $I\cup\{u\}\setminus\{x\}\in MGIS(G)$ because $G[I\cup\{u\}\setminus\{x\}]$ does not contain permanent edges, and $nb(I\cup\{u\}\setminus\{x\})\geq nb(I)+w(u)-w(x)+\sum_{v\in N_{r}(x)}p(\{v,x\})-\sum_{v\in N_{r}(u)}max(0,p(\{u,v\}))\geq nb(I)+w(u)-\sum_{v\in N_{r}(u)}max(0,p(\{u,v\}))-\max(0,\tilde{w}(x))\geq nb(I)$.
+In conclusion, $u$ is in at least one set in $MGIS(G)$.
 ∎
@@
 Given an arbitrary $I\in MGIS(G)$.
-When $w(u)\geq p(u,v)$ and $w(u)\geq 0$.
-If $u\notin I$ and $v\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})=nb(I)+w(u)-p(u,v)\geq nb(I)$.
-If $u\notin I$ and $v\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$.
-In summary, $u$ is included in some set in $MGIS(G)$.
-When $w(u)<0$ and $w(u)\geq p(u,v)$.
-If $v\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})=nb(I)+w(u)-p(u,v)\geq nb(I)$ for $u\notin I$.
+When $w(u)\geq p(\{u,v\})$ and $w(u)\geq 0$.
+If $u\notin I$ and $v\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})=nb(I)+w(u)-p(\{u,v\})\geq nb(I)$.
+If $u\notin I$ and $v\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $G[I\cup\{u\}]$ does not contain permanent edges, and $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$.
+In summary, $u$ is included in at least one set in $MGIS(G)$.
+When $w(u)<0$ and $w(u)\geq p(\{u,v\})$.
+If $v\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})=nb(I)+w(u)-p(\{u,v\})\geq nb(I)$ for $u\notin I$, and $G[I\cup\{u\}]$ does not contain permanent edges.
 If $v\notin I$, then $nb(I\cup\{u\})<nb(I\setminus\{u\})$.
-In other word, there exists a set $S\in MGIS(G)$ that $u\in S$ if and only if there exists a set $S^{\prime}\in MGIS(G^{\prime})$ that $v\in S^{\prime}$.
+In other words, there exists a set $S\in MGIS(G)$ such that $u\in S$ if and only if there exists a set $S^{\prime}\in MGIS(G^{\prime})$ such that $v\in S^{\prime}$.
 It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,v$ that could appear in a set in $MGIS(G)$.
-When $w(u)\geq 0$, $w(u)<p(u,v)$ and $w(u)\geq\tilde{w}(v)$.
-If $u\notin I,v\in I$, then $nb(I\cup\{u\})\setminus\{v\})-nb(I)=w(u)-w(v)-\sum_{x\in N_{r}(v)\cap I}max(0,-p(v,x))=w(u)-\tilde{w}(v)\geq 0$, thus $I\cup\{u\}\in MGIS(G)$.
+When $w(u)\geq 0$, $w(u)<p(\{u,v\})$ and $w(u)\geq\tilde{w}(v)$.
+If $u\notin I,v\in I$, then $nb(I\cup\{u\})\setminus\{v\})-nb(I)=w(u)-w(v)-\sum_{x\in N_{r}(v)\cap I}max(0,-p(\{v,x\}))=w(u)-\tilde{w}(v)\geq 0$, implying $I\cup\{u\}\setminus\{v\}\in MGIS(G)$.
 Otherwise, if $u\notin I,v\notin I$, then $nb(I\cup\{u\})-nb(I)=w(u)\geq 0$, thus $I\cup\{u\}\in MGIS(G)$.
-In summary, $u$ is included in some vertex set $I^{\prime}\in MGIS(G)$ and $v$ is not in $I^{\prime}$.
-When $w(u)\geq 0$, $w(u)<p(u,v)$ and $w(u)<\tilde{w}(v)$.
+In summary, $u$ is included in at least one vertex set $I^{\prime}\in MGIS(G)$ and $v$ is not in $I^{\prime}$.
+When $w(u)\geq 0$, $w(u)<p(\{u,v\})$ and $w(u)<\tilde{w}(v)$.
 $u$ and $v$ cannot be both included in $I$.
-Because if $u,v\in I$, then $nb(I\setminus\{u\})=nb(I)+p(u,v)-w(u)>nb(I)$ contradicts to the optimality of $I$.
+Indeed, if $u,v\in I$, then $nb(I\setminus\{u\})=nb(I)+p(\{u,v\})-w(u)>nb(I)$.
+Moreover, $G[I\setminus\{u\}]$ does not contain permanent edges.
+This contradicts the optimality of $I$.
 Furthermore, if $v\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})-nb(I\setminus\{u\})=w(u)\geq 0$.
@@
 It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,v$ that could appear in a set in $MGIS(G)$.
-Finally, when $w(u)<0$ and $w(u)<p(u,v)$. We have $nb(I\cup\{u\})<nb(I\setminus\{u\})$, therefore $u$ is not in any set in $MGIS(G)$ and we can directly remove $u$. ∎
+Finally, when $w(u)<0$ and $w(u)<p(\{u,v\})$. We have $nb(I\cup\{u\})<nb(I\setminus\{u\})$, therefore $u$ is not in any set in $MGIS(G)$ and we can directly remove $u$. ∎
 Proof of R8.
-Given an arbitrary $I\in MGIS(G)$. Because that $p(x,y)=M$, it holds that $x$ and $y$ cannot be both included in $I$.
-Case 1: $w(u)\geq\max(p(u,x),p(u,y))$.
+Given an arbitrary $I\in MGIS(G)$. Because that $\{x,y\}\in E_{p}$, it holds that $x$ and $y$ cannot be both included in $I$.
+Case 1: $w(u)\geq\max(p(\{u,x\}),p(\{u,y\}))$.
 When $w(u)\geq 0$.
 If both $x$ and $y$ are not included in $I$ and $u\notin I$, then $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$.
-If $u\notin I$ and either $x\in I$ or $y\in I$, $nb(I\cup\{u\})\geq nb(I)+w(u)-\max(p(u,x),p(u,y))\geq nb(I)$.
+If $u\notin I$ and either $x\in I$ or $y\in I$, $nb(I\cup\{u\})\geq nb(I)+w(u)-\max(p(\{u,x\}),p(\{u,y\}))\geq nb(I)$.
 In summary, there exists a vertex set in $MGIS(G)$ that includes $u$.
@@
 If both $x$ and $y$ are not included in $I$, then $u\notin I$ because $nb(I\cup\{u\})=nb(I\setminus\{u\})+w(u)<nb(I\setminus\{u\})$.
-If either $x\in I$ or $y\in I$ and $u\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})\geq nb(I)+w(u)-\max(p(u,x),p(u,y))\geq nb(I)$.
+If either $x\in I$ or $y\in I$ and $u\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})\geq nb(I)+w(u)-\max(p(\{u,x\}),p(\{u,y\}))\geq nb(I)$.
 We conclude that there exists a vertex set in $MGIS(G)$ that includes $u$ iff there exists a vertex set in $MGIS(G^{\prime})$ that includes either $x$ or $y$.
-In other word, we can defer the decision of $u$ after determining $x$ and $y$.
+In other words, we can defer the decision of $u$ after determining $x$ and $y$.
 It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
-Case 2: $p(u,y)\leq w(u)<p(u,x)$. In this case, if $y$ is in $I$, then $x\notin I$ and $I\cup\{u\}\in MGIS(G)$.
+Case 2: $p(\{u,y\})\leq w(u)<p(\{u,x\})$. In this case, if $y$ is in $I$, then $x\notin I$ and $I\cup\{u\}\in MGIS(G)$.
 When $w(u)\geq 0$ and $w(u)\geq\tilde{w}(x)$.
@@
 There are five cases: $x\in I$ and $u,y\notin I$, $y\in I$ and $u,x\notin I$, $u\in I$ and $x,y\notin I$, $u,y\in I$ and $x\notin I$, and $u,x\in I$ and $y\notin I$.
-But $u,x$ cannot appear in same $I$, because otherwise $nb(I\setminus\{u\})=nb(I)+p(u,x)-w(u)>0$ contradicts the optimality of $I$.
-And when $y\in I$ and $u,x\notin I$, then $I\cup\{u\}\in MGIS(G)$.
+But $u,x$ cannot appear in the same $I$, otherwise $nb(I\setminus\{u\})=nb(I)+p(\{u,x\})-w(u)>0$ without containing permanent edges, which contradicts the optimality of $I$.
+If $y\in I$ and $u,x\notin I$, then $I\cup\{u\}\in MGIS(G)$.
 In summary, there exists a vertex set $I^{\prime}\in MGIS(G)$ such that either $x\in I^{\prime}$, or both $u,y\in I^{\prime}$, or $u\in I^{\prime}$.
 Since $u$ only connects to $x$ and $y$ in $G$, we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that does not include $x$.
-In other word, we can defer $u$ after the decision of $x$.
+In other words, we can defer $u$ after the decision of $x$.
 It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
@@
 Since $u$ only connects to $x$ and $y$ in $G$, we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that includes $y$.
-In other word, we can defer $u$ after the decision of $y$.
+In other words, we can defer $u$ after the decision of $y$.
 It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
-Case 3: $w(u)<p(u,x)$ and $w(u)<p(u,y)$.
-When $w(u)\geq 0$, the three vertices form a triangle with permanent edges.
+Case 3: $w(u)<p(\{u,x\})$ and $w(u)<p(\{u,y\})$.
+When $w(u)\geq 0$, the three vertices form a triangle where every edge has penalty $M$.
 The result follows from theorem 3.3 of Gu et al.
 (2021).
-In particular, (1) if $w(u)\geq 0$ and $w(u)\geq\tilde{w}(x)$, then $u$ must be in some set in $MGIS(G)$.
-(2) If $w(u)\geq 0$ and $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$, then $y$ must not be in some set in $MGIS(G)$.
-And there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that does not include $x$.
+In particular, (1) if $w(u)\geq 0$ and $w(u)\geq\tilde{w}(x)$, then $u$ must be in at least one set in $MGIS(G)$.
+(2) If $w(u)\geq 0$ and $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$, then $y$ must not be in at least one set in $MGIS(G)$.
+There exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that does not include $x$.
 (3) If $w(u)\geq 0$ and $w(u)<\tilde{w}(y)$, then there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that does not include both $x$ and $y$.
@@
 Given an arbitrary $I\in MGIS(G)$. Notice that $x$ and $y$ can be both included in $I$.
-Case 1: $w(u)\geq p(u,x)$. In this case, if one of $x,y$ is in $I$, say $x$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})-nb(I)\geq w(u)-p(u,x)\geq 0$ when $u\notin I$.
-When $w(u)\geq 0$ and $w(u)\geq p(u,x)+p(u,y)$.
-If either $x\in I$ or $y\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})-nb(I)\geq w(u)-\max(p(u,x)+p(u,y),p(u,x))\geq 0$ for $u\notin I$.
+Case 1: $w(u)\geq p(\{u,x\})$. In this case, if one of $x,y$ is in $I$, say $x$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})-nb(I)\geq w(u)-p(\{u,x\})\geq 0$ when $u\notin I$.
+When $w(u)\geq 0$ and $w(u)\geq p(\{u,x\})+p(\{u,y\})$.
+If either $x\in I$ or $y\in I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})-nb(I)\geq w(u)-\max(p(\{u,x\})+p(\{u,y\}),p(\{u,x\}))\geq 0$ for $u\notin I$.
 If none of $x,y$ is in $I$, then $I\cup\{u\}\in MGIS(G)$ because $w(u)\geq 0$.
-Therefore, $u$ must belong to some set in $MGIS$.
-When $w(u)\geq 0$ and $w(u)<p(u,x)+p(u,y)$.
+Therefore, $u$ must belong to at least one set in $MGIS(G)$.
+When $w(u)\geq 0$ and $w(u)<p(\{u,x\})+p(\{u,y\})$.
 If both $x,y$ are in $I$, then $nb(I\cup\{u\})<nb(I)$.
 If either $x\notin I$ or $y\notin I$, then $nb(I\cup\{u\})\geq nb(I)$.
-In other word, including $u$ has negative profit only if both $x$ and $y$ are included in the $MGIS$.
-we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that either $x\notin S^{\prime}$ or $y\notin S^{\prime}$.
+In other words, including $u$ has negative profit only if both $x$ and $y$ are included in the same set in $MGIS(G)$.
+we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that either $x\notin S^{\prime}$ or $y\notin S^{\prime}$.
 It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
-When $w(u)<0$, it holds that $w(u)\geq p(u,x)+p(u,y)$ because $w(u)\geq p(u,x)$ and $p(u,y)<0$.
+When $w(u)<0$, it holds that $w(u)\geq p(\{u,x\})+p(\{u,y\})$ because $w(u)\geq p(\{u,x\})$ and $p(\{u,y\})<0$.
 If either $x\in I$ or $y\in I$, then $nb(I\cup\{u\})\geq nb(I)$.
 If both $x\notin I$ and $y\notin I$, then $nb(I\cup\{u\})<nb(I)$.
-In other word, including $u$ has positive profit if and only if either $x$ or $y$ is included.
-we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that either $x\in S^{\prime}$ or $y\in S^{\prime}$.
+In other words, including $u$ has positive profit if and only if either $x$ or $y$ is included.
+we conclude that there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that either $x\in S^{\prime}$ or $y\in S^{\prime}$.
 It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $u,x,y$ that could appear in a set in $MGIS(G)$.
-Case 2: $p(u,y)\leq w(u)<p(u,x)$. In this case, we consider whether the three vertices $u,x,y$ can be included in $I$.
-When $w(u)\geq 0$ and $w(u)\geq p(u,x)+p(u,y)$, the all three vertices may be included in $I$.
+Case 2: $p(\{u,y\})\leq w(u)<p(\{u,x\})$. In this case, we consider whether the three vertices $u,x,y$ can be included in $I$.
+When $w(u)\geq 0$ and $w(u)\geq p(\{u,x\})+p(\{u,y\})$, the all three vertices may be included in $I$.
 All combinations of $x,y,u$ may appear in $I$.
-Except that if $x$ is included and $y$ is not included in $I$, then $u\notin I$ because $nb(I\cup\{u\})=nb(I\setminus\{u\})+w(u)-p(u,x)<nb(I\setminus\{u\})$.
-In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that either $x\notin S^{\prime}$ or $y\in S^{\prime}$.
+Except that if $x$ is included and $y$ is not included in $I$, then $u\notin I$ because $nb(I\cup\{u\})=nb(I\setminus\{u\})+w(u)-p(\{u,x\})<nb(I\setminus\{u\})$.
+In other words, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that either $x\notin S^{\prime}$ or $y\in S^{\prime}$.
 It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
-When $w(u)\geq 0$ and $w(u)<p(u,x)+p(u,y)$, the three vertices can not be included in $I$.
-Moreover, $u$ and $x$ cannot be both included by $I$ because otherwise $nb(I\setminus u)>nb(I)$ contradicts the optimality of $I$.
+When $w(u)\geq 0$ and $w(u)<p(\{u,x\})+p(\{u,y\})$, the three vertices can not be included in $I$.
+Moreover, $u$ and $x$ cannot be both included by $I$ because otherwise $nb(I\setminus u)>nb(I)$ without containing permanent edges, which contradicts the optimality of $I$.
 If $x\notin I$, then $I\cup\{u\}\in MGIS(G)$ because $nb(I\cup\{u\})\geq nb(I\setminus\{u\})$.
-In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$.
+In other words, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $x\notin S^{\prime}$.
 It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
-When $w(u)<0$ and $w(u)\geq p(u,x)+p(u,y)$, the three vertices can be included in $I$.
+When $w(u)<0$ and $w(u)\geq p(\{u,x\})+p(\{u,y\})$, the three vertices can be included in $I$.
 If $y\notin I$, then $nb(I\cup\{u\})\leq nb(I\setminus\{u\})$ regardless of whether $x\in I$.
-If $y\in I$, then $nb(I\cup\{u\})\geq nb(I\setminus\{u\})$ regardless of whether $x\in I$ In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $y\in S^{\prime}$.
+If $y\in I$, then $nb(I\cup\{u\})\geq nb(I\setminus\{u\})$ regardless of whether $x\in I$ In other words, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $y\in S^{\prime}$.
 It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
-When $w(u)<0$ and $w(u)<p(u,x)+p(u,y)$, the three vertices can not be included in $I$.
-$u$ can appear in $I$ iff $y$ is included and $x$ is not included in $I$.
-In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$ and $y\in S^{\prime}$.
+When $w(u)<0$ and $w(u)<p(\{u,x\})+p(\{u,y\})$, the three vertices can not be included in $I$.
+$u$ is in $I$ iff $y$ is included and $x$ is not included in $I$.
+In other words, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $x\notin S^{\prime}$ and $y\in S^{\prime}$.
 It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
-Case 3: $w(u)<p(u,y)$.
+Case 3: $w(u)<p(\{u,y\})$.
 When $w(u)\geq 0$.
-$u$ can appear in $I$ iff neither $x$ nor $y$ is included in $I$.
-In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$ and $y\notin S^{\prime}$.
+$u$ is in at least one set $I\in MGIS(G)$ iff neither $x$ nor $y$ is included in $I$.
+In other words, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $x\notin S^{\prime}$ and $y\notin S^{\prime}$.
 It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
-When $w(u)<0$ and $w(u)<p(u,x)+p(u,y)$. $u$ cannot appear in $I$ because otherwise $nb(I\setminus\{u\})>nb(I)$ a contradiction. Therefore, we can remove $u$ directly.
-When $w(u)<0$ and $w(u)\geq p(u,x)+p(u,y)$.
-$u$ can appear in $I$ iff both $x$ and $y$ are included in $I$.
-In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\in S^{\prime}$ and $y\in S^{\prime}$.
+When $w(u)<0$ and $w(u)<p(\{u,x\})+p(\{u,y\})$.
+$u$ is not in any $I\in MGIS(G)$ because otherwise $nb(I\setminus\{u\})>nb(I)$ without containing permanent edges, a contradiction.
+Therefore, we can remove $u$ directly.
+When $w(u)<0$ and $w(u)\geq p(\{u,x\})+p(\{u,y\})$.
+$u$ is in at least one set $I\in MGIS(G)$ iff both $x$ and $y$ are included in $I$.
+In other words, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $x\in S^{\prime}$ and $y\in S^{\prime}$.
 It is easy to verify $\alpha(G^{\prime})=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
@@
 Observe that there exists a vertex set $I\in MGIS(G)$ that includes at least one of $x,u$ must be included in $I$.
-Because if none of them is included in some $I\in MGIS(G)$, then $nb(I\cup\{u\})-nb(I)\geq w(u)-\sum_{v\in N_{r}(u)}\max(0,p(u,v))\geq 0$ and therefore $I\cup\{u\}\in MGIS(G)$.
+Because if none of them is included in at least one set $I\in MGIS(G)$, then $nb(I\cup\{u\})-nb(I)\geq w(u)-\sum_{v\in N_{r}(u)}\max(0,p(\{u,v\}))\geq 0$ and therefore $I\cup\{u\}\in MGIS(G)$.
 Given an arbitrary $I\in MGIS(G)$.
-Furthermore, $x$ and $u$ cannot be both included in $I$ since $p(u,x)=M$.
+Furthermore, $x$ and $u$ cannot be both included in $I$ since $\{u,x\}\in E_{p}$.
 In summary, there are two possible states in $I$: either $x\in I,u\notin I$ or $u\in I,x\notin I$.
-Therefore, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$.
+Therefore, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $x\notin S^{\prime}$.
 It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,u$ that could appear in a set in $MGIS(G)$.
@@
 Observe that there exists a vertex set $I\in MGIS(G)$ that includes at least one of $u,x,y$.
-If none of them is included in $I$, then $nb(I\cup\{u\})-nb(I)\geq w(u)-\sum_{v\in N_{r}(u)}\max(0,p(u,v))\geq 0$ In addition, at most one of $u,x,y$ can be included in $I$ because they are permanent neighbors of each other.
+If none of them is included in $I$, then $nb(I\cup\{u\})-nb(I)\geq w(u)-\sum_{v\in N_{r}(u)}\max(0,p(\{u,v\}))\geq 0$ In addition, at most one of $u,x,y$ can be included in $I$ because $\{u,x\}\in E_{p},\{u,y\}\in E_{p},\{x,y\}\in E_{p}$.
 Given an arbitrary $I\in MGIS(G)$.
@@
 If one of $x,y$ is included in $I$,say $x$, then $nb(I\cup\{u\}\setminus\{x\})\geq nb(I)$.
-In summary, $u$ must be included in some set in $MGIS$.
+In summary, $u$ must be included in at least one set in $MGIS(G)$.
 When $\tilde{w}(y)\leq w(u)<\tilde{w}(x)$.
 If $y$ is included in $I$, then $u\notin I$ and $nb(I\cup\{u\}\setminus\{y\})\geq nb(I)$.
-Therefore, there is a vertex set $I^{\prime}\in MGIS(G)$ that $y\notin I^{\prime}$ and exactly one of $u,x$ is included in $I^{\prime}$.
-We conclude that there exists a vertex set $S\in MGIS(G)$ that $u\in S$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$.
+Therefore, there is a vertex set $I^{\prime}\in MGIS(G)$ such that $y\notin I^{\prime}$ and exactly one of $u,x$ is included in $I^{\prime}$.
+We conclude that there exists a vertex set $S\in MGIS(G)$ such that $u\in S$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $x\notin S^{\prime}$.
 It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
 When $w(u)<\tilde{w}(y)$.
-$u$ can appear in $I$ iff both $x,y$ are not included in $I$.
-In other word, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ that $x\notin S^{\prime}$ and $y\notin S^{\prime}$.
+$u$ is in at least one set $I\in MGIS(G)$ iff both $x,y$ are not included in $I$.
+In other words, there exists a vertex set $S\in MGIS(G)$ that includes $u$ iff there exists a vertex set $S^{\prime}\in MGIS(G^{\prime})$ such that $x\notin S^{\prime}$ and $y\notin S^{\prime}$.
 It is easy to verify $\alpha(G^{\prime})+w(u)=\alpha(G)$ in all combinations of $x,y,u$ that could appear in a set in $MGIS(G)$.
@@
 Therefore, $I\cup\{u\}\setminus N(u)\in MGIS(G)$.
-Since $v\in N_{p}(u)$, there must be some set in $MGIS(G)$ that does not include $v$.
-When $w(u)\geq\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}[v])+\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p(u,x))$.
+Since $v\in N_{p}(u)$, there must exist at least one set in $MGIS(G)$ that does not include $v$.
+When $w(u)\geq\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}[v])+\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p(\{u,x\}))$.
 If $v\in I$, then $u\notin I$.
-We can replace $N_{p}(u)\cap I$ with $u$ to get a larger net benefit because $nb(I\cup\{u\}\setminus N_{p}(u))-nb(I)\geq w(u)-\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}(v))+\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p(u,x))\geq w(u)-\tilde{w}(v)+w^{+}(N(u)\cap I)+\sum_{x\in N_{r}(u)\cap I}\max(0,p(u,x))\geq 0$.
-Since $v\in N_{p}(u)$, there must be some set in $MGIS(G)$ that does not include $v$.
+We can replace $N_{p}(u)\cap I$ with $u$ to get a larger net benefit because $nb(I\cup\{u\}\setminus N_{p}(u))-nb(I)\geq w(u)-\tilde{w}(v)+w^{+}(N_{p}(u)\setminus N_{p}(v))+\sum_{x\in N_{r}(u)\setminus N_{p}(v)}\max(0,p(\{u,x\}))\geq w(u)-\tilde{w}(v)+w^{+}(N(u)\cap I)+\sum_{x\in N_{r}(u)\cap I}\max(0,p(\{u,x\}))\geq 0$.
+Since $v\in N_{p}(u)$, there must exist a set in $MGIS(G)$ that does not include $v$.
 ∎
@@
 Assume, for the sake of contradiction, that neither $u$ nor $v$ is included in any set in $MGIS(G)$.
-If for any $I\in MGIS(G)$, $N(v)\cap I=\varnothing$, then $I\cup\{v\}\in MGIS(G)$ because $nb(I\cup\{v\})\geq nb(I)$, a contradiction.
+If for any $I\in MGIS(G)$, $N(v)\cap I=\varnothing$, then $I\cup\{v\}\in MGIS(G)$ because $nb(I\cup\{v\})\geq nb(I)$ without containing permanent edges, a contradiction.
 Note that $u\notin N(v)\cap I$.
 If there exists a set $I\in MGIS(G)$ such that $N(v)\cap I\neq\varnothing$, then $w(v)\geq w^{+}(N(v))-max(0,w(u))\geq w^{+}(S)$.
-$nb(I\cup\{v\}\setminus N(v))\geq nb(I)$, a contradiction.
-Therefore, there exists a set in $MGIS(G)$ that includes at least one of $u,v$. We conclude that their common permanent neighbors must not be included in some set in $MGIS(G)$. ∎
+$nb(I\cup\{v\}\setminus N(v))\geq nb(I)$ without containing permanent edges, again a contradiction.
+Therefore, there exists a set in $MGIS(G)$ that includes at least one of $u,v$.
+We conclude that $N_{p}(u)\cap N_{p}(v)$ must not be included in at least one set in $MGIS(G)$.
+∎
 Proof of R14.
@@
 If $u$ is included in $I$, then $N_{p}(v)\cap I=\varnothing$.
-And $I\cup\{v\}\in MGIS(G)$ because $nb(I\cup\{v\})-nb(I\setminus\{v\})=w(v)-\sum_{x\in N_{r}(v)}\max(0,p(v,x))\geq 0$.
+$I\cup\{v\}\in MGIS(G)$ because $nb(I\cup\{v\})-nb(I\setminus\{v\})=w(v)-\sum_{x\in N_{r}(v)}\max(0,p(\{v,x\}))\geq 0$.
 The same argument holds when $v$ is included in $I$.
 Therefore, we can fold $u$ and $v$ into a vertex $v^{\prime}$.
-There exists $S\in MGIS(G)$ that $u\in S$ and $v\in S$ iff there exists $S^{\prime}\in MGIS(G^{\prime})$ that $v^{\prime}\in S^{\prime}$.
+There exists $S\in MGIS(G)$ such that $u\in S$ and $v\in S$ iff there exists $S^{\prime}\in MGIS(G^{\prime})$ such that $v^{\prime}\in S^{\prime}$.
 ∎
@@
 $p=0.25$, Setting 1
-brock200_2_A 200 7416 2460 986 280 986 2 986 $<$1 986 $<$1 986 $<$1
-brock400_2_A 400 44847 14939 765 441 765 1 765 $<$1 765 $<$1 765 $<$1
-C125.9_A 125 5250 1713 454 $<$1 454 $<$1 454 $<$1 454 $<$1 454 $<$1
-C250.9_A 250 20940 7044 581 10 581 $<$1 581 $<$1 581 $<$1 581 $<$1
-gen200_p0.9_55_A 200 13434 4176 535 4 535 $<$1 535 $<$1 535 $<$1 535 $<$1
-gen400_p0.9_75_A 400 53671 18149 628 82 628 $<$1 628 3 628 $<$1 628 $<$1
-hamming8-4_A 256 15580 5284 1094 240 1094 $<$1 1094 $<$1 1094 $<$1 1094 $<$1
-keller4_A 171 7057 2378 941 15 941 $<$1 941 $<$1 941 $<$1 941 $<$1
-MANN_a27_A 378 53050 17501 533 28 533 $<$1 533 $<$1 533 $<$1 533 $<$1
-p_hat300-1_A 300 8239 2694 2308 10800 2680 10800 2744 $<$1 2744 $<$1 2744 $<$1
-p_hat300-2_A 300 16403 5525 2076 5182 2076 608 2076 $<$1 2076 $<$1 2076 $<$1
-p_hat300-3_A 300 25132 8258 739 100 739 $<$1 739 $<$1 739 $<$1 739 $<$1
-brock200_2_B 200 7416 2460 962 276 962 2 962 $<$1 962 $<$1 962 $<$1
-brock400_2_B 400 44847 14939 741 445 741 1 741 $<$1 741 $<$1 741 $<$1
-C125.9_B 125 5250 1713 437 1 437 $<$1 437 $<$1 437 $<$1 437 $<$1
-C250.9_B 250 20940 7044 549 11 549 $<$1 549 $<$1 549 $<$1 549 $<$1
-gen200_p0.9_55_B 200 13434 4176 510 5 510 $<$1 510 $<$1 510 $<$1 510 $<$1
-gen400_p0.9_75_B 400 53671 18149 595 80 595 $<$1 595 $<$1 595 $<$1 595 $<$1
-hamming8-4_B 256 15580 5284 1094 234 1094 $<$1 1094 $<$1 1094 $<$1 1094 $<$1
-keller4_B 171 7057 2378 941 15 941 $<$1 941 $<$1 941 $<$1 941 $<$1
-MANN_a27_B 378 53050 17501 503 31 503 $<$1 503 $<$1 503 $<$1 503 $<$1
-p_hat300-1_B 300 8239 2694 2259 10800 2643 10800 2712 $<$1 2712 $<$1 2712 $<$1
-p_hat300-2_B 300 16403 5525 2062 5194 2062 680 2062 $<$1 2062 $<$1 2062 $<$1
-p_hat300-3_B 300 25132 8258 713 102 713 $<$1 713 $<$1 713 $<$1 713 $<$1
-brock200_2_C 200 7416 2460 932 280 932 2 932 $<$1 932 $<$1 932 $<$1
-brock400_2_C 400 44847 14939 698 455 698 2 698 $<$1 698 $<$1 698 $<$1
-C125.9_C 125 5250 1713 403 1 403 $<$1 403 $<$1 403 $<$1 403 $<$1
-C250.9_C 250 20940 7044 502 12 502 $<$1 502 $<$1 502 $<$1 502 $<$1
-gen200_p0.9_55_C 200 13434 4176 467 4 467 $<$1 467 $<$1 467 $<$1 467 $<$1
-gen400_p0.9_75_C 400 53671 18149 533 81 533 $<$1 533 $<$1 533 $<$1 533 $<$1
-hamming8-4_C 256 15580 5284 1094 229 1094 1 1094 $<$1 1094 $<$1 1094 $<$1
-keller4_C 171 7057 2378 941 14 941 $<$1 941 $<$1 941 $<$1 941 $<$1
-MANN_a27_C 378 53050 17501 443 30 443 $<$1 443 $<$1 443 $<$1 443 $<$1
-p_hat300-1_C 300 8239 2694 2205 10800 2566 10800 2649 $<$1 2649 $<$1 2649 $<$1
-p_hat300-2_C 300 16403 5525 2033 5360 2033 778 2033 $<$1 2033 $<$1 2033 $<$1
-p_hat300-3_C 300 25132 8258 688 111 688 $<$1 688 $<$1 688 $<$1 688 $<$1
+brock200_2_A 200 7,416 2,460 986 280 986 2 986 <<1 986 <<1 986 <<1
+brock400_2_A 400 44,847 14,939 765 441 765 1 765 <<1 765 <<1 765 <<1
+C125.9_A 125 5,250 1,713 454 <<1 454 <<1 454 <<1 454 <<1 454 <<1
+C250.9_A 250 20,940 7,044 581 10 581 <<1 581 <<1 581 <<1 581 <<1
+gen200_p0.9_55_A 200 13,434 4,176 535 4 535 <<1 535 <<1 535 <<1 535 <<1
+gen400_p0.9_75_A 400 53,671 18,149 628 82 628 <<1 628 3 628 <<1 628 <<1
+hamming8-4_A 256 15,580 5,284 1,094 240 1,094 <<1 1,094 <<1 1,094 <<1 1,094 <<1
+keller4_A 171 7,057 2,378 941 15 941 <<1 941 <<1 941 <<1 941 <<1
+MANN_a27_A 378 53,050 17,501 533 28 533 <<1 533 <<1 533 <<1 533 <<1
+p_hat300-1_A 300 8,239 2,694 2,308 10,800 2,680 10,800 2,744 <<1 2,744 <<1 2,744 <<1
+p_hat300-2_A 300 16,403 5,525 2,076 5,182 2,076 608 2,076 <<1 2,076 <<1 2,076 <<1
+p_hat300-3_A 300 25,132 8,258 739 100 739 <<1 739 <<1 739 <<1 739 <<1
+brock200_2_B 200 7,416 2,460 962 276 962 2 962 <<1 962 <<1 962 <<1
+brock400_2_B 400 44,847 14,939 741 445 741 1 741 <<1 741 <<1 741 <<1
+C125.9_B 125 5,250 1,713 437 1 437 <<1 437 <<1 437 <<1 437 <<1
+C250.9_B 250 20,940 7,044 549 11 549 <<1 549 <<1 549 <<1 549 <<1
+gen200_p0.9_55_B 200 13,434 4,176 510 5 510 <<1 510 <<1 510 <<1 510 <<1
+gen400_p0.9_75_B 400 53,671 18,149 595 80 595 <<1 595 <<1 595 <<1 595 <<1
+hamming8-4_B 256 15,580 5,284 1,094 234 1,094 <<1 1,094 <<1 1,094 <<1 1,094 <<1
+keller4_B 171 7,057 2,378 941 15 941 <<1 941 <<1 941 <<1 941 <<1
+MANN_a27_B 378 53,050 17,501 503 31 503 <<1 503 <<1 503 <<1 503 <<1
+p_hat300-1_B 300 8,239 2,694 2,259 10,800 2,643 10,800 2,712 <<1 2,712 <<1 2,712 <<1
+p_hat300-2_B 300 16,403 5,525 2,062 5,194 2,062 680 2,062 <<1 2,062 <<1 2,062 <<1
+p_hat300-3_B 300 25,132 8,258 713 102 713 <<1 713 <<1 713 <<1 713 <<1
+brock200_2_C 200 7,416 2,460 932 280 932 2 932 <<1 932 <<1 932 <<1
+brock400_2_C 400 44,847 14,939 698 455 698 2 698 <<1 698 <<1 698 <<1
+C125.9_C 125 5,250 1,713 403 1 403 <<1 403 <<1 403 <<1 403 <<1
+C250.9_C 250 20,940 7,044 502 12 502 <<1 502 <<1 502 <<1 502 <<1
+gen200_p0.9_55_C 200 13,434 4,176 467 4 467 <<1 467 <<1 467 <<1 467 <<1
+gen400_p0.9_75_C 400 53,671 18,149 533 81 533 <<1 533 <<1 533 <<1 533 <<1
+hamming8-4_C 256 15,580 5,284 1,094 229 1,094 1 1,094 <<1 1,094 <<1 1,094 <<1
+keller4_C 171 7,057 2,378 941 14 941 <<1 941 <<1 941 <<1 941 <<1
+MANN_a27_C 378 53,050 17,501 443 30 443 <<1 443 <<1 443 <<1 443 <<1
+p_hat300-1_C 300 8,239 2,694 2,205 10,800 2,566 10,800 2,649 <<1 2,649 <<1 2,649 <<1
+p_hat300-2_C 300 16,403 5,525 2,033 5,360 2,033 778 2,033 <<1 2,033 <<1 2,033 <<1
+p_hat300-3_C 300 25,132 8,258 688 111 688 <<1 688 <<1 688 <<1 688 <<1
 $p=0.25$, Setting 2
-brock200_2_A 200 7416 2460 1489 452 1489 5 1489 $<$1 1489 $<$1 1489 $<$1
-brock400_2_A 400 44847 14939 1084 593 1084 2 1084 2 1084 $<$1 1084 $<$1
-C125.9_A 125 5250 1713 685 1 685 $<$1 685 $<$1 685 $<$1 685 $<$1
-C250.9_A 250 20940 7044 785 12 785 $<$1 785 $<$1 785 $<$1 785 $<$1
-gen200_p0.9_55_A 200 13434 4176 778 5 778 $<$1 778 $<$1 778 $<$1 778 $<$1
-gen400_p0.9_75_A 400 53671 18149 882 94 882 $<$1 882 $<$1 882 $<$1 882 $<$1
-hamming8-4_A 256 15580 5284 1790 331 1790 2 1790 $<$1 1790 $<$1 1790 $<$1
-keller4_A 171 7057 2378 1500 23 1500 $<$1 1500 $<$1 1500 $<$1 1500 $<$1
-MANN_a27_A 378 53050 17501 683 32 683 $<$1 683 $<$1 683 $<$1 683 $<$1
-p_hat300-1_A 300 8239 2694 3981 10800 4182 10800 4674 1 4674 $<$1 4674 $<$1
-p_hat300-2_A 300 16403 5525 2994 8743 2994 9402 2994 $<$1 2994 $<$1 2994 $<$1
-p_hat300-3_A 300 25132 8258 1180 137 1180 1 1180 $<$1 1180 $<$1 1180 $<$1
-brock200_2_B 200 7416 2460 739 456 739 5 739 $<$1 739 $<$1 739 $<$1
-brock400_2_B 400 44847 14939 534 611 534 2 534 2 534 $<$1 534 $<$1
-C125.9_B 125 5250 1713 335 1 335 $<$1 335 $<$1 335 $<$1 335 $<$1
-C250.9_B 250 20940 7044 385 13 385 $<$1 385 $<$1 385 $<$1 385 $<$1
-gen200_p0.9_55_B 200 13434 4176 378 5 378 $<$1 378 $<$1 378 $<$1 378 $<$1
-gen400_p0.9_75_B 400 53671 18149 432 96 432 $<$1 432 $<$1 432 $<$1 432 $<$1
-hamming8-4_B 256 15580 5284 890 333 890 2 890 $<$1 890 $<$1 890 $<$1
-keller4_B 171 7057 2378 750 22 750 $<$1 750 $<$1 750 $<$1 750 $<$1
-MANN_a27_B 378 53050 17501 333 34 333 $<$1 333 $<$1 333 $<$1 333 $<$1
-p_hat300-1_B 300 8239 2694 1981 10800 2082 10800 2324 2 2324 $<$1 2324 $<$1
-p_hat300-2_B 300 16403 5525 1494 8947 1494 10800 1494 $<$1 1494 $<$1 1494 $<$1
-p_hat300-3_B 300 25132 8258 580 134 580 1 580 1 580 $<$1 580 $<$1
-brock200_2_C 200 7416 2460 364 460 364 5 364 $<$1 364 $<$1 364 $<$1
-brock400_2_C 400 44847 14939 259 646 259 2 259 1 259 $<$1 259 $<$1
-C125.9_C 125 5250 1713 160 1 160 $<$1 160 $<$1 160 $<$1 160 $<$1
-C250.9_C 250 20940 7044 185 13 185 $<$1 185 $<$1 185 $<$1 185 $<$1
-gen200_p0.9_55_C 200 13434 4176 178 5 178 $<$1 178 $<$1 178 $<$1 178 $<$1
-gen400_p0.9_75_C 400 53671 18149 207 98 207 $<$1 207 $<$1 207 $<$1 207 $<$1
-hamming8-4_C 256 15580 5284 440 355 440 2 440 $<$1 440 $<$1 440 $<$1
-keller4_C 171 7057 2378 375 23 375 $<$1 375 $<$1 375 $<$1 375 $<$1
-MANN_a27_C 378 53050 17501 158 33 158 $<$1 158 $<$1 158 $<$1 158 $<$1
-p_hat300-1_C 300 8239 2694 981 10800 1032 10800 1149 1 1149 $<$1 1149 $<$1
-p_hat300-2_C 300 16403 5525 744 10800 744 10658 744 $<$1 744 $<$1 744 $<$1
-p_hat300-3_C 300 25132 8258 280 142 280 $<$1 280 $<$1 280 $<$1 280 $<$1
+brock200_2_A 200 7,416 2,460 1,489 452 1,489 5 1,489 <<1 1,489 <<1 1,489 <<1
+brock400_2_A 400 44,847 14,939 1,084 593 1,084 2 1,084 2 1,084 <<1 1,084 <<1
+C125.9_A 125 5,250 1,713 685 1 685 <<1 685 <<1 685 <<1 685 <<1
+C250.9_A 250 20,940 7,044 785 12 785 <<1 785 <<1 785 <<1 785 <<1
+gen200_p0.9_55_A 200 13,434 4,176 778 5 778 <<1 778 <<1 778 <<1 778 <<1
+gen400_p0.9_75_A 400 53,671 18,149 882 94 882 <<1 882 <<1 882 <<1 882 <<1
+hamming8-4_A 256 15,580 5,284 1,790 331 1,790 2 1,790 <<1 1,790 <<1 1,790 <<1
+keller4_A 171 7,057 2,378 1,500 23 1,500 <<1 1,500 <<1 1,500 <<1 1,500 <<1
+MANN_a27_A 378 53,050 17,501 683 32 683 <<1 683 <<1 683 <<1 683 <<1
+p_hat300-1_A 300 8,239 2,694 3,981 10,800 4,182 10,800 4,674 1 4,674 <<1 4,674 <<1
+p_hat300-2_A 300 16,403 5,525 2,994 8,743 2,994 9,402 2,994 <<1 2,994 <<1 2,994 <<1
+p_hat300-3_A 300 25,132 8,258 1,180 137 1,180 1 1,180 <<1 1,180 <<1 1,180 <<1
+brock200_2_B 200 7,416 2,460 739 456 739 5 739 <<1 739 <<1 739 <<1
+brock400_2_B 400 44,847 14,939 534 611 534 2 534 2 534 <<1 534 <<1
+C125.9_B 125 5,250 1,713 335 1 335 <<1 335 <<1 335 <<1 335 <<1
+C250.9_B 250 20,940 7,044 385 13 385 <<1 385 <<1 385 <<1 385 <<1
+gen200_p0.9_55_B 200 13,434 4,176 378 5 378 <<1 378 <<1 378 <<1 378 <<1
+gen400_p0.9_75_B 400 53,671 18,149 432 96 432 <<1 432 <<1 432 <<1 432 <<1
+hamming8-4_B 256 15,580 5,284 890 333 890 2 890 <<1 890 <<1 890 <<1
+keller4_B 171 7,057 2,378 750 22 750 <<1 750 <<1 750 <<1 750 <<1
+MANN_a27_B 378 53,050 17,501 333 34 333 <<1 333 <<1 333 <<1 333 <<1
+p_hat300-1_B 300 8,239 2,694 1,981 10,800 2,082 10,800 2,324 2 2,324 <<1 2,324 <<1
+p_hat300-2_B 300 16,403 5,525 1,494 8,947 1,494 10,800 1,494 <<1 1,494 <<1 1,494 <<1
+p_hat300-3_B 300 25,132 8,258 580 134 580 1 580 1 580 <<1 580 <<1
+brock200_2_C 200 7,416 2,460 364 460 364 5 364 <<1 364 <<1 364 <<1
+brock400_2_C 400 44,847 14,939 259 646 259 2 259 1 259 <<1 259 <<1
+C125.9_C 125 5,250 1,713 160 1 160 <<1 160 <<1 160 <<1 160 <<1
+C250.9_C 250 20,940 7,044 185 13 185 <<1 185 <<1 185 <<1 185 <<1
+gen200_p0.9_55_C 200 13,434 4,176 178 5 178 <<1 178 <<1 178 <<1 178 <<1
+gen400_p0.9_75_C 400 53,671 18,149 207 98 207 <<1 207 <<1 207 <<1 207 <<1
+hamming8-4_C 256 15,580 5,284 440 355 440 2 440 <<1 440 <<1 440 <<1
+keller4_C 171 7,057 2,378 375 23 375 <<1 375 <<1 375 <<1 375 <<1
+MANN_a27_C 378 53,050 17,501 158 33 158 <<1 158 <<1 158 <<1 158 <<1
+p_hat300-1_C 300 8,239 2,694 981 10,800 1,032 10,800 1,149 1 1,149 <<1 1,149 <<1
+p_hat300-2_C 300 16,403 5,525 744 10,800 744 10,658 744 <<1 744 <<1 744 <<1
+p_hat300-3_C 300 25,132 8,258 280 142 280 <<1 280 <<1 280 <<1 280 <<1
 $p=0.50$, Setting 1
-brock200_2_A 200 4916 4960 1298 8390 1298 93 1298 $<$1 1298 $<$1 1298 $<$1
-brock400_2_A 400 29711 30075 1103 10800 1123 355 1123 $<$1 1123 $<$1 1123 $<$1
-C125.9_A 125 3500 3463 627 5 627 $<$1 627 $<$1 627 $<$1 627 $<$1
-C250.9_A 250 14017 13967 817 266 817 1 817 $<$1 817 $<$1 817 $<$1
-gen200_p0.9_55_A 200 8908 9002 785 68 785 $<$1 785 $<$1 785 $<$1 785 $<$1
-gen400_p0.9_75_A 400 35823 35997 895 5699 895 29 895 $<$1 895 $<$1 895 $<$1
-hamming8-4_A 256 10329 10535 1301 7608 1301 49 1301 $<$1 1301 $<$1 1301 $<$1
-keller4_A 171 4738 4697 1118 153 1118 2 1118 $<$1 1118 $<$1 1118 $<$1
-MANN_a27_A 378 35345 35206 812 1217 812 6 812 $<$1 812 $<$1 812 $<$1
-p_hat300-1_A 300 5505 5428 2568 10800 3026 10800 3129 $<$1 3129 $<$1 3129 $<$1
-p_hat300-2_A 300 11051 10877 2063 10800 2477 10800 2477 $<$1 2477 $<$1 2477 $<$1
-p_hat300-3_A 300 16820 16570 1029 3748 1029 38 1029 $<$1 1029 $<$1 1029 $<$1
-brock200_2_B 200 4916 4960 1224 9589 1224 125 1224 $<$1 1224 $<$1 1224 $<$1
-brock400_2_B 400 29711 30075 1010 10800 1035 445 1035 $<$1 1035 $<$1 1035 $<$1
-C125.9_B 125 3500 3463 582 5 582 $<$1 582 $<$1 582 $<$1 582 $<$1
-C250.9_B 250 14017 13967 744 292 744 2 744 $<$1 744 $<$1 744 $<$1
-gen200_p0.9_55_B 200 8908 9002 716 72 716 $<$1 716 $<$1 716 $<$1 716 $<$1
-gen400_p0.9_75_B 400 35823 35997 805 6602 805 34 805 $<$1 805 $<$1 805 $<$1
-hamming8-4_B 256 10329 10535 1255 7780 1255 62 1255 $<$1 1255 $<$1 1255 $<$1
-keller4_B 171 4738 4697 1094 156 1094 2 1094 $<$1 1094 $<$1 1094 $<$1
-MANN_a27_B 378 35345 35206 707 1417 707 7 707 1 707 $<$1 707 $<$1
-p_hat300-1_B 300 5505 5428 2492 10800 2950 10800 3023 $<$1 3023 $<$1 3023 $<$1
-p_hat300-2_B 300 11051 10877 1979 10800 2405 10800 2405 $<$1 2405 $<$1 2405 $<$1
-p_hat300-3_B 300 16820 16570 967 4067 967 46 967 $<$1 967 $<$1 967 $<$1
-brock200_2_C 200 4916 4960 1101 10800 1101 204 1101 $<$1 1101 $<$1 1101 $<$1
-brock400_2_C 400 29711 30075 861 10800 892 691 892 $<$1 892 $<$1 892 $<$1
-C125.9_C 125 3500 3463 506 6 506 $<$1 506 $<$1 506 $<$1 506 $<$1
-C250.9_C 250 14017 13967 623 334 623 2 623 $<$1 623 $<$1 623 $<$1
-gen200_p0.9_55_C 200 8908 9002 597 91 597 $<$1 597 $<$1 597 $<$1 597 $<$1
-gen400_p0.9_75_C 400 35823 35997 651 7858 651 50 651 $<$1 651 $<$1 651 $<$1
-hamming8-4_C 256 10329 10535 1184 7447 1184 85 1184 $<$1 1184 $<$1 1184 $<$1
-keller4_C 171 4738 4697 1049 145 1049 3 1049 $<$1 1049 $<$1 1049 $<$1
-MANN_a27_C 378 35345 35206 552 1849 552 10 552 1 552 $<$1 552 $<$1
-p_hat300-1_C 300 5505 5428 2303 10800 2818 10800 2897 $<$1 2897 $<$1 2897 $<$1
-p_hat300-2_C 300 11051 10877 1852 10800 2263 10800 2263 $<$1 2263 $<$1 2263 $<$1
-p_hat300-3_C 300 16820 16570 851 4709 851 69 851 $<$1 851 $<$1 851 $<$1
+brock200_2_A 200 4,916 4,960 1,298 8,390 1,298 93 1,298 <<1 1,298 <<1 1,298 <<1
+brock400_2_A 400 29,711 30,075 1,103 10,800 1,123 355 1,123 <<1 1,123 <<1 1,123 <<1
+C125.9_A 125 3,500 3,463 627 5 627 <<1 627 <<1 627 <<1 627 <<1
+C250.9_A 250 14,017 13,967 817 266 817 1 817 <<1 817 <<1 817 <<1
+gen200_p0.9_55_A 200 8,908 9,002 785 68 785 <<1 785 <<1 785 <<1 785 <<1
+gen400_p0.9_75_A 400 35,823 35,997 895 5,699 895 29 895 <<1 895 <<1 895 <<1
+hamming8-4_A 256 10,329 10,535 1,301 7,608 1,301 49 1,301 <<1 1,301 <<1 1,301 <<1
+keller4_A 171 4,738 4,697 1,118 153 1,118 2 1,118 <<1 1,118 <<1 1,118 <<1
+MANN_a27_A 378 35,345 35,206 812 1,217 812 6 812 <<1 812 <<1 812 <<1
+p_hat300-1_A 300 5,505 5,428 2,568 10,800 3,026 10,800 3,129 <<1 3,129 <<1 3,129 <<1
+p_hat300-2_A 300 11,051 10,877 2,063 10,800 2,477 10,800 2,477 <<1 2,477 <<1 2,477 <<1
+p_hat300-3_A 300 16,820 16,570 1,029 3,748 1,029 38 1,029 <<1 1,029 <<1 1,029 <<1
+brock200_2_B 200 4,916 4,960 1,224 9,589 1,224 125 1,224 <<1 1,224 <<1 1,224 <<1
+brock400_2_B 400 29,711 30,075 1,010 10,800 1,035 445 1,035 <<1 1,035 <<1 1,035 <<1
+C125.9_B 125 3,500 3,463 582 5 582 <<1 582 <<1 582 <<1 582 <<1
+C250.9_B 250 14,017 13,967 744 292 744 2 744 <<1 744 <<1 744 <<1
+gen200_p0.9_55_B 200 8,908 9,002 716 72 716 <<1 716 <<1 716 <<1 716 <<1
+gen400_p0.9_75_B 400 35,823 35,997 805 6,602 805 34 805 <<1 805 <<1 805 <<1
+hamming8-4_B 256 10,329 10,535 1,255 7,780 1,255 62 1,255 <<1 1,255 <<1 1,255 <<1
+keller4_B 171 4,738 4,697 1,094 156 1,094 2 1,094 <<1 1,094 <<1 1,094 <<1
+MANN_a27_B 378 35,345 35,206 707 1,417 707 7 707 1 707 <<1 707 <<1
+p_hat300-1_B 300 5,505 5,428 2,492 10,800 2,950 10,800 3,023 <<1 3,023 <<1 3,023 <<1
+p_hat300-2_B 300 11,051 10,877 1,979 10,800 2,405 10,800 2,405 <<1 2,405 <<1 2,405 <<1
+p_hat300-3_B 300 16,820 16,570 967 4,067 967 46 967 <<1 967 <<1 967 <<1
+brock200_2_C 200 4,916 4,960 1,101 10,800 1,101 204 1,101 <<1 1,101 <<1 1,101 <<1
+brock400_2_C 400 29,711 30,075 861 10,800 892 691 892 <<1 892 <<1 892 <<1
+C125.9_C 125 3,500 3,463 506 6 506 <<1 506 <<1 506 <<1 506 <<1
+C250.9_C 250 14,017 13,967 623 334 623 2 623 <<1 623 <<1 623 <<1
+gen200_p0.9_55_C 200 8,908 9,002 597 91 597 <<1 597 <<1 597 <<1 597 <<1
+gen400_p0.9_75_C 400 35,823 35,997 651 7,858 651 50 651 <<1 651 <<1 651 <<1
+hamming8-4_C 256 10,329 10,535 1,184 7,447 1,184 85 1,184 <<1 1,184 <<1 1,184 <<1
+keller4_C 171 4,738 4,697 1,049 145 1,049 3 1,049 <<1 1,049 <<1 1,049 <<1
+MANN_a27_C 378 35,345 35,206 552 1,849 552 10 552 1 552 <<1 552 <<1
+p_hat300-1_C 300 5,505 5,428 2,303 10,800 2,818 10,800 2,897 <<1 2,897 <<1 2,897 <<1
+p_hat300-2_C 300 11,051 10,877 1,852 10,800 2,263 10,800 2,263 <<1 2,263 <<1 2,263 <<1
+p_hat300-3_C 300 16,820 16,570 851 4,709 851 69 851 <<1 851 <<1 851 <<1
 $p=0.50$, Setting 2
-brock200_2_A 200 4916 4960 2034 10800 2034 677 2034 $<$1 2034 $<$1 2034 $<$1
-brock400_2_A 400 29711 30075 1628 10800 1630 1323 1630 11 1630 $<$1 1630 $<$1
-C125.9_A 125 3500 3463 1152 7 1152 $<$1 1152 $<$1 1152 $<$1 1152 $<$1
-C250.9_A 250 14017 13967 1236 343 1236 3 1236 3 1236 $<$1 1236 $<$1
-gen200_p0.9_55_A 200 8908 9002 1151 113 1151 $<$1 1151 3 1151 $<$1 1151 $<$1
-gen400_p0.9_75_A 400 35823 35997 1335 8467 1335 69 1335 2 1335 $<$1 1335 $<$1
-hamming8-4_A 256 10329 10535 2155 10800 2155 166 2155 1 2155 $<$1 2155 $<$1
-keller4_A 171 4738 4697 1759 353 1759 5 1759 $<$1 1759 $<$1 1759 $<$1
-MANN_a27_A 378 35345 35206 1226 1950 1226 10 1226 12 1226 $<$1 1226 $<$1
-p_hat300-1_A 300 5505 5428 4740 10800 5637 10800 5637 $<$1 5637 $<$1 5637 $<$1
-p_hat300-2_A 300 11051 10877 3342 10800 3750 10800 3943 $<$1 3943 $<$1 3943 $<$1
-p_hat300-3_A 300 16820 16570 1658 7243 1658 157 1658 $<$1 1658 $<$1 1658 $<$1
-brock200_2_B 200 4916 4960 984 10800 984 757 984 $<$1 984 $<$1 984 $<$1
-brock400_2_B 400 29711 30075 765 10800 780 1385 780 10 780 $<$1 780 $<$1
-C125.9_B 125 3500 3463 552 7 552 $<$1 552 $<$1 552 $<$1 552 $<$1
-C250.9_B 250 14017 13967 586 391 586 3 586 4 586 $<$1 586 $<$1
-gen200_p0.9_55_B 200 8908 9002 551 121 551 $<$1 551 1 551 $<$1 551 $<$1
-gen400_p0.9_75_B 400 35823 35997 635 9395 635 72 635 $<$1 635 $<$1 635 $<$1
-hamming8-4_B 256 10329 10535 1055 10800 1055 163 1055 $<$1 1055 $<$1 1055 $<$1
-keller4_B 171 4738 4697 859 367 859 5 859 $<$1 859 $<$1 859 $<$1
-MANN_a27_B 378 35345 35206 576 2161 576 11 576 13 576 $<$1 576 $<$1
-p_hat300-1_B 300 5505 5428 2340 10800 2325 10800 2787 $<$1 2787 $<$1 2787 $<$1
-p_hat300-2_B 300 11051 10877 1642 10800 1850 10800 1943 $<$1 1943 $<$1 1943 $<$1
-p_hat300-3_B 300 16820 16570 808 7790 808 161 808 $<$1 808 $<$1 808 $<$1
-brock200_2_C 200 4916 4960 459 10800 459 934 459 $<$1 459 $<$1 459 $<$1
-brock400_2_C 400 29711 30075 341 10800 355 1711 355 12 355 $<$1 355 $<$1
-C125.9_C 125 3500 3463 252 8 252 $<$1 252 $<$1 252 $<$1 252 $<$1
-C250.9_C 250 14017 13967 261 501 261 3 261 3 261 $<$1 261 $<$1
-gen200_p0.9_55_C 200 8908 9002 251 146 251 1 251 2 251 $<$1 251 $<$1
-gen400_p0.9_75_C 400 35823 35997 285 10800 285 83 285 2 285 $<$1 285 $<$1
-hamming8-4_C 256 10329 10535 505 10800 505 170 505 $<$1 505 $<$1 505 $<$1
-keller4_C 171 4738 4697 409 389 409 5 409 $<$1 409 $<$1 409 $<$1
-MANN_a27_C 378 35345 35206 251 2666 251 13 251 10 251 $<$1 251 $<$1
-p_hat300-1_C 300 5505 5428 1138 10800 1125 10800 1362 $<$1 1362 $<$1 1362 $<$1
-p_hat300-2_C 300 11051 10877 788 10800 900 10800 943 $<$1 943 $<$1 943 $<$1
-p_hat300-3_C 300 16820 16570 383 9014 383 178 383 $<$1 383 $<$1 383 $<$1
+brock200_2_A 200 4,916 4,960 2,034 10,800 2,034 677 2,034 <<1 2,034 <<1 2,034 <<1
+brock400_2_A 400 29,711 30,075 1,628 10,800 1,630 1,323 1,630 11 1,630 <<1 1,630 <<1
+C125.9_A 125 3,500 3,463 1,152 7 1,152 <<1 1,152 <<1 1,152 <<1 1,152 <<1
+C250.9_A 250 14,017 13,967 1,236 343 1,236 3 1,236 3 1,236 <<1 1,236 <<1
+gen200_p0.9_55_A 200 8,908 9,002 1,151 113 1,151 <<1 1,151 3 1,151 <<1 1,151 <<1
+gen400_p0.9_75_A 400 35,823 35,997 1,335 8,467 1,335 69 1,335 2 1,335 <<1 1,335 <<1
+hamming8-4_A 256 10,329 10,535 2,155 10,800 2,155 166 2,155 1 2,155 <<1 2,155 <<1
+keller4_A 171 4,738 4,697 1,759 353 1,759 5 1,759 <<1 1,759 <<1 1,759 <<1
+MANN_a27_A 378 35,345 35,206 1,226 1,950 1,226 10 1,226 12 1,226 <<1 1,226 <<1
+p_hat300-1_A 300 5,505 5,428 4,740 10,800 5,637 10,800 5,637 <<1 5,637 <<1 5,637 <<1
+p_hat300-2_A 300 11,051 10,877 3,342 10,800 3,750 10,800 3,943 <<1 3,943 <<1 3,943 <<1
+p_hat300-3_A 300 16,820 16,570 1,658 7,243 1,658 157 1,658 <<1 1,658 <<1 1,658 <<1
+brock200_2_B 200 4,916 4,960 984 10,800 984 757 984 <<1 984 <<1 984 <<1
+brock400_2_B 400 29,711 30,075 765 10,800 780 1,385 780 10 780 <<1 780 <<1
+C125.9_B 125 3,500 3,463 552 7 552 <<1 552 <<1 552 <<1 552 <<1
+C250.9_B 250 14,017 13,967 586 391 586 3 586 4 586 <<1 586 <<1
+gen200_p0.9_55_B 200 8,908 9,002 551 121 551 <<1 551 1 551 <<1 551 <<1
+gen400_p0.9_75_B 400 35,823 35,997 635 9,395 635 72 635 <<1 635 <<1 635 <<1
+hamming8-4_B 256 10,329 10,535 1,055 10,800 1,055 163 1,055 <<1 1,055 <<1 1,055 <<1
+keller4_B 171 4,738 4,697 859 367 859 5 859 <<1 859 <<1 859 <<1
+MANN_a27_B 378 35,345 35,206 576 2,161 576 11 576 13 576 <<1 576 <<1
+p_hat300-1_B 300 5,505 5,428 2,340 10,800 2,325 10,800 2,787 <<1 2,787 <<1 2,787 <<1
+p_hat300-2_B 300 11,051 10,877 1,642 10,800 1,850 10,800 1,943 <<1 1,943 <<1 1,943 <<1
+p_hat300-3_B 300 16,820 16,570 808 7,790 808 161 808 <<1 808 <<1 808 <<1
+brock200_2_C 200 4,916 4,960 459 10,800 459 934 459 <<1 459 <<1 459 <<1
+brock400_2_C 400 29,711 30,075 341 10,800 355 1,711 355 12 355 <<1 355 <<1
+C125.9_C 125 3,500 3,463 252 8 252 <<1 252 <<1 252 <<1 252 <<1
+C250.9_C 250 14,017 13,967 261 501 261 3 261 3 261 <<1 261 <<1
+gen200_p0.9_55_C 200 8,908 9,002 251 146 251 1 251 2 251 <<1 251 <<1
+gen400_p0.9_75_C 400 35,823 35,997 285 10,800 285 83 285 2 285 <<1 285 <<1
+hamming8-4_C 256 10,329 10,535 505 10,800 505 170 505 <<1 505 <<1 505 <<1
+keller4_C 171 4,738 4,697 409 389 409 5 409 <<1 409 <<1 409 <<1
+MANN_a27_C 378 35,345 35,206 251 2,666 251 13 251 10 251 <<1 251 <<1
+p_hat300-1_C 300 5,505 5,428 1,138 10,800 1,125 10,800 1,362 <<1 1,362 <<1 1,362 <<1
+p_hat300-2_C 300 11,051 10,877 788 10,800 900 10,800 943 <<1 943 <<1 943 <<1
+p_hat300-3_C 300 16,820 16,570 383 9,014 383 178 383 <<1 383 <<1 383 <<1
 $p=0.75$, Setting 1
-brock200_2_A 200 2438 7438 1615 10800 1851 10800 1885 $<$1 1885 $<$1 1885 $<$1
-brock400_2_A 400 14751 45035 1403 10800 1667 10800 1728 5 1728 $<$1 1728 $<$1
-C125.9_A 125 1733 5230 1023 247 1023 4 1023 $<$1 1023 $<$1 1023 $<$1
-C250.9_A 250 7073 20911 1200 10800 1236 5143 1236 2 1236 $<$1 1236 $<$1
-gen200_p0.9_55_A 200 4425 13485 1186 10800 1206 601 1206 $<$1 1206 $<$1 1206 $<$1
-gen400_p0.9_75_A 400 18063 53757 1155 10800 1490 10800 1490 $<$1 1490 $<$1 1490 $<$1
-hamming8-4_A 256 5173 15691 1549 10800 1723 10800 1759 $<$1 1759 $<$1 1759 $<$1
-keller4_A 171 2400 7035 1400 10800 1434 1573 1434 $<$1 1434 $<$1 1434 $<$1
-MANN_a27_A 378 17580 52971 1313 10800 1323 10800 1323 $<$1 1323 $<$1 1323 $<$1
-p_hat300-1_A 300 2734 8199 2504 10800 3932 10800 4164 $<$1 4164 $<$1 4164 $<$1
-p_hat300-2_A 300 5603 16325 2120 10800 2902 10800 2990 1 2990 $<$1 2990 $<$1
-p_hat300-3_A 300 8388 25002 1350 10800 1561 10800 1564 $<$1 1564 $<$1 1564 $<$1
-brock200_2_B 200 2438 7438 1416 10800 1602 10800 1641 $<$1 1641 $<$1 1641 $<$1
-brock400_2_B 400 14751 45035 1129 10800 1341 10800 1386 2 1386 $<$1 1386 $<$1
-C125.9_B 125 1733 5230 856 395 856 8 856 $<$1 856 $<$1 856 $<$1
-C250.9_B 250 7073 20911 960 10800 1001 10800 1001 1 1001 $<$1 1001 $<$1
-gen200_p0.9_55_B 200 4425 13485 961 10800 983 1585 983 $<$1 983 $<$1 983 $<$1
-gen400_p0.9_75_B 400 18063 53757 920 10800 1105 10800 1120 $<$1 1120 $<$1 1120 $<$1
-hamming8-4_B 256 5173 15691 1356 10800 1486 10800 1579 $<$1 1579 $<$1 1579 $<$1
-keller4_B 171 2400 7035 1230 10800 1268 4468 1268 $<$1 1268 $<$1 1268 $<$1
-MANN_a27_B 378 17580 52971 922 10800 1011 10800 1021 $<$1 1021 $<$1 1021 $<$1
-p_hat300-1_B 300 2734 8199 2306 10800 3547 10800 3886 $<$1 3886 $<$1 3886 $<$1
-p_hat300-2_B 300 5603 16325 1762 10800 2588 10800 2782 $<$1 2782 $<$1 2782 $<$1
-p_hat300-3_B 300 8388 25002 1144 10800 1209 10800 1299 $<$1 1299 $<$1 1299 $<$1
-brock200_2_C 200 2438 7438 1158 10800 1303 10800 1321 $<$1 1321 $<$1 1321 $<$1
-brock400_2_C 400 14751 45035 793 10800 961 10800 1033 $<$1 1033 $<$1 1033 $<$1
-C125.9_C 125 1733 5230 644 10800 644 45 644 $<$1 644 $<$1 644 $<$1
-C250.9_C 250 7073 20911 688 10800 734 10800 734 1 734 $<$1 734 $<$1
-gen200_p0.9_55_C 200 4425 13485 707 10800 727 9879 727 $<$1 727 $<$1 727 $<$1
-gen400_p0.9_75_C 400 18063 53757 613 10800 756 10800 772 1 772 $<$1 772 $<$1
-hamming8-4_C 256 5173 15691 1120 10800 1248 10800 1378 $<$1 1378 $<$1 1378 $<$1
-keller4_C 171 2400 7035 1060 10800 1109 10800 1109 $<$1 1109 $<$1 1109 $<$1
-MANN_a27_C 378 17580 52971 583 10800 648 10800 651 2 651 $<$1 651 $<$1
-p_hat300-1_C 300 2734 8199 1866 10800 3191 10800 3480 $<$1 3480 $<$1 3480 $<$1
-p_hat300-2_C 300 5603 16325 1511 10800 2249 10800 2473 $<$1 2473 $<$1 2473 $<$1
-p_hat300-3_C 300 8388 25002 886 10800 919 10800 1004 $<$1 1004 $<$1 1004 $<$1
+brock200_2_A 200 2,438 7,438 1,615 10,800 1,851 10,800 1,885 <<1 1,885 <<1 1,885 <<1
+brock400_2_A 400 14,751 45,035 1,403 10,800 1,667 10,800 1,728 5 1,728 <<1 1,728 <<1
+C125.9_A 125 1,733 5,230 1,023 247 1,023 4 1,023 <<1 1,023 <<1 1,023 <<1
+C250.9_A 250 7,073 20,911 1,200 10,800 1,236 5,143 1,236 2 1,236 <<1 1,236 <<1
+gen200_p0.9_55_A 200 4,425 13,485 1,186 10,800 1,206 601 1,206 <<1 1,206 <<1 1,206 <<1
+gen400_p0.9_75_A 400 18,063 53,757 1,155 10,800 1,490 10,800 1,490 <<1 1,490 <<1 1490 <<1
+hamming8-4_A 256 5,173 15,691 1,549 10,800 1,723 10,800 1,759 <<1 1,759 <<1 1,759 <<1
+keller4_A 171 2,400 7,035 1,400 10,800 1,434 1573 1,434 <<1 1,434 <<1 1,434 <<1
+MANN_a27_A 378 17,580 52,971 1,313 10,800 1,323 10,800 1,323 <<1 1,323 <<1 1,323 <<1
+p_hat300-1_A 300 2,734 8,199 2,504 10,800 3,932 10,800 4,164 <<1 4,164 <<1 4,164 <<1
+p_hat300-2_A 300 5,603 16,325 2,120 10,800 2,902 10,800 2,990 1 2,990 <<1 2,990 <<1
+p_hat300-3_A 300 8,388 25,002 1,350 10,800 1561 10,800 1,564 <<1 1,564 <<1 1,564 <<1
+brock200_2_B 200 2,438 7,438 1,416 10,800 1,602 10,800 1,641 <<1 1,641 <<1 1,641 <<1
+brock400_2_B 400 14,751 45,035 1,129 10,800 1,341 10,800 1,386 2 1,386 <<1 1,386 <<1
+C125.9_B 125 1,733 5,230 856 395 856 8 856 <<1 856 <<1 856 <<1
+C250.9_B 250 7,073 20,911 960 10,800 1,001 10,800 1,001 1 1,001 <<1 1,001 <<1
+gen200_p0.9_55_B 200 4,425 13,485 961 10,800 983 1585 983 <<1 983 <<1 983 <<1
+gen400_p0.9_75_B 400 18,063 53,757 920 10,800 1,105 10,800 1,120 <<1 1,120 <<1 1,120 <<1
+hamming8-4_B 256 5,173 15,691 1,356 10,800 1,486 10,800 1,579 <<1 1,579 <<1 1,579 <<1
+keller4_B 171 2,400 7,035 1,230 10,800 1,268 4468 1,268 <<1 1,268 <<1 1,268 <<1
+MANN_a27_B 378 17,580 52,971 922 10,800 1,011 10,800 1,021 <<1 1,021 <<1 1,021 <<1
+p_hat300-1_B 300 2,734 8,199 2,306 10,800 3,547 10,800 3,886 <<1 3,886 <<1 3,886 <<1
+p_hat300-2_B 300 5,603 16,325 1,762 10,800 2,588 10,800 2,782 <<1 2,782 <<1 2,782 <<1
+p_hat300-3_B 300 8,388 25,002 1,144 10,800 1,209 10,800 1,299 <<1 1,299 <<1 1,299 <<1
+brock200_2_C 200 2,438 7,438 1,158 10,800 1,303 10,800 1,321 <<1 1,321 <<1 1,321 <<1
+brock400_2_C 400 14,751 45,035 793 10,800 961 10,800 1,033 <<1 1,033 <<1 1,033 <<1
+C125.9_C 125 1,733 5,230 644 10,800 644 45 644 <<1 644 <<1 644 <<1
+C250.9_C 250 7,073 20,911 688 10,800 734 10,800 734 1 734 <<1 734 <<1
+gen200_p0.9_55_C 200 4,425 13,485 707 10,800 727 9,879 727 <<1 727 <<1 727 <<1
+gen400_p0.9_75_C 400 18,063 53,757 613 10,800 756 10,800 772 1 772 <<1 772 <<1
+hamming8-4_C 256 5,173 15,691 1,120 10,800 1,248 10,800 1,378 <<1 1,378 <<1 1,378 <<1
+keller4_C 171 2,400 7,035 1,060 10,800 1,109 10,800 1,109 <<1 1,109 <<1 1,109 <<1
+MANN_a27_C 378 17,580 52,971 583 10,800 648 10,800 651 2 651 <<1 651 <<1
+p_hat300-1_C 300 2,734 8,199 1,866 10,800 3,191 10,800 3,480 <<1 3,480 <<1 3,480 <<1
+p_hat300-2_C 300 5,603 16,325 1,511 10,800 2,249 10,800 2,473 <<1 2,473 <<1 2,473 <<1
+p_hat300-3_C 300 8,388 25,002 886 10,800 919 10,800 1,004 <<1 1,004 <<1 1,004 <<1
 $p=0.75$, Setting 2
-brock200_2_A 200 2438 7438 2999 10800 3078 10800 3326 8 3326 $<$1 3326 $<$1
-brock400_2_A 400 14751 45035 2603 10800 2787 10800 2941 12 2941 1 2941 3
-C125.9_A 125 1733 5230 1837 365 1837 62.71 1837 $<$1 1837 $<$1 1837 $<$1
-C250.9_A 250 7073 20911 2011 10800 2171 10800 2171 5 2171 $<$1 2171 $<$1
-gen200_p0.9_55_A 200 4425 13485 2094 10800 2096 7832 2096 $<$1 2096 $<$1 2096 $<$1
-gen400_p0.9_75_A 400 18063 53757 2164 10800 2310 10800 2404 19 2404 $<$1 2404 1
-hamming8-4_A 256 5173 15691 2813 10800 3124 10800 3124 1 3124 $<$1 3124 $<$1
-keller4_A 171 2400 7035 2532 10800 2690 10800 2690 $<$1 2690 $<$1 2690 $<$1
-MANN_a27_A 378 17580 52971 1976 10800 2206 10800 2208 $<$1 2208 $<$1 2208 $<$1
-p_hat300-1_A 300 2734 8199 5566 10800 $6$ 10800 7899 25 7899 $<$1 7899 $<$1
-p_hat300-2_A 300 5603 16325 4102 10800 4714 10800 5343 $<$1 5343 $<$1 5343 $<$1
-p_hat300-3_A 300 8388 25002 2477 10800 2658 10800 2838 2 2838 $<$1 2838 $<$1
-brock200_2_B 200 2438 7438 1398 10800 1427 10800 1533 $<$1 1533 $<$1 1533 $<$1
-brock400_2_B 400 14751 45035 1161 10800 1210 10800 1291 16 1291 1 1291 4
-C125.9_B 125 1733 5230 837 523 837 84 837 $<$1 837 $<$1 837 $<$1
-C250.9_B 250 7073 20911 911 10800 969 10800 971 3 971 $<$1 971 $<$1
-gen200_p0.9_55_B 200 4425 13485 944 10800 946 10800 946 $<$1 946 $<$1 946 $<$1
-gen400_p0.9_75_B 400 18063 53757 936 10800 1010 10800 1052 11 1054 1 1054 1
-hamming8-4_B 256 5173 15691 1309 10800 1474 10800 1474 $<$1 1474 $<$1 1474 $<$1
-keller4_B 171 2400 7035 1190 10800 1240 10800 1254 $<$1 1254 $<$1 1254 $<$1
-MANN_a27_B 378 17580 52971 875 10800 956 10800 958 11 958 $<$1 958 $<$1
-p_hat300-1_B 300 2734 8199 2666 10800 3216 10800 3818 2 3818 $<$1 3818 $<$1
-p_hat300-2_B 300 5603 16325 1952 10800 2177 10800 2543 $<$1 2543 $<$1 2543 $<$1
-p_hat300-3_B 300 8388 25002 1101 10800 1208 10800 1288 2 1288 $<$1 1288 $<$1
-brock200_2_C 200 2438 7438 607 10800 607 10800 658 $<$1 658 $<$1 658 $<$1
-brock400_2_C 400 14751 45035 461 10800 465 10800 500 3 500 $<$1 500 $<$1
-C125.9_C 125 1733 5230 337 1519 337 204 337 $<$1 337 $<$1 337 $<$1
-C250.9_C 250 7073 20911 359 10800 370 10800 372 5 372 $<$1 372 $<$1
-gen200_p0.9_55_C 200 4425 13485 352 10800 371 10800 371 $<$1 371 $<$1 371 $<$1
-gen400_p0.9_75_C 400 18063 53757 356 10800 379 10800 395 16 395 $<$1 395 $<$1
-hamming8-4_C 256 5173 15691 556 10800 649 10800 652 $<$1 652 $<$1 652 $<$1
-keller4_C 171 2400 7035 510 10800 554 10800 558 $<$1 558 $<$1 558 $<$1
-MANN_a27_C 378 17580 52971 325 10800 335 10800 335 11 335 $<$1 335 $<$1
-p_hat300-1_C 300 2734 8199 1205 10800 1491 10800 1793 1 1793 $<$1 1793 $<$1
-p_hat300-2_C 300 5603 16325 854 10800 981 10800 1159 $<$1 1159 $<$1 1159 $<$1
-p_hat300-3_C 300 8388 25002 453 10800 494 10800 529 1 529 $<$1 529 $<$1
+brock200_2_A 200 2,438 7,438 2,999 10,800 3,078 10,800 3,326 8 3,326 <<1 3,326 <<1
+brock400_2_A 400 14,751 45,035 2,603 10,800 2,787 10,800 2,941 12 2,941 1 2,941 3
+C125.9_A 125 1,733 5,230 1,837 365 1,837 62.71 1,837 <<1 1,837 <<1 1,837 <<1
+C250.9_A 250 7,073 20,911 2,011 10,800 2,171 10,800 2,171 5 2,171 <<1 2,171 <<1
+gen200_p0.9_55_A 200 4,425 13,485 2,094 10,800 2,096 7,832 2,096 <<1 2,096 <<1 2,096 <<1
+gen400_p0.9_75_A 400 18,063 53,757 2,164 10,800 2,310 10,800 2,404 19 2,404 <<1 2,404 1
+hamming8-4_A 256 5,173 15,691 2,813 10,800 3,124 10,800 3,124 1 3,124 <<1 3,124 <<1
+keller4_A 171 2,400 7,035 2,532 10,800 2,690 10,800 2,690 <<1 2,690 <<1 2,690 <<1
+MANN_a27_A 378 17,580 52,971 1,976 10,800 2,206 10,800 2,208 <<1 2,208 <<1 2,208 <<1
+p_hat300-1_A 300 2,734 8,199 5,566 10,800 6,666 10,800 7,899 25 7,899 <<1 7,899 <<1
+p_hat300-2_A 300 5,603 16,325 4,102 10,800 4,714 10,800 5,343 <<1 5,343 <<1 5,343 <<1
+p_hat300-3_A 300 8,388 25,002 2,477 10,800 2,658 10,800 2,838 2 2,838 <<1 2,838 <<1
+brock200_2_B 200 2,438 7,438 1,398 10,800 1,427 10,800 1,533 <<1 1,533 <<1 1,533 <<1
+brock400_2_B 400 14,751 45,035 1,161 10,800 1,210 10,800 1,291 16 1,291 1 1,291 4
+C125.9_B 125 1,733 5,230 837 523 837 84 837 <<1 837 <<1 837 <<1
+C250.9_B 250 7,073 20,911 911 10,800 969 10,800 971 3 971 <<1 971 <<1
+gen200_p0.9_55_B 200 4,425 13,485 944 10,800 946 10,800 946 <<1 946 <<1 946 <<1
+gen400_p0.9_75_B 400 18,063 53,757 936 10,800 1,010 10,800 1,052 11 1,054 1 1,054 1
+hamming8-4_B 256 5,173 15,691 1,309 10,800 1,474 10,800 1,474 <<1 1,474 <<1 1,474 <<1
+keller4_B 171 2,400 7,035 1,190 10,800 1,240 10,800 1,254 <<1 1,254 <<1 1,254 <<1
+MANN_a27_B 378 17,580 52,971 875 10,800 956 10,800 958 11 958 <<1 958 <<1
+p_hat300-1_B 300 2,734 8,199 2,666 10,800 3,216 10,800 3,818 2 3,818 <<1 3,818 <<1
+p_hat300-2_B 300 5,603 16,325 1,952 10,800 2,177 10,800 2,543 <<1 2,543 <<1 2,543 <<1
+p_hat300-3_B 300 8,388 25,002 1,101 10,800 1,208 10,800 1,288 2 1,288 <<1 1,288 <<1
+brock200_2_C 200 2,438 7,438 607 10,800 607 10,800 658 <<1 658 <<1 658 <<1
+brock400_2_C 400 14,751 45,035 461 10,800 465 10,800 500 3 500 <<1 500 <<1
+C125.9_C 125 1,733 5,230 337 1,519 337 204 337 <<1 337 <<1 337 <<1
+C250.9_C 250 7,073 20,911 359 10,800 370 10,800 372 5 372 <<1 372 <<1
+gen200_p0.9_55_C 200 4,425 13,485 352 10,800 371 10,800 371 <<1 371 <<1 371 <<1
+gen400_p0.9_75_C 400 18,063 53,757 356 10,800 379 10,800 395 16 395 <<1 395 <<1
+hamming8-4_C 256 5,173 15,691 556 10,800 649 10,800 652 <<1 652 <<1 652 <<1
+keller4_C 171 2,400 7,035 510 10,800 554 10,800 558 <<1 558 <<1 558 <<1
+MANN_a27_C 378 17,580 52,971 325 10,800 335 10,800 335 11 335 <<1 335 <<1
+p_hat300-1_C 300 2,734 8,199 1,205 10,800 1,491 10,800 1,793 1 1,793 <<1 1,793 <<1
+p_hat300-2_C 300 5,603 16,325 854 10,800 981 10,800 1,159 <<1 1,159 <<1 1,159 <<1
+p_hat300-3_C 300 8,388 25,002 453 10,800 494 10,800 529 1 529 <<1 529 <<1